The OMake user guide and reference manual

Jason Hickey, Aleksey Nogin, et. al. 15th March 2007

2

Contents

1 Guide 15

2 OMake quickstart guide 17

2.1 Description . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 2.1.1 Automatic dependency analysis . . . . . . . . . . . . . . . 17 2.1.2 Content-based dependency analysis . . . . . . . . . . . . . 17

2.2 For users already familiar with make . . . . . . . . . . . . . . . . 18 2.3 Building a small C program . . . . . . . . . . . . . . . . . . . . . 18 2.4 Larger projects . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 2.5 Subdirectories . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 2.6 Other things to consider . . . . . . . . . . . . . . . . . . . . . . . 24 2.7 Building OCaml programs . . . . . . . . . . . . . . . . . . . . . . 24 2.8 The OMakefile and OMakeroot files . . . . . . . . . . . . . . . . 26 2.9 Multiple version support . . . . . . . . . . . . . . . . . . . . . . . 27 2.10 Notes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27

3 Additional build examples 29 3.1 OMakeroot vs. OMakefile . . . . . . . . . . . . . . . . . . . . . . 31 3.2 An example C project . . . . . . . . . . . . . . . . . . . . . . . . 31 3.3 An example OCaml project . . . . . . . . . . . . . . . . . . . . . 33 3.4 Handling new languages . . . . . . . . . . . . . . . . . . . . . . . 35

3.4.1 Defining a default compilation rule . . . . . . . . . . . . . 36 3.4.2 Defining a rule for linking . . . . . . . . . . . . . . . . . . 36 3.4.3 Dependency scanning . . . . . . . . . . . . . . . . . . . . 37 3.4.4 Pulling it all together . . . . . . . . . . . . . . . . . . . . 39 3.4.5 Finishing up . . . . . . . . . . . . . . . . . . . . . . . . . 41

3.5 Collapsing the hierarchy, .SUBDIRS bodies . . . . . . . . . . . . 41 3.5.1 Using glob patterns . . . . . . . . . . . . . . . . . . . . . 42 3.5.2 Simplified sub-configurations . . . . . . . . . . . . . . . . 42 3.5.3 Computing the subdirectory list . . . . . . . . . . . . . . 43 3.5.4 Temporary directories . . . . . . . . . . . . . . . . . . . . 44 4 OMake concepts and syntax 47

4.1 Variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 4.2 Adding to a variable definition . . . . . . . . . . . . . . . . . . . 48 4.3 Arrays . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 4.4 Special characters and quoting . . . . . . . . . . . . . . . . . . . 48 4.5 Function definitions . . . . . . . . . . . . . . . . . . . . . . . . . 49 4.6 Comments . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 4.7 File inclusion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 4.8 Scoping, sections . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 4.9 Conditionals . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 4.10 Matching . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 4.11 Objects . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 4.12 Classes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 4.13 Inheritance . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 4.14 Special objects/sections . . . . . . . . . . . . . . . . . . . . . . . 55 4.15 private. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 4.16 protected. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 4.17 public. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 4.18 static. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 4.19 Short syntax for scoping objects . . . . . . . . . . . . . . . . . . 59 4.20 Modular programming . . . . . . . . . . . . . . . . . . . . . . . . 60

5 Expressions and values 61 5.1 Dynamic scoping . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 5.2 Functional evaluation . . . . . . . . . . . . . . . . . . . . . . . . 63 5.3 Exporting the environment . . . . . . . . . . . . . . . . . . . . . 64 5.4 Eager evaluation . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 5.5 Objects . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 5.6 Field and method calls . . . . . . . . . . . . . . . . . . . . . . . . 67 5.7 Method override . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 5.8 Super calls . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68

6 Additional language examples 69 6.1 Strings and arrays . . . . . . . . . . . . . . . . . . . . . . . . . . 69 6.2 Quoted strings . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 6.3 Files and directories . . . . . . . . . . . . . . . . . . . . . . . . . 71 6.4 Iteration, mapping, and foreach . . . . . . . . . . . . . . . . . . . 72 6.5 Lazy expressions . . . . . . . . . . . . . . . . . . . . . . . . . . . 73

6.5.1 A larger example of lazy expressions . . . . . . . . . . . . 74 6.6 Scoping and exports . . . . . . . . . . . . . . . . . . . . . . . . . 76 6.7 Shell aliases . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 6.8 Input/output redirection on the cheap . . . . . . . . . . . . . . . 78 7 Rules 79

7.1 Implicit rules . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 80 7.2 Bounded implicit rules . . . . . . . . . . . . . . . . . . . . . . . . 80 7.3 section . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 7.4 section rule . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 7.5 Special dependencies . . . . . . . . . . . . . . . . . . . . . . . . . 81

7.5.1 :exists: . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 7.5.2 :effects: . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 7.5.3 :value: . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82

7.6 .SCANNER rules . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 7.6.1 Named scanners, and the :scanner: dependencies …. 84 7.6.2 Notes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85

7.7 .DEFAULT . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 7.8 .SUBDIRS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85 7.9 .INCLUDE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 7.10 .PHONY . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 86 7.11 Rule scoping . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 87

7.11.1 Scoping of implicit rules . . . . . . . . . . . . . . . . . . . 88 7.11.2 Scoping of .SCANNER rules . . . . . . . . . . . . . . . . . . 88 7.11.3 Scoping for .PHONY targets . . . . . . . . . . . . . . . . . 89

7.12 Running OMake from a subdirectory . . . . . . . . . . . . . . . . 90 7.12.1 Phony targets in a subdirectory . . . . . . . . . . . . . . . 90 7.12.2 Hierarchy of .PHONY targets . . . . . . . . . . . . . . . . . 91

7.13 Pathnames in rules . . . . . . . . . . . . . . . . . . . . . . . . . . 91

8 Base library 93 8.1 Builtin variables . . . . . . . . . . . . . . . . . . . . . . . . . . . 93 8.2 Logic, Boolean functions, and control flow . . . . . . . . . . . . . 94

8.2.1 not . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 8.2.2 equal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 8.2.3 and . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 8.2.4 or . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 95 8.2.5 if . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 96 8.2.6 switch, match . . . . . . . . . . . . . . . . . . . . . . . . . 96 8.2.7 try . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 97 8.2.8 raise . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 8.2.9 exit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 8.2.10 defined . . . . . . . . . . . . . . . . . . . . . . . . . . . . 98 8.2.11 defined-env . . . . . . . . . . . . . . . . . . . . . . . . . . 99 8.2.12 getenv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 8.2.13 setenv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 99 8.2.14 unsetenv . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 8.2.15 get-registry . . . . . . . . . . . . . . . . . . . . . . . . . . 100 8.2.16 getvar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 100 8.2.17 setvar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101

8.3 Arrays and sequences . . . . . . . . . . . . . . . . . . . . . . . . . 101 8.3.1 array . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 8.3.2 split . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 101 8.3.3 concat . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 8.3.4 length . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 8.3.5 nth . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 8.3.6 nth-hd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 102 8.3.7 nth-tl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 8.3.8 subrange . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 8.3.9 rev . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 8.3.10 join . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 8.3.11 string . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 103 8.3.12 string-escaped, ocaml-escaped, html-escaped, html-pre-escaped,

c-escaped, id-escaped . . . . . . . . . . . . . . . . . . . . . 104 8.3.13 decode-uri, encode-uri . . . . . . . . . . . . . . . . . . . . 104 8.3.14 quote . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 105 8.3.15 quote-argv . . . . . . . . . . . . . . . . . . . . . . . . . . 105 8.3.16 html-string . . . . . . . . . . . . . . . . . . . . . . . . . . 105 8.3.17 addsuffix . . . . . . . . . . . . . . . . . . . . . . . . . . . 105 8.3.18 mapsuffix . . . . . . . . . . . . . . . . . . . . . . . . . . . 105 8.3.19 addsuffixes . . . . . . . . . . . . . . . . . . . . . . . . . . 106 8.3.20 removeprefix . . . . . . . . . . . . . . . . . . . . . . . . . 106 8.3.21 removesuffix . . . . . . . . . . . . . . . . . . . . . . . . . . 106 8.3.22 replacesuffixes . . . . . . . . . . . . . . . . . . . . . . . . . 106 8.3.23 addprefix . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 8.3.24 mapprefix . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 8.3.25 add-wrapper . . . . . . . . . . . . . . . . . . . . . . . . . 107 8.3.26 set . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 8.3.27 mem . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 107 8.3.28 intersection . . . . . . . . . . . . . . . . . . . . . . . . . . 108 8.3.29 intersects . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 8.3.30 set-diff . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 8.3.31 filter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 108 8.3.32 filter-out . . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 8.3.33 capitalize . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 8.3.34 uncapitalize . . . . . . . . . . . . . . . . . . . . . . . . . . 109 8.3.35 uppercase . . . . . . . . . . . . . . . . . . . . . . . . . . . 109 8.3.36 lowercase . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 8.3.37 system . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 8.3.38 shell . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 8.3.39 export . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 110 8.3.40 while . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 111 8.3.41 break . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 8.3.42 random, random-init . . . . . . . . . . . . . . . . . . . . . 112

8.4 Arithmetic . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 8.4.1 int . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 8.4.2 float . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 112 8.4.3 Basic arithmetic . . . . . . . . . . . . . . . . . . . . . . . 112 8.4.4 Comparisons . . . . . . . . . . . . . . . . . . . . . . . . . 113

8.5 First-class functions . . . . . . . . . . . . . . . . . . . . . . . . . 114 8.5.1 fun . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 8.5.2 apply . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 8.5.3 applya . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 114 8.5.4 create-map, create-lazy-map . . . . . . . . . . . . . . . . . 115

8.6 Iteration and mapping . . . . . . . . . . . . . . . . . . . . . . . . 115 8.6.1 foreach . . . . . . . . . . . . . . . . . . . . . . . . . . . . 115

9 File, I/O and system operations 117

9.1 File names . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117 9.1.1 file, dir . . . . . . . . . . . . . . . . . . . . . . . . . . . . 117 9.1.2 tmpfile . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118 9.1.3 in . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 118 9.1.4 basename . . . . . . . . . . . . . . . . . . . . . . . . . . . 118 9.1.5 dirname . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 9.1.6 rootname . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 9.1.7 dirof . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 9.1.8 fullname . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 9.1.9 absname . . . . . . . . . . . . . . . . . . . . . . . . . . . . 119 9.1.10 homename . . . . . . . . . . . . . . . . . . . . . . . . . . 120 9.1.11 suffix . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120

9.2 Path search . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120 9.2.1 which . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120 9.2.2 where . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120 9.2.3 rehash . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 120 9.2.4 exists-in-path . . . . . . . . . . . . . . . . . . . . . . . . . 121 9.2.5 digest . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 121 9.2.6 find-in-path . . . . . . . . . . . . . . . . . . . . . . . . . . 121 9.2.7 digest-path . . . . . . . . . . . . . . . . . . . . . . . . . . 121

9.3 File stats . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 122 9.3.1 file-exists, target-exists, target-is-proper . . . . . . . . . . 122 9.3.2 stat-reset . . . . . . . . . . . . . . . . . . . . . . . . . . . 122 9.3.3 filter-exists, filter-targets, filter-proper-targets . . . . . . . 122 9.3.4 find-targets-in-path, find-targets-in-path-optional . . . . . 123 9.3.5 file-sort . . . . . . . . . . . . . . . . . . . . . . . . . . . . 123

9.3.5.1 sort rule . . . . . . . . . . . . . . . . . . . . . . 124 9.3.6 file-check-sort . . . . . . . . . . . . . . . . . . . . . . . . . 124

9.4 Globbing and file listings . . . . . . . . . . . . . . . . . . . . . . . 125 9.4.1 glob . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 126 9.4.2 ls . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 127 9.4.3 subdirs . . . . . . . . . . . . . . . . . . . . . . . . . . . . 127

9.5 Filesystem operations . . . . . . . . . . . . . . . . . . . . . . . . 128 9.5.1 mkdir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 128 9.5.2 Stat . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 128 9.5.3 stat, lstat . . . . . . . . . . . . . . . . . . . . . . . . . . . 129 9.5.4 unlink . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 129 9.5.5 rename . . . . . . . . . . . . . . . . . . . . . . . . . . . . 129 9.5.6 link . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 130 9.5.7 symlink . . . . . . . . . . . . . . . . . . . . . . . . . . . . 130 9.5.8 readlink . . . . . . . . . . . . . . . . . . . . . . . . . . . . 130 9.5.9 chmod . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131 9.5.10 chown . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131 9.5.11 truncate . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131 9.5.12 umask . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 131

9.6 vmount . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 132 9.6.1 vmount . . . . . . . . . . . . . . . . . . . . . . . . . . . . 132 9.6.2 add-project-directories . . . . . . . . . . . . . . . . . . . . 132 9.6.3 remove-project-directories . . . . . . . . . . . . . . . . . . 132

9.7 File predicates . . . . . . . . . . . . . . . . . . . . . . . . . . . . 133 9.7.1 test . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 133 9.7.2 find . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135

9.8 IO functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 135 9.8.1 Standard channels . . . . . . . . . . . . . . . . . . . . . . 135 9.8.2 open-in-string . . . . . . . . . . . . . . . . . . . . . . . . . 136 9.8.3 open-out-string, out-string . . . . . . . . . . . . . . . . . . 136 9.8.4 fopen . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 136 9.8.5 close . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137 9.8.6 read . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137 9.8.7 write . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 137 9.8.8 lseek . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 138 9.8.9 rewind . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 138 9.8.10 tell . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 138 9.8.11 flush . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 138 9.8.12 dup . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139 9.8.13 dup2 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139 9.8.14 set-nonblock . . . . . . . . . . . . . . . . . . . . . . . . . 139 9.8.15 set-close-on-exec-mode . . . . . . . . . . . . . . . . . . . . 139 9.8.16 pipe . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 139 9.8.17 mkfifo . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 140 9.8.18 select . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 140 9.8.19 lockf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 140 9.8.20 InetAddr . . . . . . . . . . . . . . . . . . . . . . . . . . . 141 9.8.21 Host . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141 9.8.22 gethostbyname . . . . . . . . . . . . . . . . . . . . . . . . 141 9.8.23 Protocol . . . . . . . . . . . . . . . . . . . . . . . . . . . . 141 9.8.24 getprotobyname . . . . . . . . . . . . . . . . . . . . . . . 141 9.8.25 Service . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 142 9.8.26 getservbyname . . . . . . . . . . . . . . . . . . . . . . . . 142 9.8.27 socket . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 142 9.8.28 bind . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 143 9.8.29 listen . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 143 9.8.30 accept . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 143 9.8.31 connect . . . . . . . . . . . . . . . . . . . . . . . . . . . . 143 9.8.32 getchar . . . . . . . . . . . . . . . . . . . . . . . . . . . . 144 9.8.33 gets . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 144 9.8.34 fgets . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 144

9.9 Printing functions . . . . . . . . . . . . . . . . . . . . . . . . . . 144 9.10 Value printing functions . . . . . . . . . . . . . . . . . . . . . . . 145 9.11 Higher-level IO functions . . . . . . . . . . . . . . . . . . . . . . 145

9.11.1 Regular expressions . . . . . . . . . . . . . . . . . . . . . 145 9.11.2 cat . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 147 9.11.3 grep . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 148 9.11.4 scan . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 148 9.11.5 awk . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 150 9.11.6 fsubst . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 151 9.11.7 lex . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 152 9.11.8 lex-search . . . . . . . . . . . . . . . . . . . . . . . . . . . 153 9.11.9 Lexer . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 154 9.11.10 Lexer matching . . . . . . . . . . . . . . . . . . . . . . . . 155 9.11.11 Extending lexer definitions . . . . . . . . . . . . . . . . . 155 9.11.12 Threading the lexer object . . . . . . . . . . . . . . . . . 156 9.11.13 Parser . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 156 9.11.14 Calling the parser . . . . . . . . . . . . . . . . . . . . . . 158 9.11.15 Parsing control . . . . . . . . . . . . . . . . . . . . . . . . 158 9.11.16 Extending parsers . . . . . . . . . . . . . . . . . . . . . . 158 9.11.17 Passwd . . . . . . . . . . . . . . . . . . . . . . . . . . . . 159 9.11.18 getpwnam, getpwuid . . . . . . . . . . . . . . . . . . . . . 160 9.11.19 getpwents . . . . . . . . . . . . . . . . . . . . . . . . . . . 160 9.11.20 Group . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 160 9.11.21 getgrnam, getgrgid . . . . . . . . . . . . . . . . . . . . . . 160 9.11.22 tgetstr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 161 9.11.23 xterm-escape-begin, xterm-escape-end . . . . . . . . . . . 161 9.11.24 xterm-escape . . . . . . . . . . . . . . . . . . . . . . . . . 161 9.11.25 prompt-invisible-begin, prompt-invisible-end . . . . . . . . 161 9.11.26 prompt-invisible . . . . . . . . . . . . . . . . . . . . . . . 161 9.11.27 gettimeofday . . . . . . . . . . . . . . . . . . . . . . . . . 162

10 Shell commands 163 10.1 Simple commands . . . . . . . . . . . . . . . . . . . . . . . . . . 163 10.2 Globbing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 164 10.3 Background jobs . . . . . . . . . . . . . . . . . . . . . . . . . . . 164 10.4 File redirection . . . . . . . . . . . . . . . . . . . . . . . . . . . . 164 10.5 Pipelines . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 164 10.6 Conditional execution . . . . . . . . . . . . . . . . . . . . . . . . 165 10.7 Grouping . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 165 10.8 What is a shell command? . . . . . . . . . . . . . . . . . . . . . . 165 10.9 Basic builtin functions . . . . . . . . . . . . . . . . . . . . . . . . 166

10.9.1 echo . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 166 10.9.2 cd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 166

10.10Job control builtin functions . . . . . . . . . . . . . . . . . . . . . 167 10.10.1 jobs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 167 10.10.2 bg . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 167 10.10.3 fg . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 167 10.10.4 stop . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 167 10.10.5 wait . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 167 10.10.6 kill . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 167

10.11Command history . . . . . . . . . . . . . . . . . . . . . . . . . . 167 10.11.1 history . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 167

11 The standard objects 169

11.1 Pervasives objects . . . . . . . . . . . . . . . . . . . . . . . . . . 169 11.1.1 Object . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 169 11.1.2 Map . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 170 11.1.3 Number . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 11.1.4 Int . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 11.1.5 Float . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 171 11.1.6 Sequence . . . . . . . . . . . . . . . . . . . . . . . . . . . 172 11.1.7 Array . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 172 11.1.8 String . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173 11.1.9 Fun . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173 11.1.10 Rule . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173 11.1.11 Target . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 173 11.1.12 Node . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 174 11.1.13 File . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 174 11.1.14 Dir . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 174 11.1.15 Channel . . . . . . . . . . . . . . . . . . . . . . . . . . . . 174 11.1.16 InChannel . . . . . . . . . . . . . . . . . . . . . . . . . . . 175 11.1.17 OutChannel . . . . . . . . . . . . . . . . . . . . . . . . . . 175 11.1.18 Location . . . . . . . . . . . . . . . . . . . . . . . . . . . . 175 11.1.19 Position . . . . . . . . . . . . . . . . . . . . . . . . . . . . 175 11.1.20 Exception . . . . . . . . . . . . . . . . . . . . . . . . . . . 175 11.1.21 RuntimeException . . . . . . . . . . . . . . . . . . . . . . 176 11.1.22 UnbuildableException . . . . . . . . . . . . . . . . . . . . 176 11.1.23 Shell . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 176

12 Build functions and utilities 181 12.1 Builtin .PHONY targets . . . . . . . . . . . . . . . . . . . . . . . 181 12.2 Options and versioning . . . . . . . . . . . . . . . . . . . . . . . . 182

12.2.1 OMakeFlags . . . . . . . . . . . . . . . . . . . . . . . . . 182 12.2.2 OMakeVersion . . . . . . . . . . . . . . . . . . . . . . . . 182 12.2.3 cmp-versions . . . . . . . . . . . . . . . . . . . . . . . . . 183 12.2.4 DefineCommandVars . . . . . . . . . . . . . . . . . . . . . 183 12.3 Examining the dependency graph . . . . . . . . . . . . . . . . . . 183

12.3.1 dependencies, dependencies-all, dependencies-proper … 183 12.3.2 target . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 184 12.3.3 find-build-targets . . . . . . . . . . . . . . . . . . . . . . . 184 12.3.4 project-directories . . . . . . . . . . . . . . . . . . . . . . 184 12.3.5 rule . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 185

12.4 The OMakeroot file . . . . . . . . . . . . . . . . . . . . . . . . . . 185 12.4.1 Variables . . . . . . . . . . . . . . . . . . . . . . . . . . . 186 12.4.2 System variables . . . . . . . . . . . . . . . . . . . . . . . 186

12.5 Building C and C++ code . . . . . . . . . . . . . . . . . . . . . . 187

12.5.1 Autoconfiguration variables . . . . . . . . . . . . . . . . . 187 12.5.1.1 Unix-like systems . . . . . . . . . . . . . . . . . 187 12.5.1.2 Win32 . . . . . . . . . . . . . . . . . . . . . . . . 187

12.5.2 C and C++ configuration variables . . . . . . . . . . . . . 187 12.5.3 Generated C files . . . . . . . . . . . . . . . . . . . . . . . 189

12.5.3.1 CGeneratedFiles, LocalCGeneratedFiles . . . . . 189 12.5.4 Building C programs and Libraries . . . . . . . . . . . . . 189

12.5.4.1 StaticCLibrary . . . . . . . . . . . . . . . . . . . 189 12.5.4.2 StaticCLibraryCopy . . . . . . . . . . . . . . . . 190 12.5.4.3 StaticCLibraryInstall . . . . . . . . . . . . . . . 190 12.5.4.4 StaticCObject, StaticCObjectCopy, StaticCOb-

jectInstall . . . . . . . . . . . . . . . . . . . . . . 190

12.5.4.5 CProgram . . . . . . . . . . . . . . . . . . . . . 190 12.5.4.6 CProgramCopy . . . . . . . . . . . . . . . . . . 191 12.5.4.7 CProgramInstall . . . . . . . . . . . . . . . . . . 191 12.5.4.8 CXXProgram, CXXProgramInstall . . . . . . . 191 12.5.4.9 StaticCXXLibrary, StaticCXXLibraryCopy, Stat-

icCXXLibraryInstall . . . . . . . . . . . . . . . . 191

12.6 Building OCaml code . . . . . . . . . . . . . . . . . . . . . . . . 191 12.6.1 Autoconfiguration variables for OCaml compilation …. 192 12.6.2 Configuration variables for OCaml compilation . . . . . . 192 12.6.3 OCaml command flags . . . . . . . . . . . . . . . . . . . . 193 12.6.4 Library variables . . . . . . . . . . . . . . . . . . . . . . . 194 12.6.5 Generated OCaml Files . . . . . . . . . . . . . . . . . . . 195

12.6.5.1 OCamlGeneratedFiles, LocalOCamlGenerated-

Files . . . . . . . . . . . . . . . . . . . . . . . . . 195

12.6.5.2 Automatic discovery of generated files during

dependency analysis . . . . . . . . . . . . . . . . 195

12.6.6 Using the Menhir parser generator . . . . . . . . . . . . . 196 12.6.6.1 OCamlLibrary . . . . . . . . . . . . . . . . . . . 197 12.6.6.2 OCamlPackage . . . . . . . . . . . . . . . . . . . 197 12.6.6.3 OCamlLibraryCopy . . . . . . . . . . . . . . . . 197 12.6.6.4 OCamlLibraryInstall . . . . . . . . . . . . . . . 197 12.6.6.5 OCamlProgram . . . . . . . . . . . . . . . . . . 197 12.6.6.6 OCamlProgramCopy . . . . . . . . . . . . . . . 198 12.6.6.7 OCamlProgramInstall . . . . . . . . . . . . . . . 198 12.7 Building L^{A}TEX files . . . . . . . . . . . . . . . . . . . . . . . . . . 198

12.7.1 Configuration variables . . . . . . . . . . . . . . . . . . . 198 12.7.2 Building L^{A}TEX documents . . . . . . . . . . . . . . . . . . 199

12.7.2.1 LaTeXDocument . . . . . . . . . . . . . . . . . . 199 12.7.2.2 TeXGeneratedFiles, LocalTeXGeneratedFiles . . 200 12.7.2.3 LaTeXDocumentCopy . . . . . . . . . . . . . . . 200 12.7.2.4 LaTeXDocumentInstall . . . . . . . . . . . . . . 200

13 Autoconfiguration functions and variables 201

13.1 General-purpose autoconfiguration functions . . . . . . . . . . . . 201 13.1.1 ConfMsgChecking, ConfMsgResult . . . . . . . . . . . . . 201 13.1.2 ConfMsgWarn, ConfMsgError . . . . . . . . . . . . . . . . 202 13.1.3 ConfMsgYesNo, ConfMsgFound . . . . . . . . . . . . . . . 202 13.1.4 TryCompileC, TryLinkC, TryRunC . . . . . . . . . . . . . 202 13.1.5 RunCProg . . . . . . . . . . . . . . . . . . . . . . . . . . . 202 13.1.6 CheckCHeader, VerboseCheckCHeader . . . . . . . . . . . 203 13.1.7 CheckCLib, VerboseCheckCLib . . . . . . . . . . . . . . . 203 13.1.8 CheckProg . . . . . . . . . . . . . . . . . . . . . . . . . . 203

13.2 Translating autoconf scripts . . . . . . . . . . . . . . . . . . . . 203 13.3 Predefined configuration tests . . . . . . . . . . . . . . . . . . . . 204

13.3.1 NCurses library configuration . . . . . . . . . . . . . . . . 204 13.3.2 ReadLine library configuration . . . . . . . . . . . . . . . 204 13.3.3 Snprintf configuration . . . . . . . . . . . . . . . . . . . . 205

14 The OSH shell 207 14.1 Startup . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 207 14.2 Aliases . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 208 14.3 Interactive syntax . . . . . . . . . . . . . . . . . . . . . . . . . . 208

A Synopsis 209 A.1 General usage . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 209 A.2 Output control . . . . . . . . . . . . . . . . . . . . . . . . . . . . 209

A.2.1 -s . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 209 A.2.2 -S . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 209 A.2.3 -w . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 210 A.2.4 --progress . . . . . . . . . . . . . . . . . . . . . . . . . . 210 A.2.5 --print-status . . . . . . . . . . . . . . . . . . . . . . . 210 A.2.6 --print-exit . . . . . . . . . . . . . . . . . . . . . . . . 210 A.2.7 --verbose . . . . . . . . . . . . . . . . . . . . . . . . . . 210 A.2.8 --output-normal . . . . . . . . . . . . . . . . . . . . . . 210 A.2.9 --output-postpone . . . . . . . . . . . . . . . . . . . . . 210 A.2.10 --output-only-errors . . . . . . . . . . . . . . . . . . . 211 A.2.11 --output-at-end . . . . . . . . . . . . . . . . . . . . . . 211 A.2.12 -o . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 211

A.3 Build options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 212 A.3.1 -k . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 212 A.3.2 -n . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 212 A.3.3 -p . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 212 A.3.4 -P . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 212 A.3.5 -R . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 213 A.3.6 -t . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 213 A.3.7 -U . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 213 A.3.8 --depend . . . . . . . . . . . . . . . . . . . . . . . . . . . 213 A.3.9 --configure . . . . . . . . . . . . . . . . . . . . . . . . . 213 A.3.10 --force-dotomake . . . . . . . . . . . . . . . . . . . . . . 213 A.3.11 --dotomake . . . . . . . . . . . . . . . . . . . . . . . . . . 214 A.3.12 -j . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 214 A.3.13 --print-dependencies . . . . . . . . . . . . . . . . . . . 214 A.3.14 --show-dependencies . . . . . . . . . . . . . . . . . . . . 214 A.3.15 --all-dependencies . . . . . . . . . . . . . . . . . . . . 214 A.3.16 --verbose-dependencies . . . . . . . . . . . . . . . . . . 214 A.3.17 --install . . . . . . . . . . . . . . . . . . . . . . . . . . 215 A.3.18 --install-all . . . . . . . . . . . . . . . . . . . . . . . . 215 A.3.19 --install-force . . . . . . . . . . . . . . . . . . . . . . 215 A.3.20 --absname . . . . . . . . . . . . . . . . . . . . . . . . . . 215 A.3.21 variable definition . . . . . . . . . . . . . . . . . . . . 215

A.4 Additional options . . . . . . . . . . . . . . . . . . . . . . . . . . 215 A.5 Environment variables . . . . . . . . . . . . . . . . . . . . . . . . 215

A.5.1 OMAKEFLAGS . . . . . . . . . . . . . . . . . . . . . . . . . . 215 A.5.2 OMAKELIB . . . . . . . . . . . . . . . . . . . . . . . . . . . 216

A.6 Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 216

A.6.1 OMakeFlags . . . . . . . . . . . . . . . . . . . . . . . . . . 216 A.7 Option processing . . . . . . . . . . . . . . . . . . . . . . . . . . 216 A.8 .omakerc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 217

B OMake grammar 219

B.1 OMake lexical conventions . . . . . . . . . . . . . . . . . . . . . . 219 B.1.1 Comments . . . . . . . . . . . . . . . . . . . . . . . . . . . 219 B.1.2 Special characters . . . . . . . . . . . . . . . . . . . . . . 219 B.1.3 Identifiers . . . . . . . . . . . . . . . . . . . . . . . . . . . 220 B.1.4 Command identifiers . . . . . . . . . . . . . . . . . . . . . 220 B.1.5 Variable references . . . . . . . . . . . . . . . . . . . . . . 221 B.1.6 String constants . . . . . . . . . . . . . . . . . . . . . . . 221

B.2 The OMake grammar . . . . . . . . . . . . . . . . . . . . . . . . 222

B.2.1 Expressions . . . . . . . . . . . . . . . . . . . . . . . . . . 222

B.2.1.1 Inline applications . . . . . . . . . . . . . . . . . 223 B.2.2 Statements and programs . . . . . . . . . . . . . . . . . . 223

B.2.2.1 Special forms . . . . . . . . . . . . . . . . . . . . 225 B.2.2.2 Variable definitions . . . . . . . . . . . . . . . . 227 B.2.2.3 Applications and function definitions . . . . . . 227 B.2.2.4 Objects . . . . . . . . . . . . . . . . . . . . . . . 228 B.2.2.5 Rules . . . . . . . . . . . . . . . . . . . . . . . . 229

B.2.2.6 Shell commands . . . . . . . . . . . . . . . . . . 231

B.3 Dollar modifiers . . . . . . . . . . . . . . . . . . . . . . . . . . . . 232

C References 241 C.1 See Also . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 241 C.2 Version . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 241 C.3 License and Copyright . . . . . . . . . . . . . . . . . . . . . . . . 241 C.4 Author . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 241

Chapter 1

Guide

If you are new to OMake, you the omake-quickstart presents a short introduction that describes how to set up a project. The omake-build-examples gives larger examples of build projects, and omake-language-examples presents programming examples. Quickstart 2 A quickstart guide to using omake. Build examples 3 Advanced build examples. The OMake language 4 The omake language, including a description of ob-

jects, expressions, and values. Language discussion 5 Further discussion on the language, including scop-

ing, evaluation, and objects. Language examples 6 Additional language examples. Build rules 7 Defining and using rules to build programs. Base builtin functions 8 Functions and variables in the core standard li-

brary. System functions 9 Functions on files, input/output, and system commands. Shell commands 10 Using the omake shell for command-line interpretation. The standard objects 11 Pervasives defines the built-in objects. Standard build definitions 12 The build specifications for programming lan-

guages in the OMake standard library. Standard autoconfiguration functions and variables 13 The utilities provoded

by the OMake standard library to simplify programming of autoconfiguration tests.

The interactive command interpreter 14 The osh command-line interpreter.

Appendices OMake command-line options A Command-line options for

         omake.

The OMake language grammar B A more precise specification of the OMake language.

All the documentation on a single page All the OMake documentation in a single page.

OMake quickstart guide

2.1 Description

omake is designed for building projects that might have source files in several directories. Projects are normally specified using an OMakefile in each of the project directories, and an OMakeroot file in the root directory of the project. The OMakeroot file specifies general build rules, and the OMakefiles specify the build parameters specific to each of the subdirectories. When omake runs, it walks the configuration tree, evaluating rules from all of the OMakefiles. The project is then built from the entire collection of build rules.

2.1.1 Automatic dependency analysis

Dependency analysis has always been problematic with the make(1) program. omake addresses this by adding the .SCANNER target, which specifies a command to produce dependencies. For example, the following rule

    .SCANNER: %.o: %.c
        $(CC) $(INCLUDE) -MM $<

is the standard way to generate dependencies for .c files. omake will automatically run the scanner when it needs to determine dependencies for a file.

2.1.2 Content-based dependency analysis

Dependency analysis in omake uses MD5 digests to determine whether files have changed. After each run, omake stores the dependency information in a file called .omakedb in the project root directory. When a rule is considered for execution, the command is not executed if the target, dependencies, and command sequence are unchanged since the last run of omake. As an optimization, omake does not recompute the digest for a file that has an unchanged modification time, size, and inode number.

2.2 For users already familiar with make

For users already familiar with the make(1) command, here is a list of differences to keep in mind when using omake.

• In omake, you are much less likely to define build rules of your own. The system provides many standard functions (like StaticCLibrary and CProgram), described in Chapter 12, to specify these builds more simply.

• Implicit rules using .SUFFIXES and the .suf1.suf2: are not supported. You should use wildcard patterns instead %.suf2: %.suf1.

• Scoping is significant: you should define variables and .PHONY targets (see Section 7.10) before they are used.

• Subdirectories are incorporated into a project using the .SUBDIRS: target (see Section 7.8).

2.3 Building a small C program

To start a new project, the easiest method is to change directories to the project root and use the command omake --install to install default OMakefiles.

$ cd ~/newproject
$ omake --install
*** omake: creating OMakeroot
*** omake: creating OMakefile
*** omake: project files OMakefile and OMakeroot have been installed
*** omake: you should edit these files before continuing

The default OMakefile contains sections for building C and OCaml programs. For now, we’ll build a simple C project.

Suppose we have a C file called hello_code.c containing the following code:

#include <stdio.h>
int main(int argc, char **argv)
{
    printf("Hello world\n");
    return 0;
}

To build the program a program hello from this file, we can use the CProgram function. The OMakefile contains just one line that specifies that the program hello is to be built from the source code in the hello_code.c file (note that file suffixes are not passed to these functions).

CProgram(hello, hello_code)

Now we can run omake to build the project. Note that the first time we run omake, it both scans the hello_code.c file for dependencies, and compiles it using the cc compiler. The status line printed at the end indicates how many files were scanned, how many were built, and how many MD5 digests were computed.

$ omake hello
*** omake: reading OMakefiles
*** omake: finished reading OMakefiles (0.0 sec)
- scan . hello_code.o
+ cc -I. -MM hello_code.c
- build . hello_code.o
+ cc -I. -c -o hello_code.o hello_code.c
- build . hello
+ cc -o hello hello_code.o
*** omake: done (0.5 sec, 1/6 scans, 2/6 rules, 5/22 digests)
$ omake
*** omake: reading OMakefiles
*** omake: finished reading OMakefiles (0.1 sec)
*** omake: done (0.1 sec, 0/4 scans, 0/4 rules, 0/9 digests)

If we want to change the compile options, we can redefine the CC and CFLAGS variables before the CProgram line. In this example, we will use the gcc compiler with the -g option. In addition, we will specify a .DEFAULT target to be built by default. The EXE variable is defined to be .exe on Win32 systems; it is empty otherwise.

CC = gcc
CFLAGS += -g
CProgram(hello, hello_code)
.DEFAULT: hello$(EXE)

Here is the corresponding run for omake.

$ omake
*** omake: reading OMakefiles
*** omake: finished reading OMakefiles (0.0 sec)
- scan . hello_code.o
+ gcc -g -I. -MM hello_code.c
- build . hello_code.o
+ gcc -g -I. -c -o hello_code.o hello_code.c
- build . hello
+ gcc -g -o hello hello_code.o
*** omake: done (0.4 sec, 1/7 scans, 2/7 rules, 3/22 digests)

We can, of course, include multiple files in the program. Suppose we write a new file hello_helper.c. We would include this in the project as follows.

CC = gcc
CFLAGS += -g
CProgram(hello, hello_code hello_helper)
.DEFAULT: hello$(EXE)

2.4 Larger projects

As the project grows it is likely that we will want to build libraries of code. Libraries can be built using the StaticCLibrary function. Here is an example of an OMakefile with two libraries.

CC = gcc
CFLAGS += -g
FOO_FILES = foo_a foo_b
BAR_FILES = bar_a bar_b bar_c
StaticCLibrary(libfoo, $(FOO_FILES))
StaticCLibrary(libbar, $(BAR_FILES))
# The hello program is linked with both libraries
LIBS = libfoo libbar
CProgram(hello, hello_code hello_helper)
.DEFAULT: hello$(EXE)

2.5 Subdirectories

As the project grows even further, it is a good idea to split it into several directories. Suppose we place the libfoo and libbar into subdirectories.

In each subdirectory, we define an OMakefile for that directory. For example, here is an example OMakefile for the foo subdirectory.

INCLUDES += .. ../bar
FOO_FILES = foo_a foo_b
StaticCLibrary(libfoo, $(FOO_FILES))

Note the the INCLUDES variable is defined to include the other directories in the project.

Now, the next step is to link the subdirectories into the main project. The project OMakefile should be modified to include a .SUBDIRS: target.

# Project configuration
CC = gcc

CFLAGS += -g
# Subdirectories
.SUBDIRS: foo bar
# The libraries are now in subdirectories
LIBS = foo/libfoo bar/libbar
CProgram(hello, hello_code hello_helper)
.DEFAULT: hello$(EXE)

Note that the variables CC and CFLAGS are defined before the .SUBDIRS target. These variables remain defined in the subdirectories, so that libfoo

and libbar use gcc -g.

If the two directories are to be configured differently, we have two choices. The OMakefile in each subdirectory can be modified with its configuration (this is how it would normally be done). Alternatively, we can also place the change in the root OMakefile.

# Default project configuration
CC = gcc
CFLAGS += -g
# libfoo uses the default configuration
.SUBDIRS: foo
# libbar uses the optimizing compiler
CFLAGS += -O3
.SUBDIRS: bar
# Main program
LIBS = foo/libfoo bar/libbar
CProgram(hello, hello_code hello_helper)
.DEFAULT: hello$(EXE)

Note that the way we have specified it, the CFLAGS variable also contains

the -O3 option for the CProgram, and hello_code.c and hello_helper.c file

will both be compiled with the -O3 option. If we want to make the change truly local to libbar, we can put the bar subdirectory in its own scope using the

section form.
# Default project configuration
CC = gcc
CFLAGS += -g

# libfoo uses the default configuration
.SUBDIRS: foo
# libbar uses the optimizing compiler
section
    CFLAGS += -O3
    .SUBDIRS: bar
# Main program does not use the optimizing compiler
LIBS = foo/libfoo bar/libbar
CProgram(hello, hello_code hello_helper)
.DEFAULT: hello$(EXE)

Later, suppose we decide to port this project to Win32, and we discover that we need different compiler flags and an additional library.

# Default project configuration
if $(equal $(OSTYPE), Win32)
    CC = cl /nologo
    CFLAGS += /DWIN32 /MT
    export
else
    CC = gcc
    CFLAGS += -g
    export
# libfoo uses the default configuration
.SUBDIRS: foo
# libbar uses the optimizing compiler
section
    CFLAGS += $(if $(equal $(OSTYPE), Win32), $(EMPTY), -O3)
    .SUBDIRS: bar
# Default libraries
LIBS = foo/libfoo bar/libbar
# We need libwin32 only on Win32
if $(equal $(OSTYPE), Win32)
   LIBS += win32/libwin32
   .SUBDIRS: win32
   export
# Main program does not use the optimizing compiler

CProgram(hello, hello_code hello_helper)
.DEFAULT: hello$(EXE)

Note the use of the export directives to export the variable definitions from the if-statements. Variables in omake are scoped—variables in nested blocks (blocks with greater indentation), are not normally defined in outer blocks. The export directive specifies that the variable definitions in the nested blocks should be exported to their parent block.

Finally, for this example, we decide to copy all libraries into a common lib directory. We first define a directory variable, and replace occurrences of the lib string with the variable.

# The common lib directory
LIB = $(dir lib)
# phony target to build just the libraries
.PHONY: makelibs
# Default project configuration
if $(equal $(OSTYPE), Win32)
    CC = cl /nologo
    CFLAGS += /DWIN32 /MT
    export
else
    CC = gcc
    CFLAGS += -g
    export
# libfoo uses the default configuration
.SUBDIRS: foo
# libbar uses the optimizing compiler
section
    CFLAGS += $(if $(equal $(OSTYPE), Win32), $(EMPTY), -O3)
    .SUBDIRS: bar
# Default libraries
LIBS = $(LIB)/libfoo $(LIB)/libbar
# We need libwin32 only on Win32
if $(equal $(OSTYPE), Win32)
   LIBS += $(LIB)/libwin32
   .SUBDIRS: win32
   export

    # Main program does not use the optimizing compiler
    CProgram(hello, hello_code hello_helper)
    .DEFAULT: hello$(EXE)

In each subdirectory, we modify the OMakefiles in the library directories to install them into the $(LIB) directory. Here is the relevant change

to foo/OMakefile.
    INCLUDES += .. ../bar
    FOO_FILES = foo_a foo_b
    StaticCLibraryInstall(makelib, $(LIB), libfoo, $(FOO_FILES))

Directory (and file names) evaluate to relative pathnames. Within the foo directory, the $(LIB) variable evaluates to ../lib.

As another example, instead of defining the INCLUDES variable separately in each subdirectory, we can define it in the toplevel as follows.

    INCLUDES = $(ROOT) $(dir foo bar win32)

In the foo directory, the INCLUDES variable will evaluate to the string .. . ../bar ../win32. In the bar directory, it would be .. ../foo . ../win32. In the root directory

it would be . foo bar win32.

2.6 Other things to consider

omake also handles recursive subdirectories. For example, suppose the foo directory itself contains several subdirectories. The foo/OMakefile would then contain its own .SUBDIRS target, and each of its subdirectories would contain

its own OMakefile.

2.7 Building OCaml programs

By default, omake is also configured with functions for building OCaml programs. The functions for OCaml program use the OCaml prefix. For example, suppose we reconstruct the previous example in OCaml, and we have a file called hello_code.ml that contains the following code.

   open Printf
   let () = printf "Hello world\n"

An example OMakefile for this simple project would contain the following.

# Use the byte-code compiler
BYTE_ENABLED = true
NATIVE_ENABLED = false
OCAMLCFLAGS += -g
# Build the program
OCamlProgram(hello, hello_code)
.DEFAULT: hello.run

Next, suppose the we have two library subdirectories: the foo subdirectory is written in C, the bar directory is written in OCaml, and we need to use the standard OCaml Unix module.

# Default project configuration
if $(equal $(OSTYPE), Win32)
    CC = cl /nologo
    CFLAGS += /DWIN32 /MT
    export
else
    CC = gcc
    CFLAGS += -g
    export
# Use the byte-code compiler
BYTE_ENABLED = true
NATIVE_ENABLED = false
OCAMLCFLAGS += -g
# library subdirectories
INCLUDES += $(dir foo bar)
OCAMLINCLUDES += $(dir foo bar)
.SUBDIRS: foo bar
# C libraries
LIBS = foo/libfoo
# OCaml libraries
OCAML_LIBS = bar/libbar
# Also use the Unix module
OCAML_OTHER_LIBS = unix
# The main program
OCamlProgram(hello, hello_code hello_helper)
.DEFAULT: hello

The foo/OMakefile would be configured as a C library.

    FOO_FILES = foo_a foo_b
    StaticCLibrary(libfoo, $(FOO_FILES))

The bar/OMakefile would build an ML library.

   BAR_FILES = bar_a bar_b bar_c
   OCamlLibrary(libbar, $(BAR_FILES))

2.8 The OMakefile and OMakeroot files

OMake uses the OMakefile and OMakeroot files for configuring a project. The syntax of these files is the same, but their role is slightly different. For one thing, every project must have exactly one OMakeroot file in the project root directory. This file serves to identify the project root, and it contains code that sets up the project. In contrast, a multi-directory project will often have an OMakefile in each of the project subdirectories, specifying how to build the files in that subdirectory.

Normally, the OMakeroot file is boilerplate. The following listing is a typical example.

    include $(STDLIB)/build/Common
    include $(STDLIB)/build/C
    include $(STDLIB)/build/OCaml
    include $(STDLIB)/build/LaTeX
    # Redefine the command-line variables
    DefineCommandVars(.)
    # The current directory is part of the project
    .SUBDIRS: .

The include lines include the standard configuration files needed for the project. The $(STDLIB) represents the omake library directory. The only required configuration file is Common. The others are optional; for example, the $(STDLIB)/build/OCaml file is needed only when the project contains programs written in OCaml.

The DefineCommandVars function defines any variables specified on the command line (as arguments of the form VAR=<value>). The .SUBDIRS line specifies that the current directory is part of the project (so the OMakefile should be read).

Normally, the OMakeroot file should be small and project-independent. Any project-specific configuration should be placed in the OMakefiles of the project.

2.9 Multiple version support

OMake version 0.9.6 introduced preliminary support for multiple, simultaneous versions of a project. Versioning uses the vmount(dir1, dir2) function, which defines a “virtual mount” of directory dir1 over directory dir2. A “virtual mount” is like a transparent mount in Unix, where the files from dir1 appear in the dir2 namespace, but new files are created in dir2. More precisely, the filename dir2/foo refers to: a) the file dir1/foo if it exists, or b) dir2/foo otherwise.

The vmount function makes it easy to specify multiple versions of a project. Suppose we have a project where the source files are in the directory src/, and we want to compile two versions, one with debugging support and one optimized. We create two directories, debug and opt, and mount the src directory over them.

    section
        CFLAGS += -g
        vmount(-l, src, debug)
        .SUBDIRS: debug
    section
        CFLAGS += -O3
        vmount(-l, src, opt)
        .SUBDIRS: opt

Here, we are using section blocks to define the scope of the vmount—you may not need them in your project.

The -l option is optional. It specifies that files form the src directory should be linked into the target directories (or copied, if the system is Win32). The links are added as files are referenced. If no options are given, then files are not copied or linked, but filenames are translated to refer directly to the src/ files.

Now, when a file is referenced in the debug directory, it is linked from the src directory if it exists. For example, when the file debug/OMakefile is read, the src/OMakefile is linked into the debug/ directory.

The vmount model is fairly transparent. The OMakefiles can be written as if referring to files in the src/ directory—they need not be aware of mounting. However, there are a few points to keep in mind.

2.10 Notes

• When using the vmount function for versioning, it wise to keep the source files distinct from the compiled versions. For example, suppose the source directory contained a file src/foo.o. When mounted, the foo.o file will be the same in all versions, which is probably not what you want. It is better to keep the src/ directory pristine, containing no compiled code.

• When using the vmount -l option, files are linked into the version directory only if they are referenced in the project. Functions that examine the filesystem (like $(ls ...)) may produce unexpected results.

Additional build examples

Let’s explain the OMake build model a bit more. One issue that dominates this discussion is that OMake is based on global project analysis. That means you define a configuration for the entire project, and you run one instance of omake.

For single-directory projects this doesn’t mean much. For multi-directory projects it means a lot. With GNU make, you would usually invoke the make program recursively for each directory in the project. For example, suppose you had a project with some project root directory, containing a directory of sources src, which in turn contains subdirectories lib and main. So your project looks like this nice piece of ASCII art.

    my_project/
    |--> Makefile
    `--> src/
         |---> Makefile
         |---> lib/
         |     |---> Makefile
         |     `---> source files...
         `---> main/
               |---> Makefile
               `---> source files...

Typically, with GNU make, you would start an instance of make in my_project/; this would in term start an instance of make in the src/ directory; and this would start new instances in lib/ and main/. Basically, you count up the number of Makefiles in the project, and that is the number of instances of make processes that will be created.

The number of processes is no big deal with today’s machines (sometimes contrary the the author’s opinion, we no longer live in the 1970s). The problem with the scheme was that each make process had a separate configuration, and it took a lot of work to make sure that everything was consistent. Furthermore, suppose the programmer runs make in the main/ directory, but the lib/ is outof-date. In this case, make would happily crank away, perhaps trying to rebuild files in lib/, perhaps just giving up.

With OMake this changes entirely. Well, not entirely. The source structure is quite similar, we merely add some Os to the ASCII art.

    my_project/
    |--> OMakeroot   (or Root.om)
    |--> OMakefile
    `--> src/
         |---> OMakefile
         |---> lib/
         |     |---> OMakefile
         |     `---> source files...
         `---> main/
               |---> OMakefile
               `---> source files...

The role of each <dir>/OMakefile plays the same role as each <dir>/Makefile: it describes how to build the source files in <dir>. The OMakefile retains much of syntax and structure of the Makefile, but in most cases it is much simpler.

One minor difference is the presence of the OMakeroot in the project root. The main purpose of this file is to indicate where the project root is in the first place (in case omake is invoked from a subdirectory). The OMakeroot serves as the bootstrap file; omake starts by reading this file first. Otherwise, the syntax and evaluation of OMakeroot is no different from any other OMakefile.

The big difference is that OMake performs a global analysis. Here is what happens when omake starts.

1. omake locates that OMakeroot file, and reads it.

2. Each OMakefile points to its subdirectory OMakefiles using the .SUBDIRS target. For example, my_project/OMakefile has a rule,

         .SUBDIRS: src
     and the my_project/src/OMakefile has a rule,
         .SUBDIRS: lib main

omake uses these rules to read and evaluate every OMakefile in the project. Reading and evaluation is fast. This part of the process is cheap.

3. Now that the entire configuration is read, omake determines which files are out-of-date (using a global analysis), and starts the build process. This may take a while, depending on what exactly needs to be done.

There are several advantages to this model. First, since analysis is global, it is much easier to ensure that the build configuration is consistent–after all, there is only one configuration. Another benefit is that the build configuration is inherited, and can be re-used, down the hierarchy. Typically, the root OMakefile defines some standard boilerplate and configuration, and this is inherited by subdirectories that tweak and modify it (but do not need to restate it entirely). The disadvantage of course is space, since this is global analysis after all. In practice rarely seems to be a concern; omake takes up much less space than your web browser even on large projects.

Some notes to the GNU/BSD make user.

• OMakefiles are a lot like Makefiles. The syntax is similar, and there many of the builtin functions are similar. However, the two build systems are not the same. Some evil features (in the authors’ opinions) have been dropped in OMake, and some new features have been added.

• OMake works the same way on all platforms, including Win32. The standard configuration does the right thing, but if you care about porting your code to multiple platforms, and you use some tricky features, you may need to condition parts of your build config on the $(OSTYPE) variable.

• A minor issue is that OMake dependency analysis is based on MD5 file digests. That is, dependencies are based on file contents, not file modification times. Say goodbye to false rebuilds based on spurious timestamp changes and mismatches between local time and fileserver time.

3.1 OMakeroot vs. OMakefile

Before we begin with examples, let’s ask the first question, “What is the difference between the project root OMakeroot and OMakefile?” A short answer is, there is no difference, but you must have an OMakeroot file (or Root.om file).

However, the normal style is that OMakeroot is boilerplate and is moreor-less the same for all projects. The OMakefile is where you put all your project-specific stuff.

To get started, you don’t have to do this yourself. In most cases you just perform the following step in your project root directory.

• Run omake --install in your project root.

This will create the initial OMakeroot and OMakefile files that you can edit to get started.

3.2 An example C project

To begin, let’s start with a simple example. Let’s say that we have a full directory tree, containing the following files.

my_project/
|--> OMakeroot
|--> OMakefile
`--> src/
     |---> OMakefile
     |---> lib/
     |     |---> OMakefile
     |     |---> ouch.c
     |     |---> ouch.h
     |     `---> bandaid.c
     `---> main/
           |---> OMakefile
           |---> horsefly.c
           |---> horsefly.h
           `---> main.c

Here is an example listing.

my_project/OMakeroot:
# Include the standard configuration for C applications
open build/C
# Process the command-line vars
DefineCommandVars()
# Include the OMakefile in this directory.
.SUBDIRS: .
my_project/OMakefile:
# Set up the standard configuration
CFLAGS += -g
# Include the src subdirectory
.SUBDIRS: src
my_project/src/OMakefile:
# Add any extra options you like
CFLAGS += -O2
# Include the subdirectories
.SUBDIRS: lib main
my_project/src/lib/OMakefile:
# Build the library as a static library.
# This builds libbug.a on Unix/OSX, or libbug.lib on Win32.
# Note that the source files are listed _without_suffix.

StaticCLibrary(libbug, ouch bandaid)
my_project/src/main/OMakefile:
# Some files include the .h files in ../lib
INCLUDES += ../lib
# Indicate which libraries we want to link against.
LIBS[] +=
    ../lib/libbug
# Build the program.
# Builds horsefly.exe on Win32, and horsefly on Unix.
# The first argument is the name of the executable.
# The second argument is an array of object files (without suffix)
# that are part of the program.
CProgram(horsefly, horsefly main)
# Build the program by default (in case omake is called
# without any arguments).  EXE is defined as .exe on Win32,
# otherwise it is empty.
.DEFAULT: horsefly$(EXE)

Most of the configuration here is defined in the file build/C.om (which is part of the OMake distribution). This file takes care of a lot of work, including:

• Defining the StaticCLibrary and CProgram functions, which describe the canonical way to build C libraries and programs.

• Defining a mechanism for scanning each of the source programs to discover dependencies. That is, it defines .SCANNER rules for C source files.

Variables are inherited down the hierarchy, so for example, the value of CFLAGS in src/main/OMakefile is “-g -O2”.

3.3 An example OCaml project

Let’s repeat the example, assuming we are using OCaml instead of C. This time, the directory tree looks like this.

my_project/
|--> OMakeroot
|--> OMakefile
`--> src/
     |---> OMakefile
     |---> lib/
     |     |---> OMakefile
     |     |---> ouch.ml

     |     |---> ouch.mli
     |     `---> bandaid.ml
     `---> main/
           |---> OMakefile
           |---> horsefly.ml
           |---> horsefly.mli
           `---> main.ml

The listing is only a bit different.

my_project/OMakeroot:
# Include the standard configuration for OCaml applications
open build/OCaml
# Process the command-line vars
DefineCommandVars()
# Include the OMakefile in this directory.
.SUBDIRS: .
my_project/OMakefile:
# Set up the standard configuration
OCAMLFLAGS += -Wa
# Do we want to use the bytecode compiler,
# or the native-code one?  Let's use both for
# this example.
NATIVE_ENABLED = true
BYTE_ENABLED = true
# Include the src subdirectory
.SUBDIRS: src
my_project/src/OMakefile:
# Include the subdirectories
.SUBDIRS: lib main
my_project/src/lib/OMakefile:
# Let's do aggressive inlining on native code
OCAMLOPTFLAGS += -inline 10
# Build the library as a static library.
# This builds libbug.a on Unix/OSX, or libbug.lib on Win32.
# Note that the source files are listed _without_suffix.
OCamlLibrary(libbug, ouch bandaid)

my_project/src/main/OMakefile:
# These files depend on the interfaces in ../lib
OCAMLINCLUDES += ../lib
# Indicate which libraries we want to link against.
OCAML_LIBS[] +=
    ../lib/libbug
# Build the program.
# Builds horsefly.exe on Win32, and horsefly on Unix.
# The first argument is the name of the executable.
# The second argument is an array of object files (without suffix)
# that are part of the program.
OCamlProgram(horsefly, horsefly main)
# Build the program by default (in case omake is called
# without any arguments).  EXE is defined as .exe on Win32,
# otherwise it is empty.
.DEFAULT: horsefly$(EXE)

In this case, most of the configuration here is defined in the file build/OCaml.om. In this particular configuration, files in my_project/src/lib are compiled aggressively with the option -inline 10, but files in my_project/src/lib are compiled normally.

3.4 Handling new languages

The previous two examples seem to be easy enough, but they rely on the OMake standard library (the files build/C and build/OCaml) to do all the work. What happens if we want to write a build configuration for a language that is not already supported in the OMake standard library?

For this example, let’s suppose we are adopting a new language. The language uses the standard compile/link model, but is not in the OMake standard library. Specifically, let’s say we have the following setup.

• Source files are defined in files with a .cat suffix (for Categorical Abstract Terminology).

• .cat files are compiled with the catc compiler to produce .woof files (Wicked Object-Oriented Format).

• .woof files are linked by the catc compiler with the -c option to produce a .dog executable (Digital Object Group). The catc also defines a -a option to combine several .woof files into a library.

• Each .cat can refer to other source files. If a source file a.cat contains a line open b, then a.cat depends on the file b.woof, and a.cat must be recompiled if b.woof changes. The catc function takes a -I option to define a search path for dependencies.

To define a build configuration, we have to do three things.

1. Define a .SCANNER rule for discovering dependency information for the source files.

2. Define a generic rule for compiling a .cat file to a .woof file.

3. Define a rule (as a function) for linking .woof files to produce a .dog executable.

Initially, these definitions will be placed in the project root OMakefile.

3.4.1 Defining a default compilation rule

Let’s start with part 2, defining a generic compilation rule. We’ll define the build rule as an implicit rule. To handle the include path, we’ll define a variable CAT_INCLUDES that specifies the include path. This will be an array of directories. To define the options, we’ll use a lazy variable (Section 6.5). In case there are any other standard flags, we’ll define a CAT_FLAGS variable.

   # Define the catc command, in case we ever want to override it
   CATC = catc
   # The default flags are empty
   CAT_FLAGS =
   # The directories in the include path (empty by default)
   INCLUDES[] =
   # Compute the include options from the include path
   PREFIXED_INCLUDES[] = $`(mapprefix -I, $(INCLUDES))
   # The default way to build a .woof file
   %.woof: %.cat
       $(CATC) $(PREFIXED_INCLUDES) $(CAT_FLAGS) -c $<

The final part is the build rule itself, where we call the catc compiler with the include path, and the CAT_FLAGS that have been defined. The $< variable represents the source file.

3.4.2 Defining a rule for linking

For linking, we’ll define another rule describing how to perform linking. Instead of defining an implicit rule, we’ll define a function that describes the linking step. The function will take two arguments; the first is the name of the executable (without suffix), and the second is the files to link (also without suffixes). Here is the code fragment.

    # Optional link options
    CAT_LINK_FLAGS =
    # The function that defines how to build a .dog program
    CatProgram(program, files) =
        # Add the suffixes
        file_names = $(addsuffix .woof, $(files))
        prog_name = $(addsuffix .dog, $(files))
        # The build rule
        $(prog_name): $(file_names)
            $(CATC) $(PREFIXED_INCLUDES) $(CAT_FLAGS) $(CAT_LINK_FLAGS) -o $@ $+
        # Return the program name
        value $(prog_name)

The CAT_LINK_FLAGS variable is defined just in case we want to pass additional flags specific to the link step. Now that this function is defined, whenever we want to define a rule for building a program, we simply call the rule. The previous implicit rule specifies how to compile each source file, and the CatProgram function specifies how to build the executable.

    # Build a rover.dog program from the source
    # files neko.cat and chat.cat.
    # Compile it by default.
    .DEFAULT: $(CatProgram rover, neko chat)

3.4.3 Dependency scanning

That’s it, almost. The part we left out was automated dependency scanning. This is one of the nicer features of OMake, and one that makes build specifications easier to write and more robust. Strictly speaking, it isn’t required, but you definitely want to do it.

The mechanism is to define a .SCANNER rule, which is like a normal rule, but it specifies how to compute dependencies, not the target itself. In this case, we want to define a .SCANNER rule of the following form.

    .SCANNER: %.woof: %.cat
        <commands>

This rule specifies that a .woof file may have additional dependencies that can be extracted from the corresponding .cat file by executing the <commands>. The result of executing the <commands> should be a sequence of dependencies in OMake format, printed to the standard output.

As we mentioned, each .cat file specifies dependencies on .woof files with an open directive. For example, if the neko.cat file contains a line open chat, then neko.woof depends on chat.woof. In this case, the <commands> should print the following line.

    neko.woof: chat.woof

For an analogy that might make this clearer, consider the C programming language, where a .o file is produced by compiling a .c file. If a file foo.c contains a line like #include "fum.h", then foo.c should be recompiled whenever fum.h changes. That is, the file foo.o depends on the file fum.h. In the OMake parlance, this is called an implicit dependency, and the .SCANNER <commands> would print a line like the following.

    foo.o: fum.h

Now, returning to the animal world, to compute the dependencies of neko.woof, we should scan neko.cat, line-by-line, looking for lines of the form open <name>. We could do this by writing a program, but it is easy enough to do it in omake itself. We can use the builtin awk function to scan the source file. One slight complication is that the dependencies depend on the INCLUDE path. We’ll use the find-in-path function to find them. Here we go.

    .SCANNER: %.woof: %.cat
        section
            # Scan the file
            deps[] =
            awk($<)
            case $'^open'
                deps[] += $2
                export
            # Remove duplicates, and find the files in the include path
            deps = $(find-in-path $(INCLUDES), $(set $(deps)))
            # Print the dependencies
            println($"$@: $(deps)")

Let’s look at the parts. First, the entire body is defined in a section because we are computing it internally, not as a sequence of shell commands.

We use the deps variable to collect all the dependencies. The awk function scans the source file ($<) line-by-line. For lines that match the regular expression ^open (meaning that the line begins with the word open), we add the second word on the line to the deps variable. For example, if the input line is open chat, then we would add the chat string to the deps array. All other lines in the source file are ignored.

Next, the $(set $(deps)) expression removes any duplicate values in the deps array (sorting the array alphabetically in the process). The find-in-path function then finds the actual location of each file in the include path.

The final step is print the result as the string $"$@: $(deps)" The quotations are added to flatten the deps array to a simple string.

3.4.4 Pulling it all together

To complete the example, let’s pull it all together into a single project, much like our previous example.

 my_project/
 |--> OMakeroot
 |--> OMakefile
 `--> src/
      |---> OMakefile
      |---> lib/
      |     |---> OMakefile
      |     |---> neko.cat
      |     `---> chat.cat
      `---> main/
            |---> OMakefile
            `---> main.cat

The listing for the entire project is as follows. Here, we also include a function CatLibrary to link several .woof files into a library.

my_project/OMakeroot:
 # Process the command-line vars
 DefineCommandVars()
 # Include the OMakefile in this directory.
 .SUBDIRS: .
my_project/OMakefile:
########################################################################
# Standard config for compiling .cat files
#
# Define the catc command, in case we ever want to override it
CATC = catc
# The default flags are empty
CAT_FLAGS =
# The directories in the include path (empty by default)
INCLUDES[] =
# Compute the include options from the include path
PREFIXED_INCLUDES[] = $`(mapprefix -I, $(INCLUDES))
# Dependency scanner for .cat files
.SCANNER: %.woof: %.cat

 section
     # Scan the file
     deps[] =
     awk($<)
     case $'^open'
         deps[] += $2
         export
     # Remove duplicates, and find the files in the include path
     deps = $(find-in-path $(INCLUDES), $(set $(deps)))
     # Print the dependencies
     println($"$@: $(deps)")
# The default way to compile a .cat file
%.woof: %.cat
$(CATC) $(PREFIXED_INCLUDES) $(CAT_FLAGS) -c $<
# Optional link options
CAT_LINK_FLAGS =
# Build a library for several .woof files
CatLibrary(lib, files) =
# Add the suffixes
file_names = $(addsuffix .woof, $(files))
lib_name = $(addsuffix .woof, $(lib))
# The build rule
$(lib_name): $(file_names)
    $(CATC) $(PREFIXED_INCLUDES) $(CAT_FLAGS) $(CAT_LINK_FLAGS) -a $@ $+
# Return the program name
value $(lib_name)
# The function that defines how to build a .dog program
CatProgram(program, files) =
# Add the suffixes
file_names = $(addsuffix .woof, $(files))
prog_name = $(addsuffix .dog, $(program))
# The build rule
$(prog_name): $(file_names)
    $(CATC) $(PREFIXED_INCLUDES) $(CAT_FLAGS) $(CAT_LINK_FLAGS) -o $@ $+
# Return the program name
value $(prog_name)

   ########################################################################
   # Now the program proper
   #
   # Include the src subdirectory
   .SUBDIRS: src
my_project/src/OMakefile:
   .SUBDIRS: lib main
my_project/src/lib/OMakefile:
   CatLibrary(cats, neko chat)
my_project/src/main/OMakefile:
   # Allow includes from the ../lib directory
   INCLUDES[] += ../lib
   # Build the program
   .DEFAULT: $(CatProgram main, main ../cats)

Some notes. The configuration in the project OMakeroot defines the standard configuration, including the dependency scanner, the default rule for compiling source files, and functions for building libraries and programs.

These rules and functions are inherited by subdirectories, so the .SCANNER and build rules are used automatically in each subdirectory, so you don’t need to repeat them.

3.4.5 Finishing up

At this point we are done, but there are a few things we can consider.

First, the rules for building cat programs is defined in the project OMakefile. If you had another cat project somewhere, you would need to copy the OMakeroot (and modify it as needed). Instead of that, you should consider moving the configuration to a shared library directory, in a file like Cat.om. That way, instead of copying the code, you could include the shared copy with an OMake command open Cat. The share directory should be added to your OMAKEPATH environment variable to ensure that omake knows how to find it.

Better yet, if you are happy with your work, consider submitting it as a standard configuration (by sending a request to omake@metaprl.org) so that others can make use of it too.

3.5 Collapsing the hierarchy, .SUBDIRS bodies

Some projects have many subdirectories that all have the same configuration. For instance, suppose you have a project with many subdirectories, each containing a set of images that are to be composed into a web page. Apart from the specific images, the configuration of each file is the same.

To make this more concrete, suppose the project has four subdirectories

page1, page2, page3, and page4. Each contains two files image1.jpg and

image2.jpg that are part of a web page generated by a program genhtml.

Instead of of defining a OMakefile in each directory, we can define it as a body to the .SUBDIRS command.

    .SUBDIRS: page1 page2 page3 page4
        index.html: image1.jpg image2jpg
            genhtml $+ > $@

The body of the .SUBDIRS is interpreted exactly as if it were the OMakefile, and it can contain any of the normal statements. The body is evaluated in the subdirectory for each of the subdirectories. We can see this if we add a statement that prints the current directory ($(CWD)).

    .SUBDIRS: page1 page2 page3 page4
        println($(absname $(CWD)))
        index.html: image1.jpg image2jpg
            genhtml $+ > $@
  # prints
    /home/jyh/.../page1
    /home/jyh/.../page2
    /home/jyh/.../page3
    /home/jyh/.../page4

3.5.1 Using glob patterns

Of course, this specification is quite rigid. In practice, it is likely that each subdirectory will have a different set of images, and all should be included in the web page. One of the easier solutions is to use one of the directory-listing functions, like glob or ls. The glob function takes a shell pattern, and returns an array of file with matching filenames in the current directory.

    .SUBDIRS: page1 page2 page3 page4
        IMAGES = $(glob *.jpg)
        index.html: $(IMAGES)
            genhtml $+ > $@

3.5.2 Simplified sub-configurations

Another option is to add a configuration file in each of the subdirectories that defines directory-specific information. For this example, we might define a file BuildInfo.om in each of the subdirectories that defines a list of images in that directory. The .SUBDIRS line is similar, but we include the BuildInfo file.

    .SUBDIRS: page1 page2 page3 page4
        include BuildInfo   # Defines the IMAGES variable
        index.html: $(IMAGES)
            genhtml $+ > $@

Where we might have the following configurations.

   page1/BuildInfo.om:
       IMAGES[] = image.jpg
   page2/BuildInfo.om:
       IMAGES[] = ../common/header.jpg winlogo.jpg
   page3/BuildInfo.om:
       IMAGES[] = ../common/header.jpg unixlogo.jpg daemon.jpg
   page4/BuildInfo.om:
       IMAGES[] = fee.jpg fi.jpg foo.jpg fum.jpg

3.5.3 Computing the subdirectory list

The other hardcoded specification is the list of subdirectories page1, ..., page4. Rather than editing the project OMakefile each time a directory is added, we could compute it (again with glob).

    .SUBDIRS: $(glob page*)
        index.html: $(glob *.jpg)
            genhtml $+ > $@

Alternately, the directory structure may be hierarchical. Instead of using glob, we could use the subdirs function, returns each of the directories in a hierarchy. For example, this is the result of evaluating the subdirs function in the omake project root. The P option, passed as the first argument, specifies that the listing is “proper,” it should not include the omake directory itself.

    osh> subdirs(P, .)
    - : <array
            /home/jyh/.../omake/mk : Dir
            /home/jyh/.../omake/RPM : Dir
            ...
            /home/jyh/.../omake/osx_resources : Dir>

Using subdirs, our example is now as follows.

    .SUBDIRS: $(subdirs P, .)
        index.html: $(glob *.jpg)
            genhtml $+ > $@

In this case, every subdirectory will be included in the project.

If we are using the BuildInfo.om option. Instead of including every subdirectory, we could include only those that contain a BuildInfo.om file. For this purpose, we can use the find function, which traverses the directory hierarchy looking for files that match a test expression. In our case, we want to search for files with the name BuildInfo.om. Here is an example call.

    osh> FILES = $(find . -name BuildInfo.om)
    - : <array
            /home/jyh/.../omake/doc/html/BuildInfo.om : File
            /home/jyh/.../omake/src/BuildInfo.om : File
            /home/jyh/.../omake/tests/simple/BuildInfo.om : File>
    osh> DIRS = $(dirof $(FILES))
    - : <array
            /home/jyh/.../omake/doc/html : Dir
            /home/jyh/.../omake/src : Dir
            /home/jyh/.../omake/tests/simple : Dir>

In this example, there are three BuildInfo.om files, in the doc/html, src, and tests/simple directories. The dirof function returns the directories for each of the files.

Returning to our original example, we modify it as follows.

    .SUBDIRS: $(dirof $(find . -name BuildInfo.om))
        include BuildInfo   # Defines the IMAGES variable
        index.html: $(IMAGES)
            genhtml $+ > $@

3.5.4 Temporary directories

Sometimes, your project may include temporary directories–directories where you place intermediate results. these directories are deleted whenever the project is cleanup up. This means, in particular, that you can’t place an OMakefile in a temporary directory, because it will be removed when the directory is removed.

Instead, if you need to define a configuration for any of these directories, you will need to define it using a .SUBDIRS body.

    section
        CREATE_SUBDIRS = true
        .SUBDIRS: tmp
            # Compute an MD5 digest
            %.digest: %.comments
               echo $(digest $<) > $@
            # Extract comments from the source files

            %.comments: ../src/%.src
               grep '^#' $< > $@
            .DEFAULT: foo.digest
    .PHONY: clean
    clean:
        rm -rf tmp

In this example, we define the CREATE_SUBDIRS variable as true, so that the tmp directory will be created if it does not exist. The .SUBDIRS body in this example is a bit contrived, but it illustrates the kind of specification you might expect. The clean phony-target indicates that the tmp directory should be removed when the project is cleaned up.

46 CHAPTER 3. ADDITIONAL BUILD EXAMPLES

OMake concepts and syntax

Projects are specified to omake with OMakefiles. The OMakefile has a format similar to a Makefile. An OMakefile has three main kinds of syntactic objects: variable definitions, function definitions, and rule definitions.

4.1 Variables

Variables are defined with the following syntax. The name is any sequence of alphanumeric characters, underscore _, and hyphen -.

   <name> = <value>

Values are defined as a sequence of literal characters and variable expansions. A variable expansion has the form $(<name>), which represents the value of the <name> variable in the current environment. Some examples are shown below.

   CC = gcc
   CFLAGS = -Wall -g
   COMMAND = $(CC) $(CFLAGS) -O2

In this example, the value of the COMMAND variable is the string gcc -Wall -g -O2. Unlike make(1), variable expansion is eager and functional (see also the

section on Scoping). That is, variable values are expanded immediately and new variable definitions do not affect old ones. For example, suppose we extend the previous example with following variable definitions.

   X = $(COMMAND)
   COMMAND = $(COMMAND) -O3
   Y = $(COMMAND)

In this example, the value of the X variable is the string gcc -Wall -g -O2 as before, and the value of the Y variable is gcc -Wall -g -O2 -O3.

4.2 Adding to a variable definition

Variables definitions may also use the += operator, which adds the new text to an existing definition. The following two definitions are equivalent.

   # Add options to the CFLAGS variable
   CFLAGS = $(CFLAGS) -Wall -g
   # The following definition is equivalent
   CFLAGS += -Wall -g

4.3 Arrays

Arrays can be defined by appending the [] sequence to the variable name and defining initial values for the elements as separate lines. Whitespace is significant on each line. The following code sequence prints c d e.

    X[] =
        a b
        c d e
        f
    println($(nth 2, $(X)))

4.4 Special characters and quoting

The following characters are special to omake: $():,=#\. To treat any of these characters as normal text, they should be escaped with the backslash character

\.
    DOLLAR = \$

Newlines may also be escaped with a backslash to concatenate several lines.

    FILES = a.c\
            b.c\
            c.c

Note that the backslash is not an escape for any other character, so the following works as expected (that is, it preserves the backslashes in the string).

    DOSTARGET = C:\WINDOWS\control.ini

An alternative mechanism for quoting special text is the use $"..." escapes. The number of double-quotations is arbitrary. The outermost quotations are not included in the text.

    A = $""String containing "quoted text" ""
    B = $"""Multi-line
        text.
        The # character is not special"""

4.5 Function definitions

Functions are defined using the following syntax.

   <name>(<params>) =
      <indented-body>

The parameters are a comma-separated list of identifiers, and the body must be placed on a separate set of lines that are indented from the function definition itself. For example, the following text defines a function that concatenates its arguments, separating them with a colon.

    ColonFun(a, b) =
        return($(a):$(b))

The return expression can be used to return a value from the function. A return statement is not required; if it is omitted, the returned value is the value of the last expression in the body to be evaluated. NOTE: as of version 0.9.6, return is a control operation, causing the function to immediately return. In the following example, when the argument a is true, the function f immediately returns the value 1 without evaluating the print statement.

    f(a) =
       if $(a)
          return 1
       println(The argument is false)
       return 0

In many cases, you may wish to return a value from a section or code block without returning from the function. In this case, you would use the value operator. In fact, the value operator is not limited to functions, it can be used any place where a value is required. In the following definition, the variable X is defined as 1 or 2, depending on the value of a, then result is printed, and returned from the function.

    f_value(a) =
       X =
          if $(a)
             value 1
          else
             value 2
       println(The value of X is $(X))
       value $(X)

Functions are called using the GNU-make syntax, $(<name> <args)), where <args> is a comma-separated list of values. For example, in the following program, the variable X contains the value foo:bar.

   X = $(ColonFun foo, bar)

If the value of a function is not needed, the function may also be called using standard function call notation. For example, the following program prints the string “She says: Hello world”.

    Printer(name) =
        println($(name) says: Hello world)
    Printer(She)

4.6 Comments

Comments begin with the # character and continue to the end of the line.

4.7 File inclusion

Files may be included with the include or open form. The included file must use the same syntax as an OMakefile.

    include $(Config_file)

The open operation is similar to an include, but the file is included at most once.

    open Config
    # Repeated opens are ignored, so this
    # line has no effect.
    open Config

If the file specified is not an absolute filenmame, both include and open operations search for the file based on the OMAKEPATH variable. In case of the open directive, the search is performed at parse time, and the argument to open may not contain any expressions.

4.8 Scoping, sections

Scopes in omake are defined by indentation level. When indentation is increased, such as in the body of a function, a new scope is introduced.

The section form can also be used to define a new scope. For example, the following code prints the line X = 2, followed by the line X = 1.

    X = 1
    section
        X = 2
        println(X = $(X))
    println(X = $(X))

This result may seem surprising–the variable definition within the section is not visible outside the scope of the section.

The export form, which will be described in detail in Section 5.3, can be used to circumvent this restriction by exporting variable values from an inner scope. It must be the final expression in a scope. For example, if we modify the previous example by adding an export expression, the new value for the X variable is retained, and the code prints the line X = 2 twice.

    X = 1
    section
        X = 2
        println(X = $(X))
        export
    println(X = $(X))

There are also cases where separate scoping is quite important. For example, each OMakefile is evaluated in its own scope. Since each part of a project may have its own configuration, it is important that variable definitions in one OMakefile do not affect the definitions in another.

To give another example, in some cases it is convenient to specify a separate set of variables for different build targets. A frequent idiom in this case is to use the section command to define a separate scope.

   section
      CFLAGS += -g
      %.c: %.y
          $(YACC) $<
      .SUBDIRS: foo
   .SUBDIRS: bar baz

In this example, the -g option is added to the CFLAGS variable by the foo subdirectory, but not by the bar and baz directories. The implicit rules are scoped as well and in this example, the newly added yacc rule will be inherited by the foo subdirectory, but not by the bar and baz ones; furthermore this implicit rule will not be in scope in the current directory.

4.9 Conditionals

Top level conditionals have the following form.

if <test>
<true-clause>
elseif <text>
<elseif-clause>
else
<else-clause>

The <test> expression is evaluated, and if it evaluates to a true value (see Section 8.2 for more information on logical values, and Boolean functions), the code for the <true-clause> is evaluated; otherwise the remaining clauses are evaluated. There may be multiple elseif clauses; both the elseif and else clauses are optional. Note that the clauses are indented, so they introduce new scopes.

When viewed as a predicate, a value corresponds to the Boolean false, if its string representation is the empty string, or one of the strings false, no, nil, undefined, or 0. All other values are true.

The following example illustrates a typical use of a conditional. The OSTYPE variable is the current machine architecture.

# Common suffixes for files
if $(equal $(OSTYPE), Win32)
EXT_LIB = .lib
EXT_OBJ = .obj
EXT_ASM = .asm
EXE = .exe
export
elseif $(mem $(OSTYPE), Unix Cygwin)
EXT_LIB = .a
EXT_OBJ = .o
EXT_ASM = .s
EXE =
export
else
# Abort on other architectures
eprintln(OS type $(OSTYPE) is not recognized)
exit(1)

4.10 Matching

Pattern matching is performed with the switch and match forms.

switch <string>

case <pattern1>
    <clause1>
case <pattern2>
    <clause2>
...
default
   <default-clause>

The number of cases is arbitrary. The default clause is optional; however, if it is used it should be the last clause in the pattern match.

For switch, the string is compared with the patterns literally.

switch $(HOST)
case mymachine
    println(Building on mymachine)
default
    println(Building on some other machine)

Patterns need not be constant strings. The following function tests for a literal match against pattern1, and a match against pattern2 with ## delimiters.

Switch2(s, pattern1, pattern2) =
  switch $(s)
  case $(pattern1)
      println(Pattern1)
  case $"##$(pattern2)##"
      println(Pattern2)
  default
      println(Neither pattern matched)

For match the patterns are egrep(1)-style regular expressions. The numeric variables $1, $2, ... can be used to retrieve values that are matched by \(...\) expressions.

match $(NODENAME)@$(SYSNAME)@$(RELEASE)
case $"mymachine.*@\(.*\)@\(.*\)"
    println(Compiling on mymachine; sysname $1 and release $2 are ignored)
case $".*@Linux@.*2\.4\.\(.*\)"
    println(Compiling on a Linux 2.4 system; subrelease is $1)
default
    eprintln(Machine configuration not implemented)
    exit(1)

4.11 Objects

OMake is an object-oriented language. Generally speaking, an object is a value that contains fields and methods. An object is defined with a . suffix for a variable. For example, the following object might be used to specify a point (1, 5) on the two-dimensional plane.

Coord. =
x = 1
y = 5
print(message) =
   println($"$(message): the point is ($(x), $(y)")
# Define X to be 5
X = $(Coord.x)
# This prints the string, "Hi: the point is (1, 5)"
Coord.print(Hi)

The fields x and y represent the coordinates of the point. The method print prints out the position of the point.

4.12 Classes

We can also define classes. For example, suppose we wish to define a generic Point class with some methods to create, move, and print a point. A class is really just an object with a name, defined with the class directive.

Point. =
class Point
# Default values for the fields
x = 0
y = 0
# Create a new point from the coordinates
new(x, y) =
   this.x = $(x)
   this.y = $(y)
   return $(this)
# Move the point to the right
move-right() =
   x = $(add $(x), 1)
   return $(this)

    # Print the point
    print() =
       println($"The point is ($(x), $(y)")
p1 = $(Point.new 1, 5)
p2 = $(p1.move-right)
# Prints "The point is (1, 5)"
p1.print()
# Prints "The point is (2, 5)"
p2.print()

Note that the variable $(this) is used to refer to the current object. Also, classes and objects are functional—the new and move-right methods return new objects. In this example, the object p2 is a different object from p1, which retains the original (1, 5) coordinates.

4.13 Inheritance

Classes and objects support inheritance (including multiple inheritance) with the extends directive. The following definition of Point3D defines a point with x, y, and z fields. The new object inherits all of the methods and fields of the parent classes/objects.

Z. =
   z = 0
Point3D. =
   extends $(Point)
   extends $(Z)
   class Point3D
   print() =
      println($"The 3D point is ($(x), $(y), $(z))")
# The "new" method was not redefined, so this
# defines a new point (1, 5, 0).
p = $(Point3D.new 1, 5)

4.14 Special objects/sections

Objects provide one way to manage the OMake namespace. There are also four special objects that are further used to control the namespace.

4.15 private.

The private. section is used to define variables that are private to the current file/scope. The values are not accessible outside the scope. Variables defined in a private. object can be accessed only from within the section where they are defined.

 Obj. =
    private. =
       X = 1
    print() =
       println(The value of X is: $(X))
 # Prints:
 #    The private value of X is: 1
 Obj.print()
 # This is an error--X is private in Obj
 y = $(Obj.X)

In addition, private definitions do not affect the global value of a variable.

# The public value of x is 1
x = 1
f() =
    println(The public value of x is: $(x))
# This object uses a private value of x
Obj. =
    private. =
       x = 2
    print() =
       x = 3
       println(The private value of x is: $(x))
       f()
# Prints:
#    The private value of x is: 3
#    The public value of x is: 1
Obj.print()

Private variables have two additional properties.

1. Private variables are local to the file in which they are defined.

2. Private variables are not exported by the export directive, unless they are mentioned explicitly.

        private. =
           FLAG = true
        section
           FLAG = false
           export
        # FLAG is still true
        section
           FLAG = false
           export FLAG
        # FLAG is now false

4.16 protected.

The protected. object is used to define fields that are local to an object. They can be accessed as fields, but they are not passed dynamically to other functions. The purpose of a protected variable is to prevent a variable definition within the object from affecting other parts of the project.

X = 1
f() =
   println(The public value of X is: $(X))
# Prints:
#    The public value of X is: 2
section
   X = 2
   f()
# X is a protected field in the object
Obj. =
   protected. =
      X = 3
   print() =
      println(The protected value of X is: $(X))
      f()
# Prints:

#    The protected value of X is: 3
#    The public value of X is: 1
Obj.print()
# This is legal, it defines Y as 3
Y = $(Obj.X)

In general, it is a good idea to define object variables as protected. The resulting code is more modular because variables in your object will not produce unexpected clashes with variables defined in other parts of the project.

4.17 public.

The public. object is used to specify public dynamically-scoped variables. In the following example, the public. object specifies that the value X = 4 is to be dynamically scoped. Public variables are not defined as fields of an object.

X = 1
f() =
   println(The public value of X is: $(X))
# Prints:
#    The public value of X is: 2
section
   X = 2
   f()
Obj. =
   protected. =
      X = 3
   print() =
      println(The protected value of X is: $(X))
      public. =
         X = 4
      f()
# Prints:
#    The protected value of X is: 3
#    The public value of X is: 4
Obj.print()

4.18 static.

The static. object is used to specify values that are persistent across runs of OMake. They are frequently used for configuring a project. Configuring a project can be expensive, so the static. object ensure that the configuration is performed just once. In the following (somewhat trivial) example, a static section is used to determine if the L^{A}TEX command is available. The $(where latex) function returns the full pathname for latex, or false if the command is not found.

   static. =
      LATEX_ENABLED = false
      print(--- Determining if LaTeX is installed )
      if $(where latex)
          LATEX_ENABLED = true
          export
      if $(LATEX_ENABLED)
         println($'(enabled)')
      else
         println($'(disabled)')

The OMake standard library provides a number of useful functions for programming the static. tests, as described in Chapter 13. Using the standard library, the above can be rewritten as

   open configure/Configure
   static. =
      LATEX_ENABLED = $(CheckProg latex)

As a matter of style, a static. section that is used for configuration should print what it is doing using the ConfMsgChecking and ConfMsgResult functions (of couse, most of helper functions in the standard library would do that automatically).

4.19 Short syntax for scoping objects

The usual dot-notation can be used for private, protected, and public variables (but not static variables).

    # Public definition of X
    public.X = 1
    # Private definition of X
    private.X = 2
    # Prints:
    #    The public value of X is: 1
    #    The private value of X is: 2
    println(The public value of X is: $(public.X))
    println(The private value of X is: $(private.X))

4.20 Modular programming

The scoping objects help provide a form of modularity. When you write a new file or program, explicit scoping declarations can be used to define an explicit interface for your code, and help avoid name clashes with other parts of the project. Variable definitions are public by default, but you can control this with private definitions.

    # These variables are private to this file
    private. =
       FILES = foo1 foo2 foo3
       SUFFIX = .o
       OFILES = $(addsuffix $(SUFFIX), $(FILES))
    # These variables are public
    public. =
       CFLAGS += -g
    # Build the files with the -g option
    $(OFILES):

Expressions and values

omake provides a full programming-language including many system and IO functions. The language is object-oriented – everything is an object, including the base values like numbers and strings. However, the omake language differs from other scripting languages in three main respects.

• Scoping is dynamic.

• Apart from IO, the language is entirely functional – there is no assignment operator in the language.

• Evaluation is normally eager – that is, expressions are evaluated as soon as they are encountered.

To illustrate these features, we will use the osh(1) omake program shell. The osh(1) program provides a toploop, where expressions can be entered and the result printed. osh(1) normally interprets input as command text to be executed by the shell, so in many cases we will use the value form to evaluate an expression directly.

    osh> 1
    *** omake error: File -: line 1, characters 0-1 command not found: 1
    osh> value 1
    - : "1" : Sequence
    osh> ls -l omake
    -rwxrwxr-x  1 jyh jyh 1662189 Aug 25 10:24 omake*

5.1 Dynamic scoping

Dynamic scoping means that the value of a variable is determined by the most recent binding of the variable in scope at runtime. Consider the following program.

    OPTIONS = a b c
    f() =
       println(OPTIONS = $(OPTIONS))
    g() =
       OPTIONS = d e f
       f()

If f() is called without redefining the OPTIONS variable, the function should print the string OPTIONS = a b c.

In contrast, the function g() redefines the OPTIONS variable and evaluates f() in that scope, which now prints the string OPTIONS = d e f.

The body of g defines a local scope – the redefinition of the OPTIONS variable is local to g and does not persist after the function terminates.

    osh> g()
    OPTIONS = d e f
    osh> f()
    OPTIONS = a b c

Dynamic scoping can be tremendously helpful for simplifying the code in a project. For example, the OMakeroot file defines a set of functions and rules for building projects using such variables as CC, CFLAGS, etc. However, different parts of a project may need different values for these variables. For example, we may have a subdirectory called opt where we want to use the -03 option, and a subdirectory called debug where we want to use the -g option. Dynamic scoping allows us to redefine these variables in the parts of the project without having to redefine the functions that use them.

    section
       CFLAGS = -O3
       .SUBDIRS: opt
    section
       CFLAGS = -g
       .SUBDIRS: debug

However, dynamic scoping also has drawbacks. First, it can become confusing: you might have a variable that is intended to be private, but it is accidentally redefined elsewhere. For example, you might have the following code to construct search paths.

   PATHSEP = :
   make-path(dirs) =
      return $(concat $(PATHSEP), $(dirs))
   make-path(/bin /usr/bin /usr/X11R6/bin)
   - : "/bin:/usr/bin:/usr/X11R6/bin" : String

However, elsewhere in the project, the PATHSEP variable is redefined as a directory separator /, and your function suddenly returns the string /bin//usr/bin//usr/X11R6/bin, obviously not what you want.

The private block is used to solve this problem. Variables that are defined in a private block use static scoping – that is, the value of the variable is determined by the most recent definition in scope in the source text.

   private
      PATHSEP = :
   make-path(dirs) =
      return $(concat $(PATHSEP), $(dirs))
   PATHSEP = /
   make-path(/bin /usr/bin /usr/X11R6/bin)
   - : "/bin:/usr/bin:/usr/X11R6/bin" : String

5.2 Functional evaluation

Apart from I/O, omake programs are entirely functional. This has two parts:

• There is no assignment operator.

• Functions are values, and may be passed as arguments, and returned from functions just like any other value.

The second item is straightforward. For example, the following program defines an increment function by returning a function value.

   incby(n) =
      g(i) =
         return $(add $(i), $(n))
      return $(g)
   f = $(incby 5)
   value $(f 3)
   - : 8 : Int

The first item may be the most confusing initially. Without assignment, how is it possible for a subproject to modify the global behavior of the project? In fact, the omission is intentional. Build scripts are much easier to write when there is a guarantee that subprojects do not interfere with one another.

However, there are times when a subproject needs to propagate information back to its parent object, or when an inner scope needs to propagate information back to the outer scope.

5.3 Exporting the environment

The export directive can be used to propagate all or part of an inner scope back to its parent. The export directive should be the last statement in a block. If used without arguments, the entire scope is propagated back to the parent; otherwise the arguments specify which part of the environment to propagate. The most common usage is to export the definitions in a conditional block. In the following example, the variable B is bound to 2 after the conditional. The A variable is not redefined.

    if $(test)
       A = 1
       B = $(add $(A), 1)
       export B
    else
       B = 2
       export

If the export directive is used without an argument, all of the following is exported:

• The values of all the dynamically scoped variables (as described in Section 4.17).

• The current working directory.

• The current Unix environment.

• The current implicit rules and implicit dependencies (see also Section 7.11.1).

• The current set of “phony” target declarations (see Sections 7.10 and 7.11.3).

If the export directive is used with an argument, the argument expression is evaluated and the resulting value is interpreted as follows:

• If the value is empty, everything is exported, as described above.

• If the value represents a environment (or a partial environment) captured using the export function, then the corresponding environment or partial environment is exported.

• Otherwise, the value must be a sequence of strings specifying which items are to be propagated back. The following strings have special meaning:

– .RULE — implicit rules and implicit dependencies.

– .PHONY — the set of “phony” target declarations.

All other strings are interpreted as names of the variables that need to be propagated back.

For example, in the following (somewhat artificial) example, the variables A and B will be exported, and the implicit rule will remain in the environment after the section ends, but the variable TMP and the target tmp_phony will remain unchanged.

WANT_TO_EXPORT[] = A B .RULE
section
   A = 1
   B = 2
   TMP = $(add $(A), $(B))
   .PHONY: tmp_phony
   tmp_phony:
      prepare_foo
   %.foo: %.bar tmp_phony
      compute_foo $(TMP) $< $@
   export $(WANT_TO_EXPORT)

5.4 Eager evaluation

Evaluation in omake is eager. That is, expressions are evaluated as soon as they are encountered by the evaluator. One effect of this is that the right-hand-side of a variable definition is expanded when the variable is defined.

    osh> A = 1
    - : "1"
    osh> A = $(A)$(A)
    - : "11"

In the second definition, A = $(A)$(A), the right-hand-side is evaluated first, producing the sequence 11. Then the variable A is redefined as the new value. When combined with dynamic scoping, this has many of the same properties as conventional imperative programming.

    osh> A = 1
    - : "1"
    osh> printA() =
        println($"A = $A")
    osh> A = $(A)$(A)
    - : "11"
    osh> printA()
    11

In this example, the print function is defined in the scope of A. When it is called on the last line, the dynamic value of A is 11, which is what is printed.

However, dynamic scoping and imperative programming should not be confused. The following example illustrates a difference. The second printA is not in the scope of the definition A = x$(A)$(A)x, so it prints the original value, 1.

    osh> A = 1
    - : "1"
    osh> printA() =
        println($"A = $A")
    osh> section
             A = x$(A)$(A)x
             printA()
    x11x
    osh> printA()
    1

See also Section 6.5 for further ways to control the evaluation order through the use of “lazy” expressions.

5.5 Objects

omake is an object-oriented language. Everything is an object, including base values like numbers and strings. In many projects, this may not be so apparent because most evaluation occurs in the default toplevel object, the Pervasives object, and few other objects are ever defined.

However, objects provide additional means for data structuring, and in some cases judicious use of objects may simplify your project.

Objects are defined with the following syntax. This defines name to be an object with several methods an values.

    name. =                     # += may be used as well
       extends parent-object    # optional
       class class-name         # optional
       # Fields
       X = value
       Y = value
       # Methods
       f(args) =
          body
       g(arg) =
          body

An extends directive specifies that this object inherits from the specified parent-object. The object may have any number of extends directives. If there is more than on extends directive, then fields and methods are inherited from all parent objects. If there are name conflicts, the later definitions override the earlier definitions.

The class directive is optional. If specified, it defines a name for the object that can be used in instanceof operations, as well as :: scoping directives discussed below.

The body of the object is actually an arbitrary program. The variables defined in the body of the object become its fields, and the functions defined in the body become its methods.

5.6 Field and method calls

The fields and methods of an object are named using object.name notation. For example, let’s define a one-dimensional point value.

   Point. =
      class Point
      # Default value
      x = $(int 0)
      # Create a new point
      new(x) =
         x = $(int $(x))
         return $(this)
      # Move by one
      move() =
         x = $(add $(x), 1)
         return $(this)
   osh> p1 = $(Point.new 15)
   osh> value $(p1.x)
   - : 15 : Int
   osh> p2 = $(p1.move)
   osh> value $(p2.x)
   - : 16 : Int

The $(this) variable always represents the current object. The expression $(p1.x) fetches the value of the x field in the p1 object. The expression $(Point.new 15) represents a method call to the new method of the Point object, which returns a new object with 15 as its initial value. The expression $(p1.move) is also a method call, which returns a new object at position 16.

Note that objects are functional — it is not possible to modify the fields or methods of an existing object in place. Thus, the new and move methods return new objects.

5.7 Method override

Suppose we wish to create a new object that moves by 2 units, instead of just 1. We can do it by overriding the move method.

   Point2. =
      extends $(Point)
      # Override the move method
      move() =
         x = $(add $(x), 2)
         return $(this)
   osh> p2 = $(Point2.new 15)
   osh> p3 = $(p2.move)
   osh> value $(p3.x)
   - : 17 : Int

However, by doing this, we have completely replaced the old move method.

5.8 Super calls

Suppose we wish to define a new move method that just calls the old one twice. We can refer to the old definition of move using a super call, which uses the

notation $(classname::name <args>). The classname should be the name of

the superclass, and name the field or method to be referenced. An alternative way of defining the Point2 object is then as follows.

   Point2. =
      extends $(Point)
      # Call the old method twice
      move() =
         this = $(Point::move)
         return $(Point::move)

Note that the first call to $(Point::move) redefines the current object (the this variable). This is because the method returns a new object, which is re-used for the second call.

Additional language examples

In this section, we’ll explore the core language through a series of examples (examples of the build system are the topic of the Chapter 3).

For most of these examples, we’ll use the osh command interpreter. For simplicity, the values printed by osh have been abbreviated.

6.1 Strings and arrays

The basic OMake values are strings, sequences, and arrays of values. Sequences are like arrays of values separated by whitespace; the sequences are split on demand by functions that expect arrays.

   osh> X = 1 2
   - : "1 2" : Sequence
   osh> addsuffix(.c, $X)
   - : <array 1.c 2.c> : Array

Sometimes you want to define an array explicitly. For this, use the [] brackets after the variable name, and list each array entry on a single indented line.

   osh> A[] =
           Hello world
           $(getenv HOME)
   - : <array "Hello world" "/home/jyh"> : Array

One central property of arrays is that whitespace in the elements is significant. This can be useful, especially for filenames that contain whitespace.

   # List the current files in the directory
    osh> ls -Q

"fee"  "fi"  "foo"  "fum"
osh> NAME[] =
        Hello world
- : <array "Hello world"> : Array
osh> touch $(NAME)
osh> ls -Q
"fee"  "fi"  "foo"  "fum"  "Hello world"

6.2 Quoted strings

A String is a single value; whitespace is significant in a string. Strings are introduced with quotes. There are four kinds of quoted elements; the kind is determined by the opening quote. The symbols ' (single-quote) and " (doublequote) introduce the normal shell-style quoted elements. The quotation symbols are included in the result string. Variables are always expanded within a quote of this kind. Note that the osh(1) (Chapter 14) printer escapes double-quotes within the string; these are only for printing, they are not part of the string itself.

osh> A = 'Hello "world"'
- : "'Hello \"world\"'" : String
osh> B = "$(A)"
- : "\"'Hello \"world\"'\"" : String
osh> C = 'Hello \'world\''
- : "'Hello 'world''" : String

A second kind of quote is introduced with the $' and $" quotes. The number of opening and closing quote symbols is arbitrary. These quotations have several properties:

• The quote delimiters are not part of the string.

• Backslash \ symbols within the string are treated as normal characters.

• The strings may span several lines.

• Variables are expanded within $" sequences, but not within $' sequences.

osh> A = $'''Here $(IS) an '''' \(example\) string['''
- : "Here $(IS) an '''' \\(example\\) string[" : String
osh> B = $""""A is "$(A)" """"
- : "A is \"Here $(IS) an '''' \\(example\\) string[\" " : String
osh> value $(A.length)
- : 38 : Int
osh> value $(A.nth 5)
- : "$" : String
osh> value $(A.rev)
- : "[gnirts )\\elpmaxe(\\ '''' na )SI($ ereH" : String

Strings and sequences both have the property that they can be merged with adjacent non-whitespace text.

osh> A = a b c
- : "a b c" : Sequence
osh> B = $(A).c
- : <sequence "a b c" : Sequence ".c" : Sequence> : Sequence
osh> value $(nth 2, $(B))
- : "c.c" : String
osh> value $(length $(B))
- : 3 : Int

Arrays are different. The elements of an array are never merged with adjacent text of any kind. Arrays are defined by adding square brackets [] after a variable name and defining the elements with an indented body. The elements may include whitespace.

osh> A[] =
        a b
        foo bar
- : <array
       "a b" : Sequence
       "foo bar" : Sequence>
       : Array
osh> echo $(A).c
a b foo bar .c
osh> value $(A.length)
- : 2 : Int
osh> value $(A.nth 1)
- : "foo bar" : Sequence

Arrays are quite helpful on systems where filenames often contain whitespace.

osh> FILES[] =
         c:\Documents and Settings\jyh\one file
         c:\Program Files\omake\second file
osh> CFILES = $(addsuffix .c, $(FILES))
osh> echo $(CFILES)
c:\Documents and Settings\jyh\one file.c c:\Program Files\omake\second file.c

6.3 Files and directories

OMake projects usually span multiple directories, and different parts of the project execute commands in different directories. There is a need to define a location-independent name for a file or directory.

This is done with the $(file <names>) and $(dir <names>) functions.

   osh> mkdir tmp
   osh> F = $(file fee)
   osh> section:
            cd tmp
            echo $F
   ../fee
   osh> echo $F
   fee

Note the use of a section: to limit the scope of the cd command. The section temporarily changes to the tmp directory where the name of the file is ../fee. Once the section completes, we are still in the current directory, where the name of the file is fee.

One common way to use the file functions is to define proper file names in your project OMakefile, so that references within the various parts of the project will refer to the same file.

    osh> cat OMakefile
    ROOT = $(dir .)
    TMP  = $(dir tmp)
    BIN  = $(dir bin)
    ...

6.4 Iteration, mapping, and foreach

Most builtin functions operate transparently on arrays.

    osh> addprefix(-D, DEBUG WIN32)
    - : -DDEBUG -DWIN32 : Array
    osh> mapprefix(-I, /etc /tmp)
    - : -I /etc -I /tmp : Array
    osh> uppercase(fee fi foo fum)
    - : FEE FI FOO FUM : Array

The mapprefix and addprefix functions are slightly different (the addsuffix and mapsuffix functions are similar). The addprefix adds the prefex to each array element. The mapprefix doubles the length of the array, adding the prefix as a new array element before each of the original elements.

Even though most functions work on arrays, there are times when you will want to do it yourself. The foreach function is the way to go. The foreach function has two forms, but the form with a body is most useful. In this form, the function takes two arguments and a body. The second argument is an array, and the first is a variable. The body is evaluated once for each element of the array, where the variable is bound to the element. Let’s define a function to add 1 to each element of an array of numbers.

   osh> add1(l) =
            foreach(i, $l):
                add($i, 1)
   osh> add1(7 21 75)
   - : 8 22 76 : Array

Sometimes you have an array of filenames, and you want to define a rule for each of them. Rules are not special, you can define them anywhere a statement is expected. Say we want to write a function that describes how to process each file, placing the result in the tmp/ directory.

   TMP = $(dir tmp)
   my-special-rule(files) =
      foreach(name, $(files))
         $(TMP)/$(name): $(name)
            process $< > $@

Later, in some other part of the project, we may decide that we want to use this function to process some files.

   # These are the files to process in src/lib
   MY_SPECIAL_FILES[] =
       fee.src
       fi.src
       file with spaces in its name.src
   my-special-rule($(MY_SPECIAL_FILES))

The result of calling my-special-rule is exactly the same as if we had written the following three rules explicitly.

    $(TMP)/fee.src: fee.src
        process fee > $@
    $(TMP)/fi.src: fi.src
        process fi.src > $@
    $(TMP)/$"file with spaces in its name.src": $"file with spaces in its name.src"
        process $< > $@

Of course, writing these rules is not nearly as pleasant as calling the function. The usual properties of function abstraction give us the usual benefits. The code is less redundant, and there is a single location (the my-special-rule function) that defines the build rule. Later, if we want to modify/update the rule, we need do so in only one location.

6.5 Lazy expressions

Evaluation in omake is normally eager. That is, expressions are evaluated as soon as they are encountered by the evaluator. One effect of this is that the right-hand-side of a variable definition is expanded when the variable is defined.

There are two ways to control this behavior. The $`(v) form introduces lazy behavior, and the $,(v) form restores eager behavior. Consider the following sequence.

    osh> A = 1
    - : "1" : Sequence
    osh> B = 2
    - : "2" : Sequence
    osh> C = $`(add $(A), $,(B))
    - : $(apply add $(apply A) "2" : Sequence)
    osh> println(C = $(C))
    C = 3
    osh> A = 5
    - : "5" : Sequence
    osh> B = 6
    - : "6" : Sequence
    osh> println(C = $(C))
    C = 7

The definition C = $`(add $(A), $,(B)) defines a lazy application. The add function is not applied in this case until its value is needed. Within this expression, the value $,(B) specifies that B is to be evaluated immediately, even though it is defined in a lazy expression.

The first time that we print the value of C, it evaluates to 3 since A is 1 and B is 2. The second time we evaluate C, it evaluates to 7 because A has been redefined to 5. The second definition of B has no effect, since it was evaluated at definition time.

6.5.1 A larger example of lazy expressions

Lazy expressions are not evaluated until their result is needed. Some people, including this author, frown on overuse of lazy expressions, mainly because it is difficult to know when evaluation actually happens. However, there are cases where they pay off.

One example comes from option processing. Consider the specification of “include” directories on the command line for a C compiler. If we want to include files from /home/jyh/include and ../foo, we specify it on the command

line with the options -I/home/jyh/include -I../foo.

Suppose we want to define a generic rule for building C files. We could define a INCLUDES array to specify the directories to be included, and then define a generic implicit rule in our root OMakefile.

    # Generic way to compile C files.
    CFLAGS = -g
    INCLUDES[] =
    %.o: %.c
       $(CC) $(CFLAGS) $(INCLUDES) -c $<

# The src directory builds my_widget+ from 4 source files.
# It reads include files from the include directory.
.SUBDIRS: src
    FILES = fee fi foo fum
    OFILES = $(addsuffix .o, $(FILES))
    INCLUDES[] += -I../include
    my_widget: $(OFILES)
       $(CC) $(CFLAGS) -o $@ $(OFILES)

But this is not quite right. The problem is that INCLUDES is an array of options, not directories. If we later wanted to recover the directories, we would have to strip the leading -I prefix, which is a hassle. Furthermore, we aren’t using proper names for the directories. The solution here is to use a lazy expression. We’ll define INCLUDES as a directory array, and a new

variable PREFIXED_INCLUDES that adds the -I prefix. The PREFIXED_INCLUDES

is computed lazily, ensuring that the value uses the most recent value of the INCLUDES variable.

# Generic way to compile C files.
CFLAGS = -g
INCLUDES[] =
PREFIXED_INCLUDES[] = $`(addprefix -I, $(INCLUDES))
%.o: %.c
   $(CC) $(CFLAGS) $(PREFIXED_INCLUDES) -c $<
# For this example, we define a proper name for the include directory
STDINCLUDE = $(dir include)
# The src directory builds my_widget+ from 4 source files.
# It reads include files from the include directory.
.SUBDIRS: src
    FILES = fee fi foo fum
    OFILES = $(addsuffix .o, $(FILES))
    INCLUDES[] += $(STDINCLUDE)
    my_widget: $(OFILES)
       $(CC) $(CFLAGS) -o $@ $(OFILES)

Note that there is a close connection between lazy values and functions. In the example above, we could equivalently define PREFIXED_INCLUDES as a function with zero arguments.

PREFIXED_INCLUDES() =
    addprefix(-I, $(INCLUDES))

6.6 Scoping and exports

The OMake language is functional (apart from IO and shell commands). This comes in two parts: functions are first-class, and variables are immutable (there is no assignment operator). The latter property may seem strange to users used to GNU make, but it is actually a central point of OMake. Since variables can’t be modified, it is impossible (or at least hard) for one part of the project to interfere with another.

To be sure, pure functional programming can be awkward. In OMake, each new indentation level introduces a new scope, and new definitions in that scope are lost when the scope ends. If OMake were overly strict about scoping, we would wind up with a lot of convoluted code.

osh> X = 1
osh> setenv(BOO, 12)
osh> if $(equal $(OSTYPE), Win32)
         setenv(BOO, 17)
         X = 2
osh> println($X $(getenv BOO))
1 12

The export command presents a way out. It takes care of “exporting” a value (or the entire variable environment) from an inner scope to an outer one.

osh> X = 1
osh> setenv(BOO, 12)
osh> if $(equal $(OSTYPE), Win32)
         setenv(BOO, 17)
         X = 2
         export
osh> println($X $(getenv BOO))
2 17

Exports are especially useful in loop to export values from one iteration of a loop to the next.

# Ok, let's try to add up the elements of the array
osh>sum(l) =
        total = 0
        foreach(i, $l)
            total = $(add $(total), $i)
        value $(total)
osh>sum(1 2 3)
- : 0 : Int
# Oops, that didn't work!
osh>sum(l) =

       total = 0
       foreach(i, $l)
           total = $(add $(total), $i)
           export
       value $(total)
osh>sum(1 2 3)
- : 6 : Int

A while loop is another form of loop, with an auto-export.

osh>i = 0
osh>total = 0
osh>while $(lt $i, 10)
        total = $(add $(total), $i)
        i = $(add $i, 1)
osh>println($(total))
45

6.7 Shell aliases

Sometimes you may want to define an alias, an OMake command that masquerades as a real shell command. You can do this by adding your function as a method to the Shell object.

For an example, suppose we use the awk function to print out all the comments in a file.

osh>cat comment.om
# Comment function
comments(filename) =
    awk($(filename))
    case $'^#'
        println($0)
# File finished
osh>include comment
osh>comments(comment.om)
# Comment function
# File finished

To add it as an alias, add the method (using += to preserve the existing entries in the Shell).

osh>Shell. +=
       printcom(argv) =
           comments($(nth 0, $(argv)))
osh>printcom comment.om > output.txt
osh>cat output.txt

# Comment function
# File finished

A shell command is passed an array of arguments argv. This does not include the name of the alias.

6.8 Input/output redirection on the cheap

As it turns out, scoping also provides a nice alternate way to perform redirection. Suppose you have already written a lot of code that prints to the standard output channel, but now you decide you want to redirect it. One way to do it is using the technique in the previous example: define your function as an alias, and then use shell redirection to place the output where you want.

There is an alternate method that is easier in some cases. The variables stdin, stdout, and stderr define the standard I/O channels. To redirect output, redefine these variables as you see fit. Of course, you would normally do this in a nested scope, so that the outer channels are not affected.

osh>f() =
        println(Hello world)
osh>f()
Hello world
osh>section:
        stdout = $(fopen output.txt, w)
        f()
        close($(stdout))
osh>cat output.txt
Hello world

This also works for shell commands. If you like to gamble, you can try the following example.

osh>f() =
        println(Hello world)
osh>f()
Hello world
osh>section:
        stdout = $(fopen output.txt, w)
        f()
        cat output.txt
        close($(stdout))
osh>cat output.txt
Hello world
Hello world

Rules

Rules are used by OMake to specify how to build files. At its simplest, a rule has the following form.

    <target>: <dependencies>
        <commands>

The <target> is the name of a file to be built. The <dependencies> are a list of files that are needed before the <target> can be built. The <commands> are a list of indented lines specifying commands to build the target. For example, the following rule specifies how to compile a file hello.c.

    hello.o: hello.c
        $(CC) $(CFLAGS) -c -o hello.o hello.c

This rule states that the hello.o file depends on the hello.c file. If the

hello.c file has changed, the command $(CC) $(CFLAGS) -c -o hello.o hello.c

is to be executed to update the target file hello.o.

A rule can have an arbitrary number of commands. The individual command lines are executed independently by the command shell. The commands do not have to begin with a tab, but they must be indented from the dependency line.

In addition to normal variables, the following special variables may be used in the body of a rule.

• $*: the target name, without a suffix.

• $@: the target name.

• $^: a list of the sources, in alphabetical order, with duplicates removed.

• $+: all the sources, in the original order.

• $<: the first source.

For example, the above hello.c rule may be simplified as follows.

    hello.o: hello.c
        $(CC) $(CFLAGS) -c -o $@ $<

Unlike normal values, the variables in a rule body are expanded lazily, and binding is dynamic. The following function definition illustrates some of the issues.

    CLibrary(name, files) =
        OFILES = $(addsuffix .o, $(files))
        $(name).a: $(OFILES)
            $(AR) cq $@ $(OFILES)

This function defines a rule to build a program called $(name) from a list of .o files. The files in the argument are specified without a suffix, so the first line of the function definition defines a variable OFILES that adds the .o suffix to each of the file names. The next step defines a rule to build a target library $(name).a from the $(OFILES) files. The expression $(AR) is evaluated when the function is called, and the value of the variable AR is taken from the caller’s scope (see also the section on Scoping).

7.1 Implicit rules

Rules may also be implicit. That is, the files may be specified by wildcard patterns. The wildcard character is %. For example, the following rule specifies a default rule for building .o files.

    %.o: %.c
        $(CC) $(CFLAGS) -c -o $@ $*.c

This rule is a template for building an arbitrary .o file from a .c file.

By default, implicit rules are only used for the targets in the current directory. However subdirectories included via the .SUBDIRS rules inherit all the implicit rules that are in scope (see also the section on Scoping).

7.2 Bounded implicit rules

Implicit rules may specify the set of files they apply to. The following syntax is used.

    <targets>: <pattern>: <dependencies>
        <commands>

For example, the following rule applies only to the files a.o and b.o.

   a.o b.o: %.o: %.c
        $(CC) $(CFLAGS) -DSPECIAL -c $*.c

7.3 section

Frequently, the commands in a rule body are expressions to be evaluated by the shell. omake also allows expressions to be evaluated by omake itself.

The syntax of these “computed rules” uses the section expression. The following rule uses the omake IO functions to produce the target hello.c.

    hello.c:
        section
            FP = fopen(hello.c, w)
            fprintln($(FP), $""#include <stdio.h> int main() { printf("Hello world\n"); }"")
            close($(FP))

This example uses the quotation $""..."" (see also Section B.1.6) to quote the text being printed. These quotes are not included in the output file. The fopen, fprintln, and close functions perform file IO as discussed in the IO section.

In addition, commands that are function calls, or special expressions, are interpreted correctly. Since the fprintln function can take a file directly, the above rule can be abbreviated as follows.

    hello.c:
       fprintln($@, $""#include <stdio.h> int main() { printf("Hello world\n"); }"")

7.4 section rule

Rules can also be computed using the section rule form, where a rule body is expected instead of an expression. In the following rule, the file a.c is copied onto the hello.c file if it exists, otherwise hello.c is created from the file

default.c.
    hello.c:
        section rule
           if $(target-exists a.c)
              hello.c: a.c
                 cat a.c > hello.c
           else
              hello.c: default.c
                 cp default.c hello.c

7.5 Special dependencies

7.5.1 :exists:

In some cases, the contents of a dependency do not matter, only whether the file exists or not. In this case, the :exists: qualifier can be used for the dependency.

    foo.c: a.c :exists: .flag
       if $(test -e .flag)
           $(CP) a.c $@

7.5.2 :effects:

Some commands produce files by side-effect. For example, the latex(1) command produces a .aux file as a side-effect of producing a .dvi file. In this case, the :effects: qualifier can be used to list the side-effect explicitly. omake is careful to avoid simultaneously running programs that have overlapping sideeffects.

    paper.dvi: paper.tex :effects: paper.aux
        latex paper

7.5.3 :value:

The :value: dependency is used to specify that the rule execution depends on the value of an expression. For example, the following rule

    a: b c :value: $(X)
        ...

specifies that “a” should be recompiled if the value of $(X) changes (X does not have to be a filename). This is intended to allow greater control over dependencies.

In addition, it can be used instead of other kinds of dependencies. For example, the following rule:

    a: b :exists: c
        commands

is the same as

    a: b :value: $(target-exists c)
        commands

Notes:

• The values are arbitrary (they are not limited to variables)

• The values are evaluated at rule expansion time, so expressions containing variables like $@, $^, etc are legal.

7.6 .SCANNER rules

Scanner rules define a way to specify automatic dependency scanning. A .SCANNER rule has the following form.

    .SCANNER: target: dependencies
        commands

The rule is used to compute additional dependencies that might be defined in the source files for the specified target. The result of executing the scanner commands must be a sequence of dependencies in OMake format, printed to the standard output. For example, on GNU systems the gcc -MM foo.c produces dependencies for the file foo.c (based on #include information).

We can use this to specify a scanner for C files that adds the scanned dependencies for the .o file. The following scanner specifies that dependencies for a file, say foo.o can be computed by running gcc -MM foo.c. Furthermore, foo.c is a dependency, so the scanner should be recomputed whenever the foo.c file changes.

    .SCANNER: %.o: %.c
        gcc -MM $<

Let’s suppose that the command gcc -MM foo.c prints the following line.

    foo.o: foo.h /usr/include/stdio.h

The result is that the files foo.h and /usr/include/stdio.h are considered to be dependencies of foo.o—that is, foo.o should be rebuilt if either of these files changes.

This works, to an extent. One nice feature is that the scanner will be re-run whenever the foo.c file changes. However, one problem is that dependencies in C are recursive. That is, if the file foo.h is modified, it might include other files, establishing further dependencies. What we need is to re-run the scanner if foo.h changes too.

We can do this with a value dependency. The variable $& is defined as the dependency results from any previous scan. We can add these as dependencies using the digest function, which computes an MD5 digest of the files.

    .SCANNER: %.o: %.c :value: $(digest $&)
        gcc -MM $<

Now, when the file foo.h changes, its digest will also change, and the scanner will be re-run because of the value dependency (since $& will include foo.h).

This still is not quite right. The problem is that the C compiler uses a search-path for include files. There may be several versions of the file foo.h, and the one that is chosen depends on the include path. What we need is to base the dependencies on the search path.

The $(digest-in-path-optional ...) function computes the digest based on a search path, giving us a solution that works.

    .SCANNER: %.o: %.c :value: $(digest-in-path-optional $(INCLUDES), $&)
       gcc -MM $(addprefix -I, $(INCLUDES)) $<

The standard output of the scanner rules will be captured by OMake and is not allowed to contain any content that OMake will not be able to parse as a dependency. The output is allowed to contain dependency specifications for unrelated targets, however such dependencies will be ignored. The scanner rules are allowed to produce arbitrary output on the standard error channel — such output will be handled in the same way as the output of the ordinary rules (in other words, it will be presented to the user, when dictated by the --output-… options enabled).

Additional examples of the .SCANNER rules can be found in Section 3.4.3.

7.6.1 Named scanners, and the :scanner: dependencies

Sometimes it may be useful to specify explicitly which scanner should be used in a rule. For example, we might compile .c files with different options, or (heaven help us) we may be using both gcc and the Microsoft Visual C++ compiler cl. In general, the target of a .SCANNER is not tied to a particular target, and we may name it as we like.

    .SCANNER: scan-gcc-%.c: %.c :value: $(digest-in-path-optional $(INCLUDES), $&)
        gcc -MM $(addprefix -I, $(INCLUDES)) $<
    .SCANNER: scan-cl-%.c: %.c :value: $(digest-in-path-optional $(INCLUDES), $&)
        cl --scan-dependencies-or-something $(addprefix /I, $(INCLUDES)) $<

The next step is to define explicit scanner dependencies. The :scanner: dependency is used for this. In this case, the scanner dependencies are specified explicitly.

    $(GCC_FILES): %.o: %.c :scanner: scan-gcc-%c
        gcc ...
    $(CL_FILES): %.obj: %.c :scanner: scan-cl-%c
        cl ...

Explicit :scanner: scanner specification may also be used to state that a single .SCANNER rule should be used to generate dependencies for more than one target. For example,

    .SCANNER: scan-all-c: $(GCC_FILES) :value: $(digest-in-path-optional $(INCLUDES), $&)
        gcc -MM $(addprefix -I, $(INCLUDES)) $(GCC_FILES)
    $(GCC_FILES): %.o: %.c :scanner: scan-all-c
        ...

The above has the advantage of only running gcc once and a disadvantage that when a single source file changes, all the files will end up being re-scanned.

7.6.2 Notes

In most cases, you won’t need to define scanners of your own. The standard installation includes default scanners (both explicitly and implicitly named ones) for C, OCaml, and L^{A}TEX files.

The SCANNER MODE variable controls the usage of implicit scanner dependencies.

The explicit :scanner: dependencies reduce the chances of scanner misspecifications. In large complicated projects it might be a good idea to set SCANNER_MODE to error and use only the named .SCANNER rules and explicit :scanner: specifications.

7.7 .DEFAULT

The .DEFAULT target specifies a target to be built by default if omake is run without explicit targets. The following rule instructs omake to build the program hello by default

   .DEFAULT: hello

7.8 .SUBDIRS

The .SUBDIRS target is used to specify a set of subdirectories that are part of the project. Each subdirectory should have its own OMakefile, which is evaluated in the context of the current environment.

   .SUBDIRS: src doc tests

This rule specifies that the OMakefiles in each of the src, doc, and tests directories should be read.

In some cases, especially when the OMakefiles are very similar in a large number of subdirectories, it is inconvenient to have a separate OMakefile for each directory. If the .SUBDIRS rule has a body, the body is used instead of the

OMakefile.
   .SUBDIRS: src1 src2 src3
      println(Subdirectory $(CWD))
      .DEFAULT: lib.a

In this case, the src1, src2, and src3 files do not need OMakefiles. Furthermore, if one exists, it is ignored. The following includes the file if it exists.

   .SUBDIRS: src1 src2 src3
       if $(file-exists OMakefile)
          include OMakefile
       .DEFAULT: lib.a

7.9 .INCLUDE

The .INCLUDE target is like the include directive, but it specifies a rule to build the file if it does not exist.

   .INCLUDE: config
       echo "CONFIG_READ = true" > config
    echo CONFIG_READ is $(CONFIG_READ)

You may also specify dependencies to an .INCLUDE rule.

   .INCLUDE: config: config.defaults
      cp config.defaults config

A word of caution is in order here. The usual policy is used for determining when the rule is out-of-date. The rule is executed if any of the following hold.

• the target does not exist,

• the rule has never been executed before,

• any of the following have changed since the last time the rule was executed,

– the target,

– the dependencies,

– the commands-text.

In some of the cases, this will mean that the rule is executed even if the target file already exists. If the target is a file that you expect to edit by hand (and therefore you don’t want to overwrite it), you should make the rule evaluation conditional on whether the target already exists.

   .INCLUDE: config: config.defaults
       # Don't overwrite my carefully hand-edited file
       if $(not $(file-exists config))
          cp config.defaults config

7.10 .PHONY

A “phony” target is a target that is not a real file, but exists to collect a set of dependencies. Phony targets are specified with the .PHONY rule. In the following example, the install target does not correspond to a file, but it corresponds to some commands that should be run whenever the install target is built (for example, by running omake install).

   .PHONY: install
   install: myprogram.exe
      cp myprogram.exe /usr/bin

7.11 Rule scoping

As we have mentioned before, omake is a scoped language. This provides great flexibility—different parts of the project can define different configurations without interfering with one another (for example, one part of the project might be compiled with CFLAGS=-O3 and another with CFLAGS=-g).

But how is the scope for a target file selected? Suppose we are building a file dir/foo.o. omake uses the following rules to determine the scope.

• First, if there is an explicit rule for building dir/foo.o (a rule with no wildcards), the context for that rule determines the scope for building the target.

• Otherwise, the directory dir/ must be part of the project. This normally means that a configuration file dir/OMakefile exists (although, see the .SUBDIRS section for another way to specify the OMakefile). In this case, the scope of the target is the scope at the end of the dir/OMakefile.

To illustrate rule scoping, let’s go back to the example of a “Hello world” program with two files. Here is an example OMakefile (the two definitions of CFLAGS are for illustration).

    # The executable is compiled with debugging
    CFLAGS = -g
    hello: hello_code.o hello_lib.o
       $(CC) $(CFLAGS) -o $@ $+
    # Redefine CFLAGS
    CFLAGS += -O3

In this project, the target hello is explicit. The scope of the hello target is the line beginning with hello:, where the value of CFLAGS is -g. The other two targets, hello_code.o and hello_lib.o do not appear as explicit targets, so their scope is at the end of the OMakefile, where the CFLAGS variable is defined to be -g -O3. That is, hello will be linked with CFLAGS=-g and the .o files will be compiled with CFLAGS=-g -O3.

We can change this behavior for any of the targets by specifying them as explicit targets. For example, suppose we wish to compile hello_lib.o with a preprocessor variable LIBRARY.

    # The executable is compiled with debugging
    CFLAGS = -g
    hello: hello_code.o hello_lib.o
       $(CC) $(CFLAGS) -o $@ $+
    # Compile hello_lib.o with CFLAGS = -g -DLIBRARY
    section

        CFLAGS += -DLIBRARY
        hello_lib.o:
    # Redefine CFLAGS
    CFLAGS += -O3

In this case, hello_lib.o is also mentioned as an explicit target, in a scope where CFLAGS=-g -DLIBRARY. Since no rule body is specified, it is compiled using the usual implicit rule for building .o files (in a context where

CFLAGS=-g -DLIBRARY).

7.11.1 Scoping of implicit rules

Implicit rules (rules containing wildcard patterns) are not global, they follow the normal scoping convention. This allows different parts of a project to have different sets of implicit rules. If we like, we can modify the example above to provide a new implicit rule for building hello_lib.o.

    # The executable is compiled with debugging
    CFLAGS = -g
    hello: hello_code.o hello_lib.o
       $(CC) $(CFLAGS) -o $@ $+
    # Compile hello_lib.o with CFLAGS = -g -DLIBRARY
    section
        %.o: %.c
            $(CC) $(CFLAGS) -DLIBRARY -c $<
        hello_lib.o:
    # Redefine CFLAGS
    CFLAGS += -O3

In this case, the target hello_lib.o is built in a scope with a new implicit rule for building %.o files. The implicit rule adds the -DLIBRARY option. This implicit rule is defined only for the target hello_lib.o; the target hello_code.o is built as normal.

7.11.2 Scoping of .SCANNER rules

Scanner rules are scoped the same way as normal rules. If the .SCANNER rule is explicit (containing no wildcard patterns), then the scope of the scan target is the same as the the rule. If the .SCANNER rule is implicit, then the environment is taken from the :scanner: dependency.

    # The executable is compiled with debugging
    CFLAGS = -g
    hello: hello_code.o hello_lib.o

   $(CC) $(CFLAGS) -o $@ $+
# scanner for .c files
.SCANNER: scan-c-%.c: %.c
   $(CC) $(CFLAGS) -MM $<
# Compile hello_lib.o with CFLAGS = -g -DLIBRARY
section
    CFLAGS += -DLIBRARY
    hello_lib.o: hello_lib.c :scanner: scan-c-hello_lib.c
       $(CC) $(CFLAGS) -c $<
# Compile hello_code.c with CFLAGS = -g -O3
section
    CFLAGS += -O3
    hello_code.o: hello_code.c :scanner: scan-c-hello_code.c
       $(CC) $(CFLAGS) -c $<

Again, this is for illustration—it is unlikely you would need to write a complicated configuration like this! In this case, the .SCANNER rule specifies that the C-compiler should be called with the -MM flag to compute dependencies. For the target hello_lib.o, the scanner is called with CFLAGS=-g -DLIBRARY, and

for hello_code.o it is called with CFLAGS=-g -O3.

7.11.3 Scoping for .PHONY targets

Phony targets (targets that do not correspond to files) are defined with a .PHONY: rule. Phony targets are scoped as usual. The following illustrates a common mistake, where the .PHONY target is declared after it is used.

# !!This example is broken!!
all: hello
hello: hello_code.o hello_lib.o
    $(CC) $(CFLAGS) -o $@ $+
.PHONY: all

This doesn’t work as expected because the .PHONY declaration occurs too late. The proper way to write this example is to place the .PHONY declaration first.

# Phony targets must be declared before being used
.PHONY: all
all: hello

    hello: hello_code.o hello_lib.o
        $(CC) $(CFLAGS) -o $@ $+

Phony targets are passed to subdirectories. As a practical matter, it is wise to declare all .PHONY targets in your root OMakefile, before any .SUBDIRS. This will ensure that 1) they are considered as phony targets in each of the subdirectories, and 2) you can build them from the project root.

    .PHONY: all install clean
    .SUBDIRS: src lib clib

Note that when a .PHONY target is inherited by a subdirectory via a .SUBDIRS, a whole hierarchy of .PHONY target (that are a part of the global one) is created, as described in Section 7.12.2 below.

7.12 Running OMake from a subdirectory

Running omake foo asks OMake to build the file foo in context of the whole project, even when running from a subdirectory of the project. Therefore, if bar/baz is a regular target (not a .PHONY one), then running omake bar/baz and running (cd bar; omake baz) are usually equivalent.

There are two noteworthy exceptions to the above rule:

• If the subdirectory is not a part of the project (there is no .SUBDIRS) for it, then OMake will complain if you try to run it in that directory.

• If a subdirectory contains an OMakeroot of its own, this would designate the subdirectory as a separate project (which is usually a bad idea and is not recommended).

7.12.1 Phony targets in a subdirectory

Suppose you have a .PHONY: clean declared in your root OMakefile and both the root OMakefile and the OMakefile in some of the subdirectories contain clean: rules. In this case

• Running omake clean in the root directory will execute all the rules (each in the appropriate directory);

• Running omake clean in the subdirectory will execute just its local one, as well as the ones from the subdirectories of the current directory.

The above equally applies to the built-in .PHONY targets, including .DEFAULT. Namely, if OMake is executed (without argument) in the root directory of a project, all the .DEFAULT targets in the project will be built. On the other hand, when OMake is executed (without argument) in a subdirectory, only the .DEFAULT targets defined in and under that subdirectory will be built.

The following Section explains the underlying semantics that gives rise to the above behavior.

7.12.2 Hierarchy of .PHONY targets

When the the root OMakefile contains a .PHONY: clean directive, it creates:

• A “global” phony target /.PHONY/clean (note the leading “/”);

• A “relative” phony target attached to the current directory — .PHONY/clean (note the lack of the leading “/”);

   • A dependency /.PHONY/clean: .PHONY/clean.

All the clean: ... rules in the root OMakefile following this .PHONY: clean declaration would be interpreted as rules for the .PHONY/clean target.

Now when OMake then comes across a .SUBDIRS: foo directive (when it is in scope of the above .PHONY: clean declaration), it does the following:

• Creates a new .PHONY/foo/clean “relative” phony target;

   • Creates the dependency .PHONY/clean: .PHONY/foo/clean;

• Processes the body of the .SUBDIRS: foo directive, or reads the foo/OMakefile file, if the body is empty. While doing that, it interprets its instructions relative to the foo directory. In particular, all the clean: ... rules will be taken to apply to .PHONY/foo/clean.

Now when you run omake clean in the root directory of the project, it is interpreted as omake .PHONY/clean (similar to how it happens with the normal targets), so both the rules for .PHONY/clean are executed and the rules for its

dependency .PHONY/foo/clean. Running (cd foo; omake clean) is, as for

normal targets, equivalent to running omake .PHONY/foo/clean and only those rules that apply to .PHONY/foo/clean will be executed.

7.13 Pathnames in rules

In rules, the targets and dependencies are first translated to file values (as in the file function). They are then translated to strings for the command line. This can cause some unexpected behavior. In the following example, the absname function is the absolute pathname for the file a, but the rule still prints the relative pathname.

    .PHONY: demo
    demo: $(absname a)
        echo $<
    # omake demo
    a

There is arguably a good reason for this. On Win32 systems, the / character is viewed as an “option specifier.” The pathname separator is the \ character. OMake translates the filenames automatically so that things work as expected on both systems.

demo: a/b
    echo $<
# omake demo (on a Unix system)
a/b
# omake demo (on a Win32 system)
a\b

Sometimes you may wish that target strings to be passed literally to the commands in the rule. One way to do this is to specify them literally.

 SRC = a/b $(absname c/d)
 demo: $(SRC)
     echo $(SRC)
 # omake demo (on a Win32 system)
 a/b c:\...\c\d

Alternately, you might wish that filenames be automatically expanded to absolute pathnames. For example, this might be useful when parsing the OMake output to look for errors. For this, you can use the --absname option (Section A.3.20). If you call omake with the --absname option, all filenames will be expanded to absolute names.

 # omake --absname demo (on a Unix system)
 /home/.../a/b /home/.../c/d

Alternately, the --absname option is scoped. If you want to use it for only a few rules, you can use the OMakeFlags function to control how it is applied.

section
   OMakeFlags(--absname)
   demo: a
       echo $<
# omake demo
/home/.../a

N.B. The --absname option is currently an experimental feature.

Base library

8.1 Builtin variables

OMAKE VERSION Version of OMake.

STDLIB The directory where the OMake standard library files reside. At startup, the default value is determined as follows.

• The value of the OMAKELIB environment variable, if set (must contain an absolute path, if set), otherwise

   • On Windows, the registry keys HKEY_CURRENT_USER\SOFTWARE\MetaPRL\OMake\OMAKELIB
     and HKEY_LOCAL_MACHINE\SOFTWARE\MetaPRL\OMake\OMAKELIB are looked

up and the value is used, if exist.

• Otherwise a compile-time default it used.

The current default value may be accessed by running omake --version

OMAKEPATH An array of directories specifying the lookup path for the include and open directives (see Section 4.7). The default value is an array of two elements — . and $(STDLIB).

OSTYPE Set to the machine architecture omake is running on. Possible values are Unix (for all Unix versions, including Linux and Mac OS X), Win32 (for MS-Windows, OMake compiled with MSVC++ or Mingw), and Cygwin (for MS-Windows, OMake compiled with Cygwin).

SYSNAME The name of the operating system for the current machine.

NODENAME The hostname of the current machine.

OS VERSION The operating system release.

MACHINE The machine architecture, e.g. i386, sparc, etc.

HOST Same as NODENAME.

USER The login name of the user executing the process.

HOME The home directory of the user executing the process.

PID The OMake process id.

TARGETS The command-line target strings. For example, if OMake is invoked with the following command line,

      omake CFLAGS=1 foo bar.c
   then TARGETS is defined as foo bar.c.

BUILD SUMMARY The BUILD_SUMMARY variable refers to the file that omake uses to summarize a build (the message that is printed at the very end of a build). The file is empty when the build starts. If you wish to add additional messages to the build summary, you can edit/modify this file during the build.

For example, if you want to point out that some action was taken, you can append a message to the build summary.

   foo: boo
       echo "The file foo was built" >> $(BUILD_SUMMARY)
       ...build foo...

VERBOSE Whether certain commands should be verbose. A boolean flag that is false by default and is set to true when OMake is invoked with the

--verbose option.

8.2 Logic, Boolean functions, and control flow

Boolean values in omake are represented by case-insensitive strings. The false value can be represented by the strings false, no, nil, undefined or 0, and everything else is true.

8.2.1 not

$(not e) : String
  e : String

The not function negates a Boolean value. For example, $(not false) expands to the string true, and $(not hello world)

expands to false.

8.2.2 equal

$(equal e1, e2) : String
  e1 : String
  e2 : String

The equal function tests for equality of two values.

For example $(equal a, b) expands to false, and $(equal hello world, hello world)

expands to true.

8.2.3 and

$(and e1, ..., en) : String
   e1, ..., en: Sequence

The and function evaluates to the conjunction of its arguments. For example, in the following code, X is true, and Y is false.

A = a
B = b
X = $(and $(equal $(A), a) true $(equal $(B), b))
Y = $(and $(equal $(A), a) true $(equal $(A), $(B)))

8.2.4 or

$(or e1, ..., en) : String
  e1, ..., en: String Sequence

The or function evaluates to the disjunction of its arguments. For example, in the following code, X is true, and Y is false.

A = a
B = b
X = $(or $(equal $(A), a) false $(equal $(A), $(B)))
Y = $(or $(equal $(A), $(B)) $(equal $(A), b))

8.2.5 if

    $(if e1, e2[, e3]) : value
       e1 : String
       e2, e3 : value

The if function represents a conditional based on a Boolean value. For

example $(if $(equal a, b), c, d) evaluates to d.

Conditionals may also be declared with an alternate syntax.

   if e1
      body1
   elseif e2
      body2
   ...
   else
      bodyn

If the expression e1 is not false, then the expressions in body1 are evaluated and the result is returned as the value of the conditional. Otherwise, if e1 evaluates to false, the evaluation continues with the e2 expression. If none of the conditional expressions is true, then the expressions in bodyn are evaluated and the result is returned as the value of the conditional.

There can be any number of elseif clauses; the else clause is optional.

Note that each branch of the conditional defines its own scope, so variables defined in the branches are normally not visible outside the conditional. The export command may be used to export the variables defined in a scope. For example, the following expression represents a common idiom for defining the C compiler configuration.

   if $(equal $(OSTYPE), Win32)
      CC = cl
      CFLAGS += /DWIN32
      export
   else
      CC = gcc
      CFLAGS += -g -O2
      export

8.2.6 switch, match

The switch and match functions perform pattern matching.

   $(switch <arg>, <pattern_1>, <value_1>, ..., <pattern_n>, <value_n>)
$(match <arg>, <pattern_1>, <value_1>, ..., <pattern_n>, <value_n>)

The number of <pattern>/<value> pairs is arbitrary. They strictly alternate; the total number of arguments to <match> must be odd.

The <arg> is evaluated to a string, and compared with <pattern_1>. If it matches, the result of the expression is <value_1>. Otherwise evaluation continues with the remaining patterns until a match is found. If no pattern matches, the value is the empty string.

The switch function uses string comparison to compare the argument with the patterns. For example, the following expression defines the FILE variable to be either foo, bar, or the empty string, depending on the value of the OSTYPE variable.

 FILE = $(switch $(OSTYPE), Win32, foo, Unix, bar)

The match function uses regular expression patterns (see the grep function). If a match is found, the variables $1, $2, ... are bound to the substrings matched between \( and \) delimiters. The $0 variable contains the entire match, and $* is an array of the matched substrings. to the matched substrings.

 FILE = $(match foo_xyz/bar.a, foo_\\\(.*\\\)/\\\(.*\\\)\.a, foo_$2/$1.o)

The switch and match functions also have an alternate (more usable) form.

match e
case pattern1
   body1
case pattern2
   body2
...
default
   bodyd

If the value of expression e matches pattern_i and no previous pattern, then body_i is evaluated and returned as the result of the match. The switch function uses string comparison; the match function uses regular expression matching.

match $(FILE)
case $".*\(\.[^\/.]*\)"
   println(The string $(FILE) has suffix $1)
default
   println(The string $(FILE) has no suffix)

8.2.7 try

try
   try-body
catch class1(v1)
   catch-body
when expr
   when-body
...
finally
   finally-body

The try form is used for exception handling. First, the expressions in the try-body are evaluated.

If evaluation results in a value v without raising an exception, then the expressions in the finally-body are evaluated and the value v is returned as the result.

If evaluation of the try-body results in a exception object obj, the catch clauses are examined in order. When examining catch clause catch class(v), if the exception object obj is an instance of the class name class, the variable v is bound to the exception object, and the expressions in the catch-body are evaluated.

If a when clause is encountered while a catch body is being evaluated, the predicate expr is evaluated. If the result is true, evaluation continues with the expressions in the when-body. Otherwise, the next catch clause is considered for evaluation.

If evaluation of a catch-body or when-body completes successfully, returning a value v, without encountering another when clause, then the expressions in the finally-body are evaluated and the value v is returned as the result.

There can be any number of catch clauses; the finally clause is optional.

8.2.8 raise

   raise exn
      exn : Exception

The raise function raises an exception. The exn object can be any object. However, the normal convention is to raise an Exception object.

If the exception is never caught, the whole object will be verbosely printed in the error message. However, if the object is an Exception one and contains a message field, only that field will be included in the error message.

8.2.9 exit

   exit(code)
      code : Int

The exit function terminates omake abnormally.

   $(exit <code>)

The exit function takes one integer argument, which is exit code. Non-zero values indicate abnormal termination.

8.2.10 defined

   $(defined sequence) : String
      sequence : Sequence

The defined function test whether all the variables in the sequence are currently defined. For example, the following code defines the X variable if it is not already defined.

    if $(not $(defined X))
       X = a b c
       export

8.2.11 defined-env

   $(defined-env sequence) : String
      sequence : String

The defined-env function tests whether a variable is defined as part of the process environment.

For example, the following code adds the -g compile option if the environment variable DEBUG is defined.

if $(defined-env DEBUG)
    CFLAGS += -g
    export

8.2.12 getenv

   $(getenv name) : String
   $(getenv name, default) : String

The getenv function gets the value of a variable from the process environment. The function takes one or two arguments.

In the single argument form, an exception is raised if the variable variable is not defined in the environment. In the two-argument form, the second argument is returned as the result if the value is not defined.

For example, the following code defines the variable X to be a space-separated list of elements of the PATH environment variable if it is defined, and to /bin /usr/bin otherwise.

    X = $(split $(PATHSEP), $(getenv PATH, /bin:/usr/bin))

You may also use the alternate form.

     getenv(NAME)
         default

8.2.13 setenv

   setenv(name, value)
      name : String
      value : String

The setenv function sets the value of a variable in the process environment. Environment variables are scoped like normal variables.

8.2.14 unsetenv

unsetenv(names)
   names : String Array

The unsetenv function removes some variable definitions from the process environment. Environment variables are scoped like normal variables.

8.2.15 get-registry

get-registry(hkey, key, field) : String
get-registry(hkey, key, field, default) : String
    hkey : String
    key : String
    field : String

The get-registry function retrieves a string value from the system registry on Win32. On other architectures, there is no registry.

The hive (I think that is the right word), indicates which part of the registry to use. It should be one of the following values.

• HKEY_CLASSES_ROOT
• HKEY_CURRENT_CONFIG
• HKEY_CURRENT_USER
• HKEY_LOCAL_MACHINE
• HKEY_USERS

Refer to the Microsoft documentation if you want to know what these mean.

The key is the field you want to get from the registry. It should have a form like A\B\C (if you use forward slashes, they will be converted to backslashes). The field is the sub-field of the key.

In the 4-argument form, the default is returned on failure. You may also use the alternate form.

 get-registry(hkey, key, field)
    default

8.2.16 getvar

$(getvar name) : String

The getvar function gets the value of a variable. An exception is raised if the variable variable is not defined. For example, the following code defines X to be the string abc.

 NAME = foo
 foo_1 = abc
 X = $(getvar $(NAME)_1)

8.2.17 setvar

   setvar(name, value)
      name : String
      value : String

The setvar function defines a new variable. For example, the following code defines the variable X to be the string abc.

   NAME = X
   setvar($(NAME), abc)

8.3 Arrays and sequences

8.3.1 array

    $(array elements) : Array
       elements : Sequence

The array function creates an array from a sequence. If the <arg> is a string, the elements of the array are the whitespace-separated elements of the string, respecting quotes.

In addition, array variables can be declared as follows.

    A[] =
       <val1>
       ...
       <valn>

In this case, the elements of the array are exactly <val1>, ..., <valn>, and whitespace is preserved literally.

8.3.2 split

   $(split sep, elements) : Array
      sep : String
      elements : Sequence

The split function takes two arguments, a string of separators, and a string argument. The result is an array of elements determined by splitting the elements by all occurrence of the separator in the elements sequence.

For example, in the following code, the X variable is defined to be the array

/bin /usr/bin /usr/local/bin.
    PATH = /bin:/usr/bin:/usr/local/bin
    X = $(split :, $(PATH))

The sep argument may be omitted. In this case split breaks its arguments along the white space. Quotations are not split.

8.3.3 concat

$(concat sep, elements) : String
   sep : String
   elements : Sequence

The concat function takes two arguments, a separator string, and a sequence of elements. The result is a string formed by concatenating the elements, placing the separator between adjacent elements.

For example, in the following code, the X variable is defined to be the string

foo_x_bar_x_baz.
 X = foo  bar     baz
 Y = $(concat _x_, $(X))

8.3.4 length

$(length sequence) : Int
   sequence : Sequence

The length function returns the number of elements in its argument. For example, the expression $(length a b "c d") evaluates to 3.

8.3.5 nth

$(nth i, sequence) : value
   i : Int
   sequence : Sequence
raises RuntimeException

The nth function returns the nth element of its argument, treated as a list. Counting starts at 0. An exception is raised if the index is not in bounds.

For example, the expression $(nth 1, a "b c" d) evaluates to "b c".

8.3.6 nth-hd

$(nth-hd i, sequence) : value
   i : Int
   sequence : Sequence
raises RuntimeException

The nth-hd function returns the first i elements of the sequence. An exception is raised if the sequence is not at least i elements long.

For example, the expression $(nth-hd 2, a "b c" d) evaluates to a "b c".

8.3.7 nth-tl

   $(nth-tl i, sequence) : value
      i : Int
      sequence : Sequence
   raises RuntimeException

The nth-tl function skips i elements of the sequence and returns the rest. An exception is raised if the sequence is not at least i elements long.

For example, the expression $(nth-tl 1, a "b c" d) evaluates to "b c" d.

8.3.8 subrange

   $(subrange off, len, sequent) : value
      off : Int
      len : Int
      sequence : Sequence
   raises RuntimeException

The subrange function returns a subrange of the sequence. Counting starts at 0. An exception is raised if the specified range is not in bounds.

For example, the expression $(subrange 1, 2, a "b c" d e) evaluates to

"b c" d.

8.3.9 rev

    $(rev sequence) : Sequence
       sequence : Sequence

The rev function returns the elements of a sequence in reverse order. For example, the expression $(rev a "b c" d) evaluates to d "b c" a.

8.3.10 join

   $(join sequence1, sequence2) : Sequence
      sequence1 : Sequence
      sequence2 : Sequence

The join function joins together the elements of the two sequences. For

example, $(join a b c, .c .cpp .h) evaluates to a.c b.cpp c.h. If the two

input sequences have different lengths, the remainder of the longer sequence is copied at the end of the output unmodified.

8.3.11 string

   $(string sequence) : String
      sequence : Sequence

The string function flattens a sequence into a single string. This is similar to the concat function, but the elements are separated by whitespace. The result is treated as a unit; whitespace is significant.

8.3.12 string-escaped, ocaml-escaped, html-escaped, htmlpre-escaped, c-escaped, id-escaped

   $(string-escaped sequence) : String Array
   $(ocaml-escaped sequence) : String Array
   $(html-escaped sequence) : String Array
   $(html-pre-escaped sequence) : String Array
   $(c-escaped sequence) : String Array
   $(hex-escaped sequence) : StringArray
      sequence : Array

The string-escaped function converts each element of its argument to a string, escaping it, if it contains symbols that are special to OMake. The special characters include :()\,$'"# and whitespace. This function can be used in scanner rules to escape file names before printing then to stdout.

The ocaml-escaped function converts each element of its argument to a string, escaping characters that are special to OCaml.

The c-escaped function converts a string to a form that can be used as a string constant in C.

The id-escaped function turns a string into an identifier that may be used in OMake.

The html-escaped function turns a literal string into a form acceptable as HTML. The html-pre-escaped function is similar, but it does not translate newlines into <br>.

    println($(string $(string-escaped $"a b" $"y:z")))
    a\ b y\:z

8.3.13 decode-uri, encode-uri

    $(decode-uri sequence) : sequence
        sequence : Sequence

These two functions perform URI encoding, where special characters are represented by hexadecimal characters.

    osh> s = $(encode-uri $'a b~c')
    "a+b%7ec"
    osh> decode-uri($s)
    "a b~c"

8.3.14 quote

   $(quote sequence) : String
      sequence : Sequence

The quote function flattens a sequence into a single string and adds quotes around the string. Inner quotation symbols are escaped.

For example, the expression $(quote a "b c" d) evaluates to "a \"b c\" d",

and $(quote abc) evaluates to "abc".

8.3.15 quote-argv

   $(quote-argv sequence) : String
      sequence : Sequence

The quote-argv function flattens a sequence into a single string, and adds quotes around the string. The quotation is formed so that a command-line parse can separate the string back into its components.

8.3.16 html-string

   $(html-string sequence) : String
      sequence : Sequence

The html-string function flattens a sequence into a single string, and escaped special HTML characters. This is similar to the concat function, but the elements are separated by whitespace. The result is treated as a unit; whitespace is significant.

8.3.17 addsuffix

   $(addsuffix suffix, sequence) : Array
      suffix : String
      sequence : Sequence

The addsuffix function adds a suffix to each component of sequence. The number of elements in the array is exactly the same as the number of elements in the sequence.

   For example, $(addsuffix .c, a b "c d") evaluates to a.c b.c "c d".c.

8.3.18 mapsuffix

   $(mapsuffix suffix, sequence) : Array
      suffix : value
      sequence : Sequence

The mapsuffix function adds a suffix to each component of sequence. It is similar to addsuffix, but uses array concatenation instead of string concatenation. The number of elements in the array is twice the number of elements in the sequence.

   For example, $(mapsuffix .c, a b "c d") evaluates to a .c b .c "c d" .c.

8.3.19 addsuffixes

   $(addsuffixes suffixes, sequence) : Array
      suffixes : Sequence
      sequence : Sequence

The addsuffixes function adds all suffixes in its first argument to each component of a sequence. If suffixes has n elements, and sequence has m elements, the the result has n * m elements.

For example, the $(addsuffixes .c .o, a b c) expressions evaluates to

a.c a.o b.c b.o c.o c.a.

8.3.20 removeprefix

   $(removeprefix prefix, sequence) : Array
      prefix : String
      sequence : Array

The removeprefix function removes a prefix from each component of a sequence.

8.3.21 removesuffix

   $(removesuffix sequence) : Array
      sequence : String

The removesuffix function removes the suffixes from each component of a sequence.

   For example, $(removesuffix a.c b.foo "c d") expands to a b "c d".

8.3.22 replacesuffixes

   $(replacesuffixes old-suffixes, new-suffixes, sequence) : Array
      old-suffixes : Sequence
      new-suffixes : Sequence
      sequence : Sequence

The replacesuffixes function modifies the suffix of each component in sequence. The old-suffixes and new-suffixes sequences should have the same length.

   For example, $(replacesuffixes .h .c, .o .o, a.c b.h c.z) expands
to a.o b.o c.z.

8.3.23 addprefix

$(addprefix prefix, sequence) : Array
prefix : String
sequence : Sequence

The addprefix function adds a prefix to each component of a sequence. The number of element in the result array is exactly the same as the number of elements in the argument sequence.

For example, $(addprefix foo/, a b "c d") evaluates to foo/a foo/b foo/"c d".

8.3.24 mapprefix

$(mapprefix prefix, sequence) : Array
prefix : String
sequence : Sequence

The mapprefix function adds a prefix to each component of a sequence. It is similar to addprefix, but array concatenation is used instead of string concatenation. The result array contains twice as many elements as the argument sequence.

For example, $(mapprefix foo, a b "c d") expands to foo a foo b foo "c d".

8.3.25 add-wrapper

$(add-wrapper prefix, suffix, sequence) : Array
prefix : String
suffix : String
sequence : Sequence

The add-wrapper functions adds both a prefix and a suffix to each component of a sequence. For example, the expression $(add-wrapper dir/, .c, a b) evaluates to dir/a.c dir/b.c. String concatenation is used. The array result has the same number of elements as the argument sequence.

8.3.26 set

$(set sequence) : Array
sequence : Sequence

The set function sorts a set of string components, eliminating duplicates.

For example, $(set z y z "m n" w a) expands to "m n" a w y z.

8.3.27 mem

$(mem elem, sequence) : Boolean
elem : String
sequence : Sequence

The mem function tests for membership in a sequence.

   For example, $(mem "m n", y z "m n" w a) evaluates to true, while $(mem m n, y z "m n" w a)

evaluates to false.

8.3.28 intersection

   $(intersection sequence1, sequence2) : Array
      sequence1 : Sequence
      sequence2 : Sequence

The intersection function takes two arguments, treats them as sets of strings, and computes their intersection. The order of the result is undefined, and it may contain duplicates. Use the set function to sort the result and eliminate duplicates in the result if desired.

For example, the expression $(intersection c a b a, b a) evaluates to

a b a.

8.3.29 intersects

   $(intersects sequence1, sequence2) : Boolean
      sequence1 : Sequence
      sequence2 : Sequence

The intersects function tests whether two sets have a non-empty intersection. This is slightly more efficient than computing the intersection and testing whether it is empty.

For example, the expression $(intersects a b c, d c e) evaluates to

true, and $(intersects a b c a, d e f) evaluates to false.

8.3.30 set-diff

   $(set-diff sequence1, sequence2) : Array
      sequence1 : Sequence
      sequence2 : Sequence

The set-diff function takes two arguments, treats them as sets of strings, and computes their difference (all the elements of the first set that are not present in the second one). The order of the result is undefined and it may contain duplicates. Use the set function to sort the result and eliminate duplicates in the result if desired.

For example, the expression $(set-diff c a b a e, b a) evaluates to

c e.

8.3.31 filter

   $(filter patterns, sequence) : Array
      patterns : Sequence
      sequence : Sequence

The filter function picks elements from a sequence. The patterns is a nonempty sequence of patterns, each may contain one occurrence of the wildcard % character.

   For example $(filter %.h %.o, a.c x.o b.h y.o "hello world".c) eval-
uates to x.o b.h y.o.

8.3.32 filter-out

   $(filter-out patterns, sequence) : Array
      patterns : Sequence
      sequence : Sequence

The filter-out function removes elements from a sequence. The patterns is a non-empty sequence of patterns, each may contain one occurrence of the wildcard % character.

   For example $(filter-out %.c %.h, a.c x.o b.h y.o "hello world".c)

evaluates to x.o y.o.

8.3.33 capitalize

   $(capitalize sequence) : Array
      sequence : Sequence

The capitalize function capitalizes each word in a sequence. For example,

$(capitalize through the looking Glass) evaluates to Through The Looking Glass.

8.3.34 uncapitalize

   $(uncapitalize sequence) : Array
      sequence : Sequence

The uncapitalize function uncapitalizes each word in its argument.

   For example, $(uncapitalize through the looking Glass) evaluates to
through the looking glass.

8.3.35 uppercase

   $(uppercase sequence) : Array
      sequence : Sequence

The uppercase function converts each word in a sequence to uppercase. For

example, $(uppercase through the looking Glass) evaluates to THROUGH THE LOOKING GLASS.

8.3.36 lowercase

   $(lowercase sequence) : Array
      sequence : Sequence

The lowercase function reduces each word in its argument to lowercase.

   For example, $(lowercase through tHe looking Glass) evaluates to through the looking glass.

8.3.37 system

   system(s)
      s : Sequence

The system function is used to evaluate a shell expression. This function is used internally by omake to evaluate shell commands.

For example, the following program is equivalent to the expression system(ls foo).

   ls foo

8.3.38 shell

   $(shell command) : Array
   $(shella command) : Array
   $(shell-code command) : Int
      command : Sequence

The shell function evaluates a command using the command shell, and returns the whitespace-separated words of the standard output as the result.

The shella function acts similarly, but it returns the lines as separate items in the array.

The shell-code function returns the exit code. The output is not diverted.

For example, if the current directory contains the files OMakeroot, OMakefile,

and hello.c, then $(shell ls) evaluates to hello.c OMakefile OMakeroot

(on a Unix system).

8.3.39 export

The export function allows one to capture the current environment in a variable. For example, the following code:

A = 1
B = 1
C = 1
SAVE_ENV = $(export A B)
A = 2
B = 2
C = 2
export $(SAVE_ENV)
println($A $B $C)

will print 1 1 2.

The arguments to this function are interpreted the exact same way as the arguments to the export special form (see Section 5.3).

8.3.40 while

while <test>
  <body>

–or–

while <test>
case <test1>
   <body1>
...
case <testn>
   <bodyn>
default
   <bodyd>

The loop is executed while the test is true. In the first form, the <body> is executed on every loop iteration. In the second form, the body <bodyI> is selected, as the first case where the test <testI> is true. If none apply, the optional default case is evaluated. If no cases are true, the loop exits. The environment is automatically exported.

Examples. Iterate for i from 0 to 9.

i = 0
while $(lt $i, 10)
   echo $i
   i = $(add $i, 1)

The following example is equivalent.

i = 0
while true
case $(lt $i, 10)
  echo $i
  i = $(add $i, 1)

The following example is similar, but some special cases are printed. value is printed.

i = 0
while $(lt $i, 10)
case $(equal $i, 0)
   echo zero

 case $(equal $i, 1)
    echo one
 default
    echo $i

The break function can be used to break out of the while loop early.

8.3.41 break

break

Terminate execution of the innermost loop, returning the current state.

8.3.42 random, random-init

 random-init(i)
     i : Int
 random() : Int

Produce a random number. The numbers are pseudo-random, and are not cryptographically secure.

The generator is initialized form semi-random system data. Subsequent runs should produce different results. The rando-init function can be used to return the generator to a known state.

8.4 Arithmetic

8.4.1 int

The int function can be used to create integers. It returns an Int object.

$(int 17).

8.4.2 float

The float function can be used to create floating-point numbers. It returns a Float object.

$(float 3.1415926).

8.4.3 Basic arithmetic

The following functions can be used to perform basic arithmetic.

• $(neg <numbers>): arithmetic inverse

• $(add <numbers>): addition.
• $(sub <numbers>): subtraction.

8.4. ARITHMETIC 113 • $(mul <numbers>): multiplication.

• $(div <numbers>): division.
• $(mod <numbers>): remainder.

• $(lnot <numbers>): bitwise inverse.

• $(land <numbers>): bitwise and.
• $(lor <numbers>): bitwise or.

• $(lxor <numbers>): bitwise exclusive-or.

• $(lsl <numbers>): logical shift left.

• $(lsr <numbers>): logical shift right.

• $(asr <numbers>): arithmetic shift right.

• $(min <numbers>): smallest element.

• $(max <numbers>): largest element.

8.4.4 Comparisons

The following functions can be used to perform numerical comparisons.

• $(lt <numbers>): less then.
• $(le <numbers>): no more than.
• $(eq <numbers>): equal.
• $(ge <numbers>): no less than.
• $(gt <numbers>): greater than.

• $(ult <numbers>): unsigned less than.

• $(ule <numbers>): unsigned greater than.

• $(uge <numbers>): unsigned greater than or equal.

• $(ugt <numbers>): unsigned greater than.

8.5 First-class functions

8.5.1 fun

The fun form introduces anonymous functions.

$(fun <v1>, ..., <vn>, <body>)

The last argument is the body of the function. The other arguments are the parameter names.

The three following definitions are equivalent.

F(X, Y) =
   return($(addsuffix $(Y), $(X)))
F = $(fun X, Y, $(addsuffix $(Y), $(X)))
F =
   fun(X, Y)
      value $(addsuffix $(Y), $(X))

8.5.2 apply

The apply operator is used to apply a function.

$(apply <fun>, <args>)

Suppose we have the following function definition.

F(X, Y) =
   return($(addsuffix $(Y), $(X)))

The the two expressions below are equivalent.

X = F(a b c, .c)
X = $(apply $(F), a b c, .c)

8.5.3 applya

The applya operator is used to apply a function to an array of arguments.

$(applya <fun>, <args>)

For example, in the following program, the value of Z is file.c.

F(X, Y) =
   return($(addsuffix $(Y), $(X)))
args[] =
   file
   .c
Z = $(applya $(F), $(args))

8.5.4 create-map, create-lazy-map

The create-map is a simplified form for creating Map objects. The create-map function takes an even number of arguments that specify key/value pairs. For example, the following values are equivalent.

X = $(create-map name1, xxx, name2, yyy)
X. =
    extends $(Map)
    $|name1| = xxx
    $|name2| = yyy

The create-lazy-map function is similar, but the values are computed lazily. The following two definitions are equivalent.

Y = $(create-lazy-map name1, $(xxx), name2, $(yyy))
Y. =
    extends $(Map)
    $|name1| = $`(xxx)
    $|name2| = $`(yyy)

The create-lazy-map function is used in rule construction.

8.6 Iteration and mapping

8.6.1 foreach

The foreach function maps a function over a sequence.

$(foreach <fun>, <args>)
foreach(<var>, <args>)
   <body>

For example, the following program defines the variable X as an array a.c b.c c.c.

X =
   foreach(x, a b c)
      value $(x).c
# Equivalent expression
X = $(foreach $(fun x, $(x).c), abc)

There is also an abbreviated syntax. The export form can also be used in a foreach body. The final value of X

is a.c b.c c.c.

X =
foreach(x, a b c)
   X += $(x).c
   export

The break function can be used to break out of the loop early.

File, I/O and system operations

9.1 File names

9.1.1 file, dir

   $(file sequence) : File Sequence
      sequence : Sequence
   $(dir sequence) : Dir Sequence
      sequence : Sequence

The file and dir functions define location-independent references to files and directories. In omake, the commands to build a target are executed in the target’s directory. Since there may be many directories in an omake project, the build system provides a way to construct a reference to a file in one directory, and use it in another without explicitly modifying the file name. The functions have the following syntax, where the name should refer to a file or directory.

For example, we can construct a reference to a file foo in the current directory.

   FOO = $(file foo)
   .SUBDIRS: bar

If the FOO variable is expanded in the bar subdirectory, it will expand to

../foo.

These commands are often used in the top-level OMakefile to provide locationindependent references to top-level directories, so that build commands may refer to these directories as if they were absolute.

   ROOT = $(dir .)
   LIB  = $(dir lib)
   BIN  = $(dir bin)

Once these variables are defined, they can be used in build commands in subdirectories as follows, where $(BIN) will expand to the location of the bin directory relative to the command being executed.

   install: hello
cp hello $(BIN)

9.1.2 tmpfile

    $(tmpfile prefix) : File
    $(tmpfile prefix, suffix) : File
        prefix : String
        suffix : String

The tmpfile function returns the name of a fresh temporary file in the temporary directory.

9.1.3 in

   $(in dir, exp) : String Array
      dir : Dir
      exp : expression

The in function is closely related to the dir and file functions. It takes a directory and an expression, and evaluates the expression in that effective directory. For example, one common way to install a file is to define a symbol link, where the value of the link is relative to the directory where the link is created.

The following commands create links in the $(LIB) directory.

    FOO = $(file foo)
    install:
       ln -s $(in $(LIB), $(FOO)) $(LIB)/foo

Note that the in function only affects the expansion of Node (File and Dir) values.

9.1.4 basename

   $(basename files) : String Sequence
      files : String Sequence

The basename function returns the base names for a list of files. The basename is the filename with any leading directory components removed.

   For example, the expression $(basename dir1/dir2/a.out /etc/modules.conf /foo.ml)
evaluates to a.out modules.conf foo.ml.

9.1.5 dirname

   $(dirname files) : String Sequence
      files : String Sequence

The dirname function returns the directory name for a list of files. The directory name is the filename with the basename removed. If a name does not have a directory part, the directory is “.”

   For example, the expression $(dirname dir1\dir2\a.out /etc/modules.conf /foo.ml bar.ml)
evaluates to dir1/dir2 /etc / ..

Note: this function is different from the dirof function. The function dirname is simple a function over strings, while dirof is a function on filenames.

9.1.6 rootname

   $(rootname files) : String Sequence
      files : String Sequence

The rootname function returns the root name for a list of files. The rootname is the filename with the final suffix removed.

   For example, the expression $(rootname dir1/dir2/a.out /etc/a.b.c /foo.ml)
evaluates to dir1/dir2/a /etc/a.b /foo.

9.1.7 dirof

   $(dirof files) : Dir Sequence
      files : File Sequence

The dirof function returns the directory for each of the listed files.

   For example, the expression $(dirof dir/dir2/a.out /etc/modules.conf /foo.ml)

evaluates to the directories dir1/dir2 /etc /.

9.1.8 fullname

   $(fullname files) : String Sequence
      files : File Sequence

The fullname function returns the pathname relative to the project root for each of the files or directories.

9.1.9 absname

   $(absname files) : String Sequence
      files : File Sequence

The absname function returns the absolute pathname for each of the files or directories.

9.1.10 homename

   $(homename files) : String Sequence
      files : File Sequence

The homename function returns the name of a file in tilde form, if possible. The unexpanded forms are computed lazily: the homename function will usually evaluate to an absolute pathname until the first tilde-expansion for the same directory.

9.1.11 suffix

   $(suffix files) : String Sequence
      files : StringSequence

The suffix function returns the suffixes for a list of files. If a file has no suffix, the function returns the empty string.

   For example, the expression $(suffix dir1/dir2/a.out /etc/a /foo.ml)

evaluates to .out .ml.

9.2 Path search

9.2.1 which

   $(which files) : File Sequence
      files : String Sequence

The which function searches for executables in the current command search path, and returns file values for each of the commands. It is an error if a command is not found.

9.2.2 where

The where function is similar to which, except it returns the list of all the locations of the given executable (in the order in which the corresponding directories appear in $PATH). In case a command is handled internally by the Shell object, the first string in the output will describe the command as a built-in function.

    % where echo
    echo is a Shell object method (a built-in function)
    /bin/echo

9.2.3 rehash

    rehash()

The rehash function resets all search paths.

9.2.4 exists-in-path

   $(exists-in-path files) : String
      files : String Sequence

The exists-in-path function tests whether all executables are present in the current search path.

9.2.5 digest

     $(digest files) : String Array
        file : File Array
     raises RuntimeException
     $(digest-optional files) : String Array
        file : File Array

The digest and digest-optional functions compute MD5 digests of files. The digest function raises an exception if a file does no exist. The digest-optional returns false if a file does no exist. MD5 digests are cached.

9.2.6 find-in-path

    $(find-in-path path, files) : File Array
       path : Dir Array
       files : String Array
    raises RuntimeException
    $(find-in-path-optional path, files) : File Array

The find-in-path function searches for the files in a search path. Only the tail of the filename is significant. The find-in-path function raises an exception if the file can’t be found. The find-in-path-optional function silently removes files that can’t be found.

9.2.7 digest-path

    $(digest-in-path path, files) : String/File Array
       path : Dir Array
       files : String Array
    raises RuntimeException
    $(digest-in-path-optional path, files) : String/File Array

The digest-in-path function searches for the files in a search path and returns the file and digest for each file. Only the tail of the filename is significant. The digest-in-path function raises an exception if the file can’t be found. The digest-in-path-optional function silently removes elements that can’t be found.

9.3 File stats

9.3.1 file-exists, target-exists, target-is-proper

   $(file-exists files) : String
   $(target-exists files) : String
   $(target-is-proper files) : String
       files : File Sequence

The file-exists function checks whether the files listed exist. The target-exists function is similar to the file-exists function. However, it returns true if the file exists or if it can be built by the current project. The target-is-proper returns true only if the file can be generated in the current project.

9.3.2 stat-reset

   $(stat-reset files) : String
       files : File Sequence

OMake uses a stat-cache. The stat-reset function reset the stat information for the given files, forcing the stat information to be recomputed the next time it is requested.

9.3.3 filter-exists, filter-targets, filter-proper-targets

   $(filter-exists files) : File Sequence
   $(filter-targets files) : File Sequence
   $(filter-proper-targets) : File Sequence
      files : File Sequence
   The filter-exists, filter-targets, and filter-proper-targets func-

tions remove files from a list of files.

• filter-exists: the result is the list of files that exist.

• filter-targets: the result is the list of files either exist, or can be built by the current project.

• filter-proper-targets: the result is the list of files that can be built in the current project.

One way to create a simple “clean” rule that removes generated files from the project is by removing all files that can be built in the current project. CAUTION: you should be careful before you do this. The rule removes any file that can potentially be reconstructed. There is no check to make sure that the commands to rebuild the file would actually succeed. Also, note that no file outside the current project will be deleted.

    .PHONY: clean
    clean:
        rm $(filter-proper-targets $(ls R, .))

See the dependencies-proper function to see an alternate method for removing intermediate files.

If you use CVS, you may wish to use the cvs_realclean program that is distributed with omake.

9.3.4 find-targets-in-path, find-targets-in-path-optional

    $(find-targets-in-path path files) : File Array
    $(find-targets-in-path-optional path, files) : File Array
        path : Dir Array
        files : File Sequence

The find-target-in-path function searches for targets in the search path. For each file file in the file list, the path is searched sequentially for a directory dir such that the target dir/file exists. If so, the file dir/file is returned.

For example, suppose you are building a C project, and project contains a subdirectory src/ containing only the files fee.c and foo.c. The following expression evaluates to the files src/fee.o src/foo.o even if the files have not already been built.

    $(find-targets-in-path lib src, fee.o foo.o)
    # Evaluates to
    src/fee.o src/foo.o

The find-targets-in-path function raises an exception if the file can’t be found. The find-targets-in-path-optional function silently removes targets that can’t be found.

    $(find-targets-in-path-optional lib src, fee.o foo.o fum.o)
    # Evaluates to
    src/fee.o src/foo.o

9.3.5 file-sort

   $(file-sort order, files) : File Sequence
      order : String
      files : File Sequence

The file-sort function sorts a list of filenames by build order augmented by a set of sort rules. Sort rules are declared using the .ORDER target. The .BUILDORDER defines the default order.

$(file-sort <order>, <files>)

For example, suppose we have the following set of rules.

a: b c
b: d
c: d
.DEFAULT: a b c d
   echo $(file-sort .BUILDORDER, a b c d)

In the case, the sorter produces the result d b c a. That is, a target is sorted after its dependencies. The sorter is frequently used to sort files that are to be linked by their dependencies (for languages where this matters).

There are three important restrictions to the sorter:

• The sorter can be used only within a rule body. The reason for this is that all dependencies must be known before the sort is performed.

• The sorter can only sort files that are buildable in the current project.

• The sorter will fail if the dependencies are cyclic.

9.3.5.1 sort rule

It is possible to further constrain the sorter through the use of sort rules. A sort rule is declared in two steps. The target must be listed as an .ORDER target; and then a set of sort rules must be given. A sort rule defines a pattern constraint.

.ORDER: .MYORDER
.MYORDER: %.foo: %.bar
.MYORDER: %.bar: %.baz
.DEFAULT: a.foo b.bar c.baz d.baz
   echo $(sort .MYORDER, a.foo b.bar c.baz d.baz)

In this example, the .MYORDER sort rule specifies that any file with a suffix .foo should be placed after any file with suffix .bar, and any file with suffix .bar should be placed after a file with suffix .baz.

In this example, the result of the sort is d.baz c.baz b.bar a.foo.

9.3.6 file-check-sort

file-check-sort(files)
   files : File Sequence
raises RuntimeException

The file-check-sort function checks whether a list of files is in sort order. If so, the list is returned unchanged. If not, the function raises an exception.

$(file-check-sort <order>, <files>)

9.4 Globbing and file listings

OMake commands are “glob-expanded” before being executed. That is, names may contain patterns that are expanded to sequences of file and directory names. The syntax follows the standard bash(1), csh(1), syntax, with the following rules.

• A pathname is a sequence of directory and file names separated by one of the / or \ characters. For example, the following pathnames refer to the

same file: /home/jyh/OMakefile and /home\jyh/OMakefile.

• Glob-expansion is performed on the components of a path. If a path contains occurrences of special characters (listed below), the path is viewed as a pattern to be matched against the actual files in the system. The expansion produces a sequence of all file/directory names that match.

For the following examples, suppose that a directory /dir contains files named a, -a, a.b, and b.c.

* Matches any sequence of zero-or-more characters. For example, the

    pattern /dir/a* expands to /dir/a /dir/aa /dir/a.b.

? Matches exactly one character. The pattern /dir/?a expands the file-

    name /dir/-a.

[...] Square brackets denote character sets and ranges in the ASCII character set. The pattern may contain individual characters c or character ranges c_{1}-c_{2}. The pattern matches any of the individual characters specified, or any characters in the range. A leading “hat” inverts the send of the pattern. To specify a pattern that contains the literal characters -, the - should occur as the first character in the range.

Pattern  Expansion

         /dir/[a-b]*   /dir/a /dir/a.b /dir/b.c
         /dir/[-a-b]*  /dir/a /dir/-a /dir/a.b /dir/b.c
         /dir/[-a]*    /dir/a /dir/-a /dir/a.b

{s1,...,sN} Braces indicate brace-expansion. The braces delimit a sequence of strings separated by commas. Given N strings, the result produces N copies of the pattern, one for each of the strings s_{i}.

Pattern  Expansion

                 a{b,c,d}         ab ac ad
                 a{b{c,d},e}      abc abd ae
                 a{?{[A-Z],d},*}  a?[A-Z] a?d a*

The tilde is used to specify home directories. Depending on your system, these might be possible expansions.

Pattern  Expansion

               ~jyh      /home/jyh
               ~bob/*.c  c:\Documents and Settings\users\bob

The \ character is both a pathname separator and an escape character. If followed by a special glob character, the \ changes the sense of the following character to non-special status. Otherwise, \ is viewed as a pathname separator.

Pattern  Expansion

          ~jyh/\*                 ~jyh/* (* is literal)
          /dir/\[a-z?             /dir/[a-z? ([ is literal, ? is a pattern).
          c:\Program Files\[A-z]  c:\Program Files[A-z]*

Note that the final case might be considered to be ambiguous (where \ should be viewed as a pathname separator, not as an escape for the subsequent [ character. If you want to avoid this ambiguity on Win32, you should use the forward slash / even for Win32 pathnames (the / is translated to \ in the output).

          Pattern  Expansion c:/Program Files/[A-z]*  c:\Program Files\WindowsUpdate ...

9.4.1 glob

   $(glob strings) : Node Array
      strings : String Sequence
   $(glob options, strings) : Node Array
      options : String
      strings : String Sequence

The glob function performs glob-expansion.

The . and .. entries are always ignored.

The options are:

b Do not perform csh(1)-style brace expansion.

e The \ character does not escape special characters.

n If an expansion fails, return the expansion literally instead of aborting.

i If an expansion fails, it expands to nothing.

. Allow wildcard patterns to match files beginning with a .

A Return all files, including files that begin with a .

F Match only normal files (any file that is not a directory).

D Match only directory files.

C Ignore files according to cvs(1) rules.

P Include only proper subdirectories.

In addition, the following variables may be defined that affect the behavior

of glob.

GLOB OPTIONS A string containing default options.

GLOB IGNORE A list of shell patterns for filenames that glob should ignore.

GLOB ALLOW A list of shell patterns. If a file does not match a pattern in GLOB_ALLOW, it is ignored.

The returned files are sorted by name.

9.4.2 ls

   $(ls files) : Node Array
      files : String Sequence
   $(ls options, files) : Node Array
      files : String Sequence

The ls function returns the filenames in a directory.

The . and .. entries are always ignored. The patterns are shell-style patterns, and are glob-expanded.

The options include all of the options to the glob function, plus the following.

R Perform a recursive listing.

The GLOB_ALLOW and GLOB_IGNORE variables can be defined to control the globbing behavior. The returned files are sorted by name.

9.4.3 subdirs

   $(subdirs dirs) : Dir Array
      dirs : String Sequence
   $(subdirs options, dirs) : Dir Array
      options : String
      dirs : String Sequence

The subdirs function returns all the subdirectories of a list of directories, recursively.

The possible options are the following:

A Return directories that begin with a .

C Ignore files according to .cvsignore rules.

P Include only proper subdirectories.

128 CHAPTER 9. FILE, I/O AND SYSTEM OPERATIONS

9.5 Filesystem operations

9.5.1 mkdir

   mkdir(mode, node...)
      mode : Int
      node : Node
   raises RuntimeException
   mkdir(node...)
      node : Node
   raises RuntimeException

The mkdir function creates a directory, or a set of directories. The following options are supported. -m mode Specify the permissions of the created directory. -p Create parent directories if they do not exist.

– Interpret the remaining names literally.

9.5.2 Stat

The Stat object represents an information about a filesystem node, as returned by the stat and lstat functions. It contains the following fields. dev : the device number. ino : the inode number. kind : the kind of the file, one of the following: REG (regular file), DIR (di-

rectory), CHR (character device), BLK (block device), LNK (symbolic link), FIFO (named pipe), SOCK (socket).

perm : access rights, represented as an integer. nlink : number of links. uid : user id of the owner. gid : group id of the file’s group. rdev : device minor number. size : size in bytes. atime : last access time, as a floating point number. mtime : last modification time, as a floating point number. ctime : last status change time, as a floating point number.

Not all of the fields will have meaning on all operating systems.

9.5.3 stat, lstat

    $(stat node...) : Stat
       node : Node or Channel
    $(lstat node...) : Stat
       node : Node or Channel
    raises RuntimeException

The stat functions return file information. If the file is a symbolic link, the stat function refers to the destination of the link; the lstat function refers to the link itself.

9.5.4 unlink

   $(unlink file...)
      file : File
   #(rm file...)
      file : File
   $(rmdir dir...)
      dir : Dir
   raises RuntimeException

The unlink and rm functions remove a file. The rmdir function removes a directory.

The following options are supported for rm and rmdir.

-f ignore nonexistent files, never prompt.

-i prompt before removal.

-r remove the contents of directories recursively.

-v explain what is going on.

– the rest of the values are interpreted literally.

9.5.5 rename

    rename(old, new)
       old : Node
       new : Node
    mv(nodes... dir)
       nodes : Node Sequence
       dir   : Dir
    cp(nodes... dir)
       nodes : Node Sequence
       dir   : Dir
    raises RuntimeException

The rename function changes the name of a file or directory named old to

new.

The mv function is similar, but if new is a directory, and it exists, then the files specified by the sequence are moved into the directory. If not, the behavior of mv is identical to rename. The cp function is similar, but the original file is not removed.

The mv and cp functions take the following options.

-f Do not prompt before overwriting.

-i Prompt before overwriting.

-v Explain what it happening.

-r Copy the contents of directories recursively.

– Interpret the remaining arguments literally.

9.5.6 link

   link(src, dst)
      src : Node
      dst : Node
   raises RuntimeException

The link function creates a hard link named dst to the file or directory src. Hard links may work under Win32 when NTFS is used. Normally, only the superuser can create hard links to directories.

9.5.7 symlink

   symlink(src, dst)
      src : Node
      dst : Node
   raises RuntimeException

The symlink function creates a symbolic link dst that points to the src file. The link name is computed relative to the target directory. For example,

the expression $(symlink a/b, c/d) creates a link named c/d -> ../a/b.

Symbolic links are not supported in Win32. Consider using the ln-or-cp Shell alias for cross-platform portable linking/copying.

9.5.8 readlink

   $(readlink node...) : Node
      node : Node

The readlink function reads the value of a symbolic link.

9.5.9 chmod

chmod(mode, dst...)
mode : Int
dst : Node or Channel
chmod(mode dst...)
mode : String
dst : Node Sequence
raises RuntimeException

The chmod function changes the permissions of the targets. Options:

-v Explain what is happening.

-r Change files and directories recursively.

-f Continue on errors.

– Interpret the remaining argument literally.

9.5.10 chown

chown(uid, gid, node...)
uid : Int
gid : Int
node : Node or Channel
chown(uid, node...)
uid : Int
node : Node or Channel
raises RuntimeException

The chown function changes the user and group id of the file. If the gid is not specified, it is not changed. If either id is -1, that id is not changed.

9.5.11 truncate

truncate(length, node...)
 length : Int
 node : Node or Channel
raises RuntimeException

The truncate function truncates a file to the given length.

9.5.12 umask

$(umask mode) : Int
 mode : Int
raises RuntimeException

Sets the file mode creation mask. The previous mask is returned. This value is not scoped, changes have global effect.

9.6 vmount

9.6.1 vmount

    vmount(src, dst)
       src, dst : Dir
    vmount(flags, src, dst)
       flags : String
       src, dst : Dir

“Mount” the src directory on the dst directory. This is a virtual mount, changing the behavior of the $(file ...) function. When the $(file str) function is used, the resulting file is taken relative to the src directory if the file exists. Otherwise, the file is relative to the current directory.

The main purpose of the vmount function is to support multiple builds with separate configurations or architectures.

The options are as follows.

l Create symbolic links to files in the src directory.

c Copy files from the src directory.

Mount operations are scoped.

9.6.2 add-project-directories

    add-project-directories(dirs)
       dirs : Dir Array

Add the directories to the set of directories that omake considers to be part of the project. This is mainly used to avoid omake complaining that the current directory is not part of the project.

9.6.3 remove-project-directories

    remove-project-directories(dirs)
       dirs : Dir Array

Removed the directories from the set of directories that omake considers to be part of the project. This is mainly used to cancel a .SUBDIRS from including a directory if it is determined that the directory does not need to be compiled.

9.7. FILE PREDICATES 133

9.7 File predicates

9.7.1 test

test(exp) : Bool
   exp : String Sequence

The expression grammar is as follows:

• ! expression : expression is not true

• expression1 -a expression2 : both expressions are true

• expression1 -o expression2 : at least one expression is true

• ( expression ) : expression is true The base expressions are:

• -n string : The string has nonzero length

• -z string : The string has zero length

• string = string : The strings are equal

• string != string : The strings are not equal

• int1 -eq int2 : The integers are equal

• int1 -ne int2 : The integers are not equal

• int1 -gt int2 : int1 is larger than int2

• int1 -ge int2 : int2 is not larger than int1

• int1 -lt int2 : int1 is smaller than int2

• int1 -le int2 : int1 is not larger than int2

• file1 -ef file2 : On Unix, file1 and file2 have the same device and inode number. On Win32, file1 and file2 have the same name.

• file1 -nt file2 : file1 is newer than file2

• file1 -ot file2 : file1 is older than file2

• -b file : The file is a block special file

• -c file : The file is a character special file

• -d file : The file is a directory

• -e file : The file exists • -f file : The file is a normal file

• -g file : The set-group-id bit is set on the file

• -G file : The file’s group is the current effective group

• -h file : The file is a symbolic link (also -L)

• -k file : The file’s sticky bit is set

• -L file : The file is a symbolic link (also -h)

• -O file : The file’s owner is the current effective user

• -p file : The file is a named pipe

• -r file : The file is readable

• -s file : The file is empty

• -S file : The file is a socket

• -u file : The set-user-id bit is set on the file

• -w file : The file is writable

• -x file : The file is executable

A string is any sequence of characters; leading - characters are allowed. An int is a string that can be interpreted as an integer. Unlike traditional

versions of the test program, the leading characters may specify an arity. The prefix 0b means the numbers is in binary; the prefix 0o means the number is in octal; the prefix 0x means the number is in hexadecimal. An int can also be specified as -l string, which evaluates to the length of the string.

A file is a string that represents the name of a file. The syntax mirrors that of the test(1) program. If you are on a Unix

system, the man page explains more. Here are some examples.

 # Create an empty file
 osh> touch foo
 # Is the file empty?
 osh> test(-e foo)
 - : true
 osh> test(! -e foo)
 - : false
 # Create another file
 osh> touch boo
 # Is the newer file newer?
 osh> test(boo -nt foo)
 - : true
 # A more complex query

# boo is newer than foo, and foo is empty
osh> test(\( boo -nt foo \) -a -e foo)
- : true

9.7.2 find

find(exp) : Node Array
exp : String Sequence

The find function searches a directory recursively, returning the files for which the expression evaluates to true.

The expression argument uses the same syntax as the test function, with the following exceptions.

1. The expression may begin with a directory. If not specified, the current directory is searched.

2. The {} string expands to the current file being examined.

The syntax of the expression is the same as test, with the following additions.

• -name string : The current file matches the glob expression (see Section 9.4).

The find function performs a recursive scan of all subdirectories. The following call is being run from the root of the omake source directory.

osh> find(. -name fo* )
- : <array
/home/jyh/.../omake/mk/.svn/format
/home/jyh/.../omake/RPM/.svn/format
...
/home/jyh/.../omake/osx_resources/installer_files/.svn/format>

Another example, listing only those files that are normal files or symbolic links.

osh> find(. -name fo* -a \( -f {} -o -L {} \))
- : <array
/home/jyh/.../omake/mk/.svn/format
/home/jyh/.../omake/RPM/.svn/format
...
/home/jyh/.../omake/osx_resources/installer_files/.svn/format>

9.8 IO functions

9.8.1 Standard channels

The following variables define the standard channels.

136 CHAPTER 9. FILE, I/O AND SYSTEM OPERATIONS

stdin

stdin : InChannel

The standard input channel, open for reading. stdout

stdout : OutChannel

The standard output channel, open for writing. stderr

stderr : OutChannel

The standard error channel, open for writing.

9.8.2 open-in-string

The open-in-string treats a string as if it were a file and returns a channel for reading.

   $(open-in-string s) : Channel
       s : String

9.8.3 open-out-string, out-string

The open-out-string creates a channel that writes to a string instead of a file. The string may be retrieved with the out-string function.

   $(open-out-string) : Channel
   $(out-string chan) : String
       chan : OutChannel

9.8.4 fopen

The fopen function opens a file for reading or writing.

   $(fopen file, mode) : Channel
      file : File
      mode : String

The file is the name of the file to be opened. The mode is a combination of the following characters. r Open the file for reading; it is an error if the file does not exist. w Open the file for writing; the file is created if it does not exist. a Open the file in append mode; the file is created if it does not exist.

+ Open the file for both reading an writing.

t Open the file in text mode (default).

b Open the file in binary mode.

n Open the file in nonblocking mode.

x Fail if the file already exists.

Binary mode is not significant on Unix systems, where text and binary modes are equivalent.

9.8.5 close

    $(close channel...)
       channel : Channel

The close function closes a file that was previously opened with fopen.

9.8.6 read

   $(read channel, amount) : String
      channel : InChannel
      amount  : Int
   raises RuntimeException

The read function reads up to amount bytes from an input channel, and returns the data that was read. If an end-of-file condition is reached, the function raises a RuntimeException exception.

9.8.7 write

   $(write channel, buffer, offset, amount) : String
      channel : OutChannel
      buffer  : String
      offset  : Int
      amount  : Int
   $(write channel, buffer) : String
      channel : OutChannel
      buffer  : String
   raises RuntimeException

In the 4-argument form, the write function writes bytes to the output channel channel from the buffer, starting at position offset. Up to amount bytes are written. The function returns the number of bytes that were written.

The 3-argument form is similar, but the offset is 0.

In the 2-argument form, the offset is 0, and the amount if the length of the

buffer.

If an end-of-file condition is reached, the function raises a RuntimeException exception.

9.8.8 lseek

    $(lseek channel, offset, whence) : Int
       channel : Channel
       offset  : Int
       whence  : String
    raises RuntimeException

The lseek function repositions the offset of the channel channel according to the whence directive, as follows:

SEEK SET The offset is set to offset.

SEEK CUR The offset is set to its current position plus offset bytes.

SEEK END The offset is set to the size of the file plus offset bytes.

The lseek function returns the new position in the file.

9.8.9 rewind

   rewind(channel...)
      channel : Channel

The rewind function set the current file position to the beginning of the file.

9.8.10 tell

    $(tell channel...) : Int...
       channel : Channel
    raises RuntimeException

The tell function returns the current position of the channel.

9.8.11 flush

   $(flush channel...)
      channel : OutChannel

The flush function can be used only on files that are open for writing. It flushes all pending data to the file.

9.8.12 dup

    $(dup channel) : Channel
       channel : Channel
    raises RuntimeException

The dup function returns a new channel referencing the same file as the argument.

9.8.13 dup2

   dup2(channel1, channel2)
      channel1 : Channel
      channel2 : Channel
   raises RuntimeException

The dup2 function causes channel2 to refer to the same file as channel1.

9.8.14 set-nonblock

   set-nonblock-mode(mode, channel...)
      channel : Channel
      mode : String

The set-nonblock-mode function sets the nonblocking flag on the given channel. When IO is performed on the channel, and the operation cannot be completed immediately, the operations raises a RuntimeException.

9.8.15 set-close-on-exec-mode

   set-close-on-exec-mode(mode, channel...)
      channel : Channel
      mode : String
   raises RuntimeException

The set-close-on-exec-mode function sets the close-on-exec flags for the given channels. If the close-on-exec flag is set, the channel is not inherited by child processes. Otherwise it is.

9.8.16 pipe

   $(pipe) : Pipe
   raises RuntimeException

The pipe function creates a Pipe object, which has two fields. The read field is a channel that is opened for reading, and the write field is a channel that is opened for writing.

9.8.17 mkfifo

   mkfifo(mode, node...)
      mode : Int
      node : Node

The mkfifo function creates a named pipe.

9.8.18 select

   $(select rfd..., wfd..., wfd..., timeout) : Select
      rfd : InChannel
      wfd : OutChannel
      efd : Channel
      timeout : float
   raises RuntimeException

The select function polls for possible IO on a set of channels. The rfd are a sequence of channels for reading, wfd are a sequence of channels for writing, and efd are a sequence of channels to poll for error conditions. The timeout specifies the maximum amount of time to wait for events.

On successful return, select returns a Select object, which has the following fields:

read An array of channels available for reading.

write An array of channels available for writing.

error An array of channels on which an error has occurred.

9.8.19 lockf

    lockf(channel, command, len)
       channel : Channel
       command : String
       len : Int
    raises RuntimeException

The lockf function places a lock on a region of the channel. The region starts at the current position and extends for len bytes.

The possible values for command are the following.

F ULOCK Unlock a region.

F LOCK Lock a region for writing; block if already locked.

F TLOCK Lock a region for writing; fail if already locked.

F TEST Test a region for other locks.

F RLOCK Lock a region for reading; block if already locked.

F TRLOCK Lock a region for reading; fail is already locked.

9.8.20 InetAddr

The InetAddr object describes an Internet address. It contains the following fields.

addr String: the Internet address.

port Int: the port number.

9.8.21 Host

A Host object contains the following fields.

name String: the name of the host.

aliases String Array: other names by which the host is known.

addrtype String: the preferred socket domain.

addrs InetAddr Array: an array of Internet addresses belonging to the host.

9.8.22 gethostbyname

   $(gethostbyname host...) : Host...
      host : String
   raises RuntimeException

The gethostbyname function returns a Host object for the specified host. The host may specify a domain name or an Internet address.

9.8.23 Protocol

The Protocol object represents a protocol entry. It has the following fields.

name String: the canonical name of the protocol.

aliases String Array: aliases for the protocol.

proto Int: the protocol number.

9.8.24 getprotobyname

   $(getprotobyname name...) : Protocol...
      name : Int or String
   raises RuntimeException

The getprotobyname function returns a Protocol object for the specified protocol. The name may be a protocol name, or a protocol number.

9.8.25 Service

The Service object represents a network service. It has the following fields.

name String: the name of the service.

aliases String Array: aliases for the service.

port Int: the port number of the service.

proto Protocol: the protocol for the service.

9.8.26 getservbyname

   $(getservbyname service...) : Service...
      service : String or Int
   raises RuntimeException

The getservbyname function gets the information for a network service. The service may be specified as a service name or number.

9.8.27 socket

   $(socket domain, type, protocol) : Channel
      domain : String
      type : String
      protocol : String
   raises RuntimeException

The socket function creates an unbound socket.

The possible values for the arguments are as follows.

The domain may have the following values.

PF UNIX or unix Unix domain, available only on Unix systems.

PF INET or inet Internet domain, IPv4.

PF INET6 or inet6 Internet domain, IPv6.

The type may have the following values.

SOCK STREAM or stream Stream socket.

SOCK DGRAM or dgram Datagram socket.

SOCK RAW or raw Raw socket.

SOCK SEQPACKET or seqpacket Sequenced packets socket

The protocol is an Int or String that specifies a protocol in the protocols database.

9.8.28 bind

bind(socket, host, port)
socket : InOutChannel
host : String
port : Int
bind(socket, file)
socket : InOutChannel
file : File
raise RuntimeException

The bind function binds a socket to an address. The 3-argument form specifies an Internet connection, the host specifies a

host name or IP address, and the port is a port number.

The 2-argument form is for Unix sockets. The file specifies the filename for the address.

9.8.29 listen

listen(socket, requests)
socket : InOutChannel
requests : Int
raises RuntimeException

The listen function sets up the socket for receiving up to requests number of pending connection requests.

9.8.30 accept

$(accept socket) : InOutChannel
socket : InOutChannel
raises RuntimeException

The accept function accepts a connection on a socket.

9.8.31 connect

connect(socket, addr, port)
 socket : InOutChannel
 addr : String
 port : int
connect(socket, name)
 socket : InOutChannel
 name : File
raise RuntimeException

The connect function connects a socket to a remote address.

The 3-argument form specifies an Internet connection. The addr argument is the Internet address of the remote host, specified as a domain name or IP address. The port argument is the port number.

The 2-argument form is for Unix sockets. The name argument is the filename of the socket.

9.8.32 getchar

    $(getc) : String
    $(getc file) : String
       file : InChannel or File
    raises RuntimeException

The getc function returns the next character of a file. If the argument is not specified, stdin is used as input. If the end of file has been reached, the function returns false.

9.8.33 gets

   $(gets) : String
   $(gets channel) : String
      channel : InChannel or File
   raises RuntimeException

The gets function returns the next line from a file. The function returns the empty string if the end of file has been reached. The line terminator is removed.

9.8.34 fgets

   $(fgets) : String
   $(fgets channel) : String
      channel : InChannel or File
   raises RuntimeException

The fgets function returns the next line from a file that has been opened for reading with fopen. The function returns the empty string if the end of file has been reached. The returned string is returned as literal data. The line terminator is not removed.

9.9 Printing functions

Output is printed with the print and println functions. The println function adds a terminating newline to the value being printed, the print function does not.

fprint(<file>, <string>)
print(<string>)
eprint(<string>)
fprintln(<file>, <string>)
println(<string>)
eprintln(<string>)

The fprint functions print to a file that has been previously opened with fopen. The print functions print to the standard output channel, and the eprint functions print to the standard error channel.

9.10 Value printing functions

Values can be printed with the printv and printvln functions. The printvln function adds a terminating newline to the value being printed, the printv function does not.

fprintv(<file>, <string>)
printv(<string>)
eprintv(<string>)
fprintvln(<file>, <string>)
printvln(<string>)
eprintvln(<string>)

The fprintv functions print to a file that has been previously opened with fopen. The printv functions print to the standard output channel, and the eprintv functions print to the standard error channel.

9.11 Higher-level IO functions

9.11.1 Regular expressions

Many of the higher-level functions use regular expressions. Regular expressions are defined by strings with syntax nearly identical to awk(1).

Strings may contain the following character constants.

• \\ : a literal backslash.

• \a : the alert character ^G.

• \b : the backspace character ^H.

• \f : the formfeed character ^L.

• \n : the newline character ^J.

• \r : the carriage return character ^M.

146 CHAPTER 9. FILE, I/O AND SYSTEM OPERATIONS

• \t : the tab character ^I.

• \v : the vertical tab character.

• \xhh... : the character represented by the string of hexadecimal digits h. All valid hexadecimal digits following the sequence are considered to be part of the sequence.

• \ddd : the character represented by 1, 2, or 3 octal digits. Regular expressions are defined using the special characters .\^$[(){}*?+.

• c : matches the literal character c if c is not a special character.

• \c : matches the literal character c, even if c is a special character.

• . : matches any character, including newline.

• ^: matches the beginning of a line.

• $ : matches the end of line.

• [abc...] : matches any of the characters abc...

• [^abc...] : matches any character except abc...

• r1|r2 : matches either r1 or r2.

• r1r2 : matches r1 and then r2.

• r+ : matches one or more occurrences of r.

• r* : matches zero or more occurrences of r.

• r? : matches zero or one occurrence of r.

• (r) : parentheses are used for grouping; matches r.

• \(r\) : also defines grouping, but the expression matched within the parentheses is available to the output processor through one of the variables $1, $2, ...

• r{n} : matches exactly n occurrences of r.

• r{n,} : matches n or more occurrences of r.

• r{n,m} : matches at least n occurrences of r, and no more than m occurrences.

• \y: matches the empty string at either the beginning or end of a word.

• \B: matches the empty string within a word.

• \<: matches the empty string at the beginning of a word.

9.11. HIGHER-LEVEL IO FUNCTIONS 147 • \>: matches the empty string at the end of a word.

• \w: matches any character in a word.

• \W: matches any character that does not occur within a word.

• \`: matches the empty string at the beginning of a file.

• \': matches the empty string at the end of a file. Character classes can be used to specify character sequences abstractly.

Some of these sequences can change depending on your LOCALE.

• [:alnum:] Alphanumeric characters.

• [:alpha:] Alphabetic characters.

• [:lower:] Lowercase alphabetic characters.

• [:upper:] Uppercase alphabetic characters.

• [:cntrl:] Control characters.

• [:digit:] Numeric characters.

• [:xdigit:] Numeric and hexadecimal characters.

• [:graph:] Characters that are printable and visible.

• [:print:] Characters that are printable, whether they are visible or not.

• [:punct:] Punctuation characters.

• [:blank:] Space or tab characters.

• [:space:] Whitespace characters.

9.11.2 cat

 cat(files) : Sequence
    files : File or InChannel Sequence

The cat function concatenates the output from multiple files and returns it as a string.

9.11.3 grep

   grep(pattern) : String  # input from stdin, default options
      pattern : String
   grep(pattern, files) : String  # default options
      pattern : String
      files   : File Sequence
   grep(options, pattern, files) : String
     options : String
     pattern : String
     files   : File Sequence

The grep function searches for occurrences of a regular expression pattern in a set of files, and prints lines that match. This is like a highly-simplified version of grep(1).

The options are:

q If specified, the output from grep is not displayed.

h If specified, output lines will not include the filename (default, when only one input file is given).

n If specified, output lines include the filename (default, when more than one input file is given).

v If specified, search for lines without a match instead of lines with a match,

The pattern is a regular expression.

If successful (grep found a match), the function returns true. Otherwise, it returns false.

9.11.4 scan

   scan(input-files)
   case string1
      body1
   case string2
      body2
   ...
   default
      bodyd

The scan function provides input processing in command-line form. The function takes file/filename arguments. If called with no arguments, the input is taken from stdin. If arguments are provided, each specifies an InChannel, or the name of a file for input. Output is always to stdout.

The scan function operates by reading the input one line at a time, and processing it according to the following algorithm.

For each line, the record is first split into fields, and the fields are bound to the variables $1, $2, .... The variable $0 is defined to be the entire line, and $* is an array of all the field values. The $(NF) variable is defined to be the number of fields.

Next, a case expression is selected. If string_i matches the token $1, then body_i is evaluated. If the body ends in an export, the state is passed to the next clause. Otherwise the value is discarded.

For example, here is an scan function that acts as a simple command processor.

    calc() =
       i = 0
       scan(script.in)
       case print
          println($i)
       case inc
          i = $(add $i, 1)
          export
       case dec
          i = $(sub $i, 1)
          export
       case addconst
          i = $(add $i, $2)
          export
       default
          eprintln($"Unknown command: $1")

The scan function also supports several options.

    scan(options, files)
    ...

A Parse each line as an argument list, where arguments may be quoted. For example, the following line has three words, “ls”, “-l”, “Program Files”.

            ls -l "Program Files"

O Parse each line using white space as the separator, using the usual OMake algorithm for string parsing. This is the default.

x Once each line is split, reduce each word using the hex representation. This is the usual hex representation used in URL specifiers, so the string “Program Files” may be alternately represented in the form Program-Program+Files.

Note, if you want to redirect the output to a file, the easiest way is to redefine the stdout variable. The stdout variable is scoped the same way as other variables, so this definition does not affect the meaning of stdout outside the calc function.

    calc() =
        stdout = $(fopen script.out, w)
        scan(script.in)
           ...
        close(stdout)

9.11.5 awk

   awk(input-files)
   case pattern1:
      body1
   case pattern2:
      body2
   ...
   default:
      bodyd

or

   awk(options, input-files)
   case pattern1:
      body1
   case pattern2:
      body2
   ...
   default:
      bodyd

The awk function provides input processing similar to awk(1), but more limited. The input-files argument is a sequence of values, each specifies an InChannel, or the name of a file for input. If called with no options and no file arguments, the input is taken from stdin. Output is always to stdout.

The variables RS and FS define record and field separators as regular expressions. The default value of RS is the regular expression \r|\n|\r\n. The default value of FS is the regular expression [ \t]+.

The awk function operates by reading the input one record at a time, and processing it according to the following algorithm.

For each line, the record is first split into fields using the field separator FS, and the fields are bound to the variables $1, $2, .... The variable $0 is defined to be the entire line, and $* is an array of all the field values. The $(NF) variable is defined to be the number of fields.

Next, the cases are evaluated in order. For each case, if the regular expression pattern_i matches the record $0, then body_i is evaluated. If the body ends in an export, the state is passed to the next clause. Otherwise the value is discarded. If the regular expression contains \(r\) expression, those expression override the fields $1, $2, ....

For example, here is an awk function to print the text between two delimiters

\begin{<name>} and \end{<name>}, where the <name> must belong to a set

passed as an argument to the filter function.

    filter(names) =
       print = false
       awk(Awk.in)
       case $"^\\end\{\([:alpha:]+\)\}"
          if $(mem $1, $(names))
             print = false
             export
          export
       default
          if $(print)
             println($0)
       case $"^\\begin\{\([:alpha:]+\)\}"
          print = $(mem $1, $(names))
          export

Note, if you want to redirect the output to a file, the easiest way is to redefine the stdout variable. The stdout variable is scoped the same way as other variables, so this definition does not affect the meaning of stdout outside the filter function.

    filter(names) =
        stdout = $(fopen file.out, w)
        awk(Awk.in)
           ...
        close(stdout)

Options.

b “Break” when evaluating cases. Only the first case that matches will be selected.

The break function can be used to abort the loop, exiting the awk function immediately.

9.11.6 fsubst

   fsubst(files)
   case pattern1 [options]

      body1
   case pattern2 [options]
      body2
   ...
   default
      bodyd

The fsubst function provides a sed(1)-like substitution function. Similar to awk, if fsubst is called with no arguments, the input is taken from stdin. If arguments are provided, each specifies an InChannel, or the name of a file for input.

The RS variable defines a regular expression that determines a record separator, The default value of RS is the regular expression \r|\n|\r\n.

The fsubst function reads the file one record at a time.

For each record, the cases are evaluated in order. Each case defines a substitution from a substring matching the pattern to replacement text defined by the body.

Currently, there is only one option: g. If specified, each clause specifies a global replacement, and all instances of the pattern define a substitution. Otherwise, the substitution is applied only once.

Output can be redirected by redefining the stdout variable.

For example, the following program replaces all occurrences of an expression word. with its capitalized form.

    section
       stdout = $(fopen Subst.out, w)
       fsubst(Subst.in)
       case $"\<\([[:alnum:]]+\)\." g
          value $(capitalize $1).
       close(stdout)

9.11.7 lex

   lex(files)
   case pattern1
      body1
   case pattern2
      body2
   ...
   default
      bodyd

The lex function provides a simple lexical-style scanner function. The input is a sequence of files or channels. The cases specify regular expressions. Each time the input is read, the regular expression that matches the longest prefix of the input is selected, and the body is evaluated.

If two clauses both match the same input, the last one is selected for execution. The default case matches the regular expression .; you probably want to place it first in the pattern list.

If the body end with an export directive, the state is passed to the next clause.

For example, the following program collects all occurrences of alphanumeric words in an input file.

    collect-words($(files)) =
       words[] =
       lex($(files))
       default
          # empty
       case $"[[:alnum:]]+" g
          words[] += $0
          export

The default case, if one exists, matches single characters. Since It is an error if the input does not match any of the regular expressions. The break function can be used to abort the loop.

9.11.8 lex-search

   lex-search(files)
   case pattern1
      body1
   case pattern2
      body2
   ...
   default
      bodyd

The lex-search function is like the lex function, but input that does not match any of the regular expressions is skipped. If the clauses include a default case, then the default matches any skipped text.

For example, the following program collects all occurrences of alphanumeric words in an input file, skipping any other text.

    collect-words($(files)) =
       words[] =
       lex-search($(files))
       default
          eprintln(Skipped $0)
       case $"[[:alnum:]]+" g
          words[] += $0
          export

The default case, if one exists, matches single characters. Since It is an error if the input does not match any of the regular expressions. The break function can be used to abort the loop.

9.11.9 Lexer

The Lexer object defines a facility for lexical analysis, similar to the lex(1) and flex(1) programs.

In omake, lexical analyzers can be constructed dynamically by extending the Lexer class. A lexer definition consists of a set of directives specified with method calls, and set of clauses specified as rules.

For example, consider the following lexer definition, which is intended for lexical analysis of simple arithmetic expressions for a desktop calculator.

lexer1. =
extends $(Lexer)
other: .
eprintln(Illegal character: $* )
lex()
white: $"[[:space:]]+"
lex()
op: $"[-+*/()]"
switch $*
case +
   Token.unit($(loc), plus)
case -
   Token.unit($(loc), minus)
case *
   Token.unit($(loc), mul)
case /
   Token.unit($(loc), div)
case $"("
   Token.unit($(loc), lparen)
case $")"
   Token.unit($(loc), rparen)
number: $"[[:digit:]]+"
Token.pair($(loc), exp, $(int $* ))
eof: $"\'"
Token.unit($(loc), eof)

This program defines an object lexer1 the extends the Lexer object, which defines lexing environment.

The remainder of the definition consists of a set of clauses, each with a method name before the colon; a regular expression after the colon; and in this case, a body. The body is optional, if it is not specified, the method with the given name should already exist in the lexer definition.

NB The clause that matches the longest prefix of the input is selected. If two clauses match the same input prefix, then the last one is selected. This is unlike most standard lexers, but makes more sense for extensible grammars.

The first clause matches any input that is not matched by the other clauses. In this case, an error message is printed for any unknown character, and the input is skipped. Note that this clause is selected only if no other clause matches.

The second clause is responsible for ignoring white space. If whitespace is found, it is ignored, and the lexer is called recursively.

The third clause is responsible for the arithmetic operators. It makes use of the Token object, which defines three fields: a loc field that represents the source location; a name; and a value.

The lexer defines the loc variable to be the location of the current lexeme in each of the method bodies, so we can use that value to create the tokens.

The Token.unit($(loc), name) method constructs a new Token object with the given name, and a default value.

The number clause matches nonnegative integer constants. The Token.pair($(loc), name, value) constructs a token with the given name and value.

Lexer object operate on InChannel objects. The method lexer1.lex-channel(channel) reads the next token from the channel argument.

9.11.10 Lexer matching

During lexical analysis, clauses are selected by longest match. That is, the clause that matches the longest sequence of input characters is chosen for evaluation. If no clause matches, the lexer raises a RuntimeException. If more than one clause matches the same amount of input, the first one is chosen for evaluation.

9.11.11 Extending lexer definitions

Suppose we wish to augment the lexer example so that it ignores comments. We will define comments as any text that begins with the string (*, ends with *), and comments may be nested.

One convenient way to do this is to define a separate lexer just to skip comments.

   lex-comment. =
      extends $(Lexer)
      level = 0
      other: .
         lex()

      term: $"[*][)]"
         if $(not $(eq $(level), 0))
            level = $(sub $(level), 1)
            lex()
      next: $"[(][*]"
         level = $(add $(level), 1)
         lex()
      eof: $"\'"
         eprintln(Unterminated comment)

This lexer contains a field level that keeps track of the nesting level. On encountering a (* string, it increments the level, and for *), it decrements the level if nonzero, and continues.

Next, we need to modify our previous lexer to skip comments. We can do this by extending the lexer object lexer1 that we just created.

   lexer1. +=
      comment: $"[(][*]"
         lex-comment.lex-channel($(channel))
         lex()

The body for the comment clause calls the lex-comment lexer when a comment is encountered, and continues lexing when that lexer returns.

9.11.12 Threading the lexer object

Clause bodies may also end with an export directive. In this case the lexer object itself is used as the returned token. If used with the Parser object below, the lexer should define the loc, name and value fields in each export clause. Each time the Parser calls the lexer, it calls it with the lexer returned from the previous lex invocation.

9.11.13 Parser

The Parser object provides a facility for syntactic analysis based on context-free grammars.

Parser objects are specified as a sequence of directives, specified with method calls; and productions, specified as rules.

For example, let’s finish building the desktop calculator started in the Lexer example.

   parser1. =
      extends $(Parser)

#
# Use the main lexer
#
lexer = $(lexer1)
#
# Precedences, in ascending order
#
left(plus minus)
left(mul div)
right(uminus)
#
# A program
#
start(prog)
prog: exp eof
   return $1
#
# Simple arithmetic expressions
#
exp: minus exp :prec: uminus
   neg($2)
exp: exp plus exp
   add($1, $3)
exp: exp minus exp
   sub($1, $3)
exp: exp mul exp
   mul($1, $3)
exp: exp div exp
   div($1, $3)
exp: lparen exp rparen
   return $2

Parsers are defined as extensions of the Parser class. A Parser object must have a lexer field. The lexer is not required to be a Lexer object, but it must provide a lexer.lex() method that returns a token object with name and value fields. For this example, we use the lexer1 object that we defined previously.

The next step is to define precedences for the terminal symbols. The precedences are defined with the left, right, and nonassoc methods in order of increasing precedence.

The grammar must have at least one start symbol, declared with the start method.

Next, the productions in the grammar are listed as rules. The name of the production is listed before the colon, and a sequence of variables is listed to the right of the colon. The body is a semantic action to be evaluated when the production is recognized as part of the input.

In this example, these are the productions for the arithmetic expressions recognized by the desktop calculator. The semantic action performs the calculation. The variables $1, $2, ... correspond to the values associated with each of the variables on the right-hand-side of the production.

9.11.14 Calling the parser

The parser is called with the $(parser1.parse-channel start, channel) or
$(parser1.parse-file start, file) functions. The start argument is the

start symbol, and the channel or file is the input to the parser.

9.11.15 Parsing control

The parser generator generates a pushdown automation based on LALR(1) tables. As usual, if the grammar is ambiguous, this may generate shift/reduce or reduce/reduce conflicts. These conflicts are printed to standard output when the automaton is generated.

By default, the automaton is not constructed until the parser is first used.

The build(debug) method forces the construction of the automaton. While not required, it is wise to finish each complete parser with a call to the build(debug) method. If the debug variable is set, this also prints with parser table together with any conflicts.

The loc variable is defined within action bodies, and represents the input range for all tokens on the right-hand-side of the production.

9.11.16 Extending parsers

Parsers may also be extended by inheritance. For example, let’s extend the grammar so that it also recognizes the << and >> shift operations.

First, we extend the lexer so that it recognizes these tokens. This time, we choose to leave lexer1 intact, instead of using the += operator.

   lexer2. =
      extends $(lexer1)
      lsl: $"<<"
         Token.unit($(loc), lsl)

     asr: $">>"
        Token.unit($(loc), asr)

Next, we extend the parser to handle these new operators. We intend that the bitwise operators have lower precedence than the other arithmetic operators. The two-argument form of the left method accomplishes this.

  parser2. =
     extends $(parser1)
     left(plus, lsl lsr asr)
     lexer = $(lexer2)
     exp: exp lsl exp
        lsl($1, $3)
     exp: exp asr exp
        asr($1, $3)

In this case, we use the new lexer lexer2, and we add productions for the new shift operations.

9.11.17 Passwd

The Passwd object represents an entry in the system’s user database. It contains the following fields.

pw name: the login name.

pw passwd: the encrypted password.

pw uid: user id of the user.

pw gid: group id of the user.

pw gecos: the user name or comment field.

pw dir: the user’s home directory.

pw shell: the user’s default shell.

Not all the fields will have meaning on all operating systems.

9.11.18 getpwnam, getpwuid

    $(getpwnam name...) : Passwd
       name : String
    $(getpwuid uid...) : Passwd
       uid : Int
    raises RuntimeException

The getpwnam function looks up an entry by the user’s login and the getpwuid function looks up an entry by user’s numerical id (uid). If no entry is found, an exception will be raised.

9.11.19 getpwents

    $(getpwents) : Array

The getpwents function returns an array of Passwd objects, one for every user fund in the system user database. Note that depending on the operating system and on the setup of the user database, the returned array may be incomplete or even empty.

9.11.20 Group

The Group object represents an entry in the system’s user group database. It contains the following fields.

gr name: the group name.

gr group: the encrypted password.

gr gid: group id of the group.

gr mem: the group member’s user names.

Not all the fields will have meaning on all operating systems.

9.11.21 getgrnam, getgrgid

    $(getgrnam name...) : Group
       name : String
    $(getgrgid gid...) : Group
       gid : Int
    raises RuntimeException

The getgrnam function looks up a group entry by the group’s name and the getgrgid function looks up an entry by groups’s numerical id (gid). If no entry is found, an exception will be raised.

9.11.22 tgetstr

   $(tgetstr id) : String
      id : String

The tgetstr function looks up the terminal capability with the indicated id. This assumes the terminfo to lookup is given in the TERM environment variable. This function returns an empty value if the given terminal capability is not defined.

Note: if you intend to use the value returned by tgetstr inside the shell prompt, you need to wrap it using the prompt-invisible function.

9.11.23 xterm-escape-begin, xterm-escape-end

   $(xterm-escape-begin) : String
   $(xterm-escape-end) : String
   The xterm-escape-begin and xterm-escape-end functions return the es-

cape sequences that can be used to set the XTerm window title. Will return empty values if this capability is not available.

Note: if you intend to use these strings inside the shell prompt, you need to

use $(prompt_invisible_begin)$(xterm-escape-begin) and $(xterm-escape-end)$(prompt_invisible_end).

9.11.24 xterm-escape

   $(xterm-escape s) : Sequence

When the TERM environment variable indicates that the XTerm title setting

capability is available, $(xterm-escape s) is equivalent to $(xterm-escape-begin)s$(xterm-escape-end).

Otherwise, it returns an empty value.

Note: if you intend to use the value returned by xterm-escape inside the shell prompt, you need to wrap it using the prompt-invisible function.

9.11.25 prompt-invisible-begin, prompt-invisible-end

   $(prompt-invisible-begin) : String
   $(prompt-invisible-end) : String
   The prompt-invisible-begin and prompt-invisible-end functions re-

turn the escape sequences that must used to mark the “invisible” sections of the shell prompt (such as various escape sequences).

9.11.26 prompt-invisible

   $(prompt-invisible s) : Sequence
   The prompt-invisible will wrap its argument with $(prompt-invisible-begin)

and $(prompt-invisible-end). All the ‘invisible” sections of the shell prompt (such as various escape sequences) must be wrapped this way.

162 CHAPTER 9. FILE, I/O AND SYSTEM OPERATIONS

9.11.27 gettimeofday

   $(gettimeofday) : Float

The gettimeofday function returns the time of day in seconds since January 1, 1970.

Shell commands

Shell commands (commands to be executed by the operating system) can be freely mixed with other code.

NOTE: the syntax and shell usage is identical on all platforms, including Win32. To avoid portability problems on Win32, it is recommended that you avoid the use of the native shell interpreter cmd.

LIB = $(dir lib)
println(The contents of the $(LIB) directory is:)
ls $(LIB)

10.1 Simple commands

The syntax of shell commands is similar to the syntax used by the Unix shell bash. In general, a command is a pipeline. A basic command is part of a pipeline. It is specified with the name of an executable and some arguments. Here are some examples.

ls
ls -AF .
echo Hello world

The command is found using the current search path in the variable PATH[], which should define an array of directories containing executables.

A command may also be prefixed by environment variable definitions.

# Prints "Hello world"
env X="Hello world" Y=2 printenv X
# Pass the include path to the Visual C++
env include="c:\Program Files\Microsoft SDK\include" cl foo.cpp

10.2 Globbing

Commands may contain wildcard patterns. A pattern specifies a set of files through a limited kind of regular expression. Patterns are expanded before the function is executed.

   # List all files with a .c suffix
   ls *.c
   # List all files with a single character prefix, and .c suffix
   ls ?.c
   # Rename the file hello.ml to foo.ml
   mv {hello,foo}.ml

A comprehensive description of OMake glob patterns is given in Section 9.4.

10.3 Background jobs

The command may also be placed in the background by placing an ampersand after the command. Control returns to the shell without waiting for the job to complete. The job continues to run in the background.

    gcc -o hugeprogram *.c &

10.4 File redirection

Input and output can be redirected to files by using the <, >, and >& directives after the command.

    # Write to the "foo" file
    echo Hello world > foo
    # Redirect input from the foo file
    cat < foo
    # Redirect standard output and errors to the foo file
    gcc -o boo *.c >& foo

10.5 Pipelines

Pipelines are sequences of commands, where the output from each command is sent to the next. Pipelines are defined with the | and |& syntax. With | the output is redirected, but errors are not. With |& both output and errors are redirected.

   # Send the output of the ls command to the printer
   ls *.c | lpr
   # Send output and errors to jyh as email
   gcc -o hugefile *.c |& mail jyh

10.6 Conditional execution

Commands may also be composed though conditional evaluation using the || and && syntax. Every command has an integer exit code, which may be zero or some other integer. A command is said to succeed if its exit code is zero.

The expression command1 && command2 executes command2 only if command1

succeeds. The expression command1 || command2 executes command2 only if

command1 fails.
   # Display the x/y file if possible
   cd x && cat y
   # Run foo.exe, or print an error message
   (test -x foo.exe && foo.exe) || echo "foo.exe is not executable"

10.7 Grouping

Parenthesis are used for grouping in a pipeline or conditional command. In the following expression, the test function is used to test whether the foo.exe file is executable. If it is, the foo.exe file is executed. If the file is not executable (or

if the foo.exe command fails), the message "foo.exe is not executable" is

printed.

   # Run foo.exe, or print an error message
   (test -x foo.exe && foo.exe) || echo "foo.exe is not executable"

10.8 What is a shell command?

Syntactially, shell commands are any line that is not one of the following:

• A variable definition of the form VAR=string

• A function call f(...) or method call o.f(...)

• A rule definition containing a colon string: ...

• A special command, including the following:

       – if ...
       – switch ...

       – match ...
       – section ...
       – return ...

Commands may also be builtin (aliases). See the documentation for the Shell object for more information.

10.9 Basic builtin functions

10.9.1 echo

The echo function prints a string.

$(echo <args>)
echo <args>

10.9.2 cd

The cd function changes the current directory.

    cd(dir)
       dir : Dir

The cd function also supports a 2-argument form:

    $(cd dir, e)
       dir : Dir
       e : expression

In the two-argument form, expression e is evaluated in the directory dir. The current directory is not changed otherwise.

The behavior of the cd function can be changed with the CDPATH variable, which specifies a search path for directories. This is normally useful only in the osh command interpreter.

    CDPATH : Dir Sequence

For example, the following will change directory to the first directory ./foo,

~/dir1/foo, ~/dir2/foo.
    CDPATH[] =
       .
       $(HOME)/dir1
       $(HOME)/dir2
    cd foo

10.10. JOB CONTROL BUILTIN FUNCTIONS 167

10.10 Job control builtin functions

10.10.1 jobs

The jobs function prints a list of jobs.

   jobs

10.10.2 bg

The bg function places a job in the background.

   bg <pid...>

10.10.3 fg

The fg function brings a job to the foreground.

   fg <pid...>

10.10.4 stop

The stop function suspends a job.

   stop <pid...>

10.10.5 wait

The wait function waits for a job to finish. If no process identifiers are given, the shell waits for all jobs to complete.

   wait <pid...>

10.10.6 kill

The kill function signals a job.

   kill [signal] <pid...>

10.11 Command history

10.11.1 history

    $(history-index) : Int
    $(history) : String Sequence
    history-file : File
    history-length : Int

The history variables manage the command-line history in osh. They have no effect in omake.

The history-index variable is the current index into the command-line history. The history variable is the current command-line history.

The history-file variable can be redefined if you want the command-line history to be saved. The default value is ~/.omake/osh_history.

The history-length variable can be redefined to specify the maximum number of lines in the history that you want saved. The default value is 100.

The standard objects

Pervasives defines the objects that are defined in all programs. The following objects are defined.

11.1 Pervasives objects

11.1.1 Object

Parent objects: none.

The Object object is the root object. Every class is a subclass of Object. It provides the following fields:

• $(o.object-length): the number of fields and methods in the object.

• $(o.object-mem <var>): returns true iff the <var> is a field or method

of the object.

• $(o.object-add <var>, <value>): adds the field to the object, returning a new object.

• $(o.object-find <var>): fetches the field or method from the object; it is equivalent to $(o.<var>), but the variable can be non-constant.

• $(o.object-map <fun>): maps a function over the object. The function should take two arguments; the first is a field name, the second is the value of that field. The result is a new object constructed from the values returned by the function.

• o.object-foreach: the object-foreach form is equivalent to object-map,

but with altered syntax.

     o.object-foreach(<var1>, <var2>)
        <body>

For example, the following function prints all the fields of an object o.

   PrintObject(o) =
      o.object-foreach(v, x)
         println($(v) = $(x))

The export form is valid in a object-foreach body. The following function collects just the field names of an object.

   FieldNames(o) =
      names[] =
      o.object-foreach(v, x)
         names[] += $(v)
         export
      return $(names)

11.1.2 Map

Parent objects: Object.

A Map object is a dictionary from values to values. The <key> values are restricted to simple values: integers, floating-point numbers, strings, files, directories, and arrays of simple values.

The Map object provides the following methods.

• $(o.length): the number of items in the map.

• $(o.mem <key>): returns true iff the <key> is defined in the map.

• $(o.add <key>, <value>): adds the field to the map, returning a new map.

• $(o.find <key>): fetches the field from the map.

• $(o.keys): fetches an array of all the keys in the map, in alphabetical order.

• $(o.values): fetches an array of all the values in the map, in the alphabetical order of the corresponding keys.

• $(o.map <fun>): maps a function over the map. The function should take two arguments; the first is a field name, the second is the value of that field. The result is a new object constructed from the values returned by the function.

• o.foreach: the foreach form is equivalent to map, but with altered syntax.

        o.foreach(<var1>, <var2>)
           <body>

For example, the following function prints all the fields of an object o.

        PrintObject(o) =
           o.foreach(v, x)
              println($(v) = $(x))

The export form is valid in a foreach body. The following function collects just the field names of the map.

        FieldNames(o) =
           names =
           o.foreach(v, x)
              names += $(v)
              export
           return $(names)

There is also simpler syntax when the key is a string. The table can be defined using definitions with the form $|key| (the number of pipe symbols | is allowed to vary).

    $|key 1| = value1
    $||key1|key2|| = value2    # The key is key1|key2
    X = $|key 1|               # Define X to be the value of field $|key 1|

The usual modifiers are also allowed. The expression $`|key| represents lazy evaluation of the key, and $,|key| is normal evaluation.

11.1.3 Number

Parent objects: Object.

The Number object is the parent object for integers and floating-point numbers.

11.1.4 Int

Parent objects: Number.

The Int object represents integer values.

11.1.5 Float

Parent objects: Number.

The Float object represents floating-point numbers.

11.1.6 Sequence

Parent objects: Object.

The Sequence object represents a generic object containing sequential elements. It provides the following methods.

• $(s.length): the number of elements in the sequence.

• $(s.map <fun>): maps a function over the fields in the sequence. The function should take one argument. The result is a new sequence constructed from the values returned by the function.

• s.foreach: the foreach form is equivalent to map, but with altered syntax.

     s.foreach(<var>)
        <body>

For example, the following function prints all the elements of the sequence.

     PrintSequence(s) =
        s.foreach(x)
           println(Elem = $(x))

The export form is valid in a foreach body. The following function counts the number of zeros in the sequence.

     Zeros(s) =
        count = $(int 0)
        s.foreach(v)
           if $(equal $(v), 0)
              count = $(add $(count), 1)
              export
           export
        return $(count)

11.1.7 Array

Parent objects: Sequence.

The Array is a random-access sequence. It provides the following additional methods.

• $(s.nth <i>): returns element i of the sequence.

• $(s.rev <i>): returns the reversed sequence.

11.1. PERVASIVES OBJECTS 173

11.1.8 String

Parent objects: Array.

11.1.9 Fun

Parent objects: Object.

The Fun object provides the following methods.

• $(f.arity): the arity if the function.

11.1.10 Rule

Parent objects: Object.

The Rule object represents a build rule. It does not currently have any methods.

11.1.11 Target

Parent object: Object.

The Target object contains information collected for a specific target file.

• target: the target file.

• effects: the files that may be modified by a side-effect when this target is built.

• scanner_deps: static dependencies that must be built before this target can be scanned.

• static-deps: statically-defined build dependencies of this target.

• build-deps: all the build dependencies for the target, including static and scanned dependencies.

• build-values: all the value dependencies associated with the build.

• build-commands: the commands to build the target.

• output-file: if output was diverted to a file, with one of the --output-* options A, this field names that file. Otherwise it is false.

The object supports the following methods.

• find(file): returns a Target object for the given file. Raises a RuntimeException if the specified target is not part of the project.

• find-optional(file): returns a Target object for the given file, or false if the file is not part of the project.

NOTE: the information for a target is constructed dynamically, so it is possible that the Target object for a node will contain different values in different contexts. The easiest way to make sure that the Target information is complete is to compute it within a rule body, where the rule depends on the target file, or the dependencies of the target file.

11.1.12 Node

Parent objects: Object.

The Node object is the parent object for files and directories. It supports the following operations.

• $(node.stat): returns a stat object for the file. If the file is a symbolic link, the stat information is for the destination of the link, not the link itself.

• $(node.lstat): returns a stat object for the file or symbolic link.

• $(node.unlink): removes the file.

• $(node.rename <file>): renames the file.

• $(node.link <file>): creates a hard link <dst> to this file.

• $(node.symlink <file>): create a symbolic link <dst> to this file.

• $(node.chmod <perm>): change the permission of this file.

• $(node.chown <uid>, <gid>): change the owner and group id of this file.

11.1.13 File

Parent objects: Node.

The file object represents the name of a file.

11.1.14 Dir

Parent objects: Node.

The Dir object represents the name of a directory.

11.1.15 Channel

Parent objects: Object.

A Channel is a generic IO channel. It provides the following methods.

• $(o.close): close the channel.

11.1.16 InChannel

Parent objects: Channel.

A InChannel is an input channel. The variable stdin is the standard input channel.

It provides the following methods.

• $(InChannel.fopen <file>): open a new input channel.
• $(InChannel.of-string <string>): open a new input channel, using a

string as input.

11.1.17 OutChannel

Parent object: Channel.

A OutChannel is an output channel. The variables stdout and stderr are the standard output and error channels.

It provides the following methods.

• $(OutChannel.fopen <file>): open a new output channel.

• $(OutChannel.string): open a new output channel, writing to a string.

• $(OutChannel.to-string): get the current string of output, for an out-

  put channel created as OutChannel.open-string.

• $(OutChannel.append <file>): opens a new output channel, appending to the file.

• $(c.flush): flush the output channel.

• $(c.print <string>): print a string to the channel.

• $(c.println <string>): print a string to the channel, followed by a line terminator.

11.1.18 Location

Parent objects: Location.

The Location object represents a location in a file.

11.1.19 Position

Parent objects: Position.

The Position object represents a stack trace.

11.1.20 Exception

Parent objects: Object.

The Exception object is used as the base object for exceptions. It has no fields.

11.1.21 RuntimeException

Parent objects: Exception.

The RuntimeException object represents an exception from the runtime system. It has the following fields.

• position: a string representing the location where the exception was raised.

• message: a string containing the exception message.

11.1.22 UnbuildableException

Parent objects: Exception.

The UnbuildableException object should be used to signal that a target is not buildable. It will be caught by functions such as target-exists. This exception has the following fields:

• target: indicates which target is not buildable.

• message: a string containing the exception message.

11.1.23 Shell

Parent objects: Object.

The Shell object contains the collection of builtin functions available as shell commands.

You can define aliases by extending this object with additional methods. All methods in this class are called with one argument: a single array containing an argument list.

• echo

The echo function prints its arguments to the standard output channel.

• jobs

The jobs method prints the status of currently running commands.

• cd

The cd function changes the current directory. Note that the current directory follows the usual scoping rules. For example, the following program lists the files in the foo directory, but the current directory is not changed.

     section
        echo Listing files in the foo directory...
        cd foo
        ls

   echo Listing files in the current directory...
   ls
• bg

The bg method places a job in the background. The job is resumed if it has been suspended.

• fg

The fg method brings a job to the foreground. The job is resumed if it has been suspended.

• stop

The stop method suspends a running job.

• wait

The wait function waits for a running job to terminate. It is not possible to wait for a suspended job. The job is not brought to the foreground. If the wait is interrupted, the job continues to run in the background.

• kill

The kill function signal a job.

kill [signal] <pid...>.

The signals are either numeric, or symbolic. The symbolic signals are named as follows. ABRT, ALRM, HUP, ILL, KILL, QUIT, SEGV, TERM, USR1, USR2, CHLD, STOP, TSTP, TTIN, TTOU, VTALRM, PROF.

• exit

The exit function terminates the current session.

• which, where

See the documentation for the corresponding functions.

• rehash

Reset the search path.

• ln-or-cp src dst

Links or copies src to dst, overwriting dst. Namely, ln-or-cp would first delete the dst file (unless it is a directory), if it exists. Next it would try to create a symbolic link dst poiting to src (it will make all the necessary adjustmnents of relative paths). If symbolic link can not be created (e.g. the OS or the filesystem does not support symbolic links), it will try to create a hard link. If that fails too, it will try to forcibly copy src to dst.

• history

Print the current command-line history.

• digest

Print the digests of the given files.

• Win32 functions.

Win32 doesn’t provide very many programs for scripting, except for the functions that are builtin to the DOS cmd.exe. The following functions are defined on Win32 and only on Win32. On other systems, it is expected that these programs already exist.

  – grep
       grep [-q] [-n] pattern files...

The grep function calls the omake grep function.

• Internal versions of standard system commands.

By default, omake uses internal versions of the following commands: cp, mv, cat, rm, mkdir, chmod, test, find. If you really want to use the standard system versions of these commands, set the USE_SYSTEM_COMMANDS as one of the first definitions in your OMakeroot file.

  – mkdir
        mkdir [-m <mode>] [-p] files

The mkdir function is used to create directories. The -verb+-m+ option can be used to specify the permission mode of the created directory. If the -p option is specified, the full path is created.

  – cp
  – mv
        cp [-f] [-i] [-v] src dst
        cp [-f] [-i] [-v] files dst
        mv [-f] [-i] [-v] src dst
        mv [-f] [-i] [-v] files dst

The cp function copies a src file to a dst file, overwriting it if it already exists. If more than one source file is specified, the final file must be a directory, and the source files are copied into the directory.

-f Copy files forcibly, do not prompt.

-i Prompt before removing destination files.

-v Explain what is happening.

  – rm

   rm [-f] [-i] [-v] [-r] files
   rmdir [-f] [-i] [-v] [-r] dirs

The rm function removes a set of files. No warnings are issued if the files do not exist, or if they cannot be removed. Options:

-f Forcibly remove files, do not prompt. -i Prompt before removal. -v Explain what is happening. -r Remove contents of directories recursively.

– chmod
    chmod [-r] [-v] [-f] mode files

The chmod function changes the permissions on a set of files or directories. This function does nothing on Win32. The mode may be specified as an octal number, or in symbolic form [ugoa]*[- =][rwxXstugo]+. See the man page for chmod for details. Options:

-r Change permissions of all files in a directory recursively. -v Explain what is happening. -f Continue on errors.

– cat
   cat files...

The cat function prints the contents of the files to stdout

– test

test expression [ expression +]+

[ --help
[ --version

See the documentation for the test function.

– find
   find \emph{expression}

See the documentation for the find function.

180 CHAPTER 11. THE STANDARD OBJECTS

Build functions and utilities

12.1 Builtin .PHONY targets

The complete set of builtin .PHONY targets include the following. .PHONY Declares new phony targets (Section 7.10). .DEFAULT Declare the default build targets (Section 7.7). .SUBDIRS Include a directory as part of the project (Section 7.8). .SCANNER Define a dependency scanner (Section 7.8). .INCLUDE Include a file (Section 7.9). .ORDER Define a file-dependency ordering rule (Section 9.3.5). .BUILD BEGIN Commands to be executed at the beginning of a build. .BUILD SUCCESS Commands to be executed if the build is successful. .BUILD FAILURE Commands to be executed if the build fails.

The .BUILD targets can be used to specify commands to be executed at the beginning and end of the build. The .BUILD_BEGIN target is built at the beginning of a project build, and one of .BUILD_FAILURE or .BUILD_SUCCESS is executed when the build terminates.

For example, the following set of rules simply print additional messages about the status of the build.

   .BUILD_BEGIN:
       echo Build starting
   .BUILD_SUCCESS:
       echo The build was successful

   .BUILD_FAILURE:
       println($"The build failed: $(length $(find-build-targets Failed)) targets could not be built")

Another common use is to define notifications to be performed when the build completes. For example, the following rule will create a new X terminal displaying the summary of the build (using the BUILD SUMMARY variable).

    .BUILD_FAILURE:
        xterm -e vi $(BUILD_SUMMARY)

If you do not wish to add these rules directly to your project (which is probably a good idea if you work with others), you can define them in your .omakerc (see Section A.8).

The find-build-targets function is useful for obtaining a firther summary of the build. Note that when output diversions are in effect (with the --output-* options — see Chapter A), any output produced by the commands is copied to a file. The name of the file is specified by the output-file field of the Target object. You may find this useful in defining custom build summaries.

12.2 Options and versioning

12.2.1 OMakeFlags

   OMakeFlags(options)
      options : String

The OMakeFlags function is used to set omake options from within OMakefiles. The options have exactly the same format as options on the command line.

For example, the following code displays the progress bar unless the VERBOSE environment variable is defined.

    if $(not $(defined-env VERBOSE))
        OMakeFlags(-S --progress)
        export

12.2.2 OMakeVersion

   OMakeVersion(version1)
   OMakeVersion(version1, version2)
      version1, version2 : String

The OMakeVersion function is used for version checking in OMakefiles. It takes one or two arguments.

In the one argument form, if the omake version number is less than <version1>, then an exception is raised. In the two argument form, the version must lie be-

tween version1 and version2.

12.2.3 cmp-versions

   $(cmp-versions version1, version2)
      version1, version2 : String

The cmp-versions\ functions can be used to compare arbitrary version strings. It returns 0 when the two version strings are equal, a negative number when the first string represents an earlier version, and a positive number otherwise.

12.2.4 DefineCommandVars

   DefineCommandVars()

The DefineCommandVars function redefines the variables passed on the commandline. Variables definitions are passed on the command line in the form name=value. This function is primarily for internal use by omake to define these variables for the first time.

12.3 Examining the dependency graph

12.3.1 dependencies, dependencies-all, dependencies-proper

   $(dependencies targets) : File Array
   $(dependencies-all targets) : File Array
   $(dependencies-proper targets) : File Array
      targets : File Array
   raises RuntimeException

The dependencies function returns the set of immediate dependencies of the given targets. This function can only be used within a rule body and all the arguments to the dependency function must also be dependencies of this rule. This restriction ensures that all the dependencies are known when this function is executed.

The dependencies-all function is similar, but it expands the dependencies recursively, returning all of the dependencies of a target, not just the immediate ones.

The dependencies-proper function returns all recursive dependencies, except the dependencies that are leaf targets. A leaf target is a target that has no dependencies and no build commands; a leaf target corresponds to a source file in the current project.

In all three functions, files that are not part of the current project are silently discarded.

One purpose of the dependencies-proper function is for “clean” targets. For example, one way to delete all intermediate files in a build is with a rule that uses the dependencies-proper. Note however, that the rule requires building the project before it can be deleted. For a shorter form, see the

filter-proper-targets function.
    .PHONY: clean
    APP = ...     # the name of the target application
    clean: $(APP)
       rm $(dependencies-proper $(APP))

12.3.2 target

   $(target targets) : Target Array
      targets : File Sequence
   raises RuntimeException

The target function returns the Target object associated with each of the targets. See the Target object for more information.

12.3.3 find-build-targets

    $(find-build-targets tag) : Target Array
       tag : Succeeded | Failed

The find-build-targets allow the results of the build to be examined. The tag must specifies which targets are to be returned; the comparison is case-insensitive.

Succeeded The list of targets that were built successfully.

Failed The list of targets that could not be built.

These are used mainly in conjuction with the .BUILD_SUCCESS (Section 12.1) and .BUILD_FAILURE (Section 12.1) phony targets. For example, adding the following to your project OMakefile will print the number of targets that failed (if the build failed).

    .BUILD_FAILURE:
        echo "Failed target count: $(length $(find-build-targets Failed))"

12.3.4 project-directories

   $(project-directories) : Dir Array

The project-directories function returns the list of all directories that are considered to be part of the project.

To get the complete directory list, this function should be called from within a rule body.

12.3.5 rule

The rule function is called whenever a build rule is defined. It is unlikely that you will need to redefine this function, except in very exceptional cases.

   rule(multiple, target, pattern, sources, options, body) : Rule
      multiple : String
      target   : Sequence
      pattern  : Sequence
      sources  : Sequence
      options  : Array
      body     : Body

The rule function is called when a rule is evaluated.

multiple A Boolean value indicating whether the rule was defined with a double colon ::.

target The sequence of target names.

pattern The sequence of patterns. This sequence will be empty for two-part rules.

sources The sequence of dependencies.

options An array of options. Each option is represented as a Map object associating each specified option with a value.

body The body expression of the rule.

Consider the following rule.

   target: pattern: sources :name1: option1 :name2: option2
      expr1
      expr2

This expression represents the following function call, where square brackets are used to indicate arrays, and the curly brackets represent a Map object.

   rule(false, target, pattern, sources,
        { $|:name1:| = option1; $|:name2:| = option2 }
        [expr1; expr2])

12.4 The OMakeroot file

The standard OMakeroot file defines the functions are rules for building standard projects.

12.4.1 Variables

ROOT The root directory of the current project.

CWD The current working directory (the directory is set for each OMakefile in the project).

EMPTY The empty string.

STDROOT The name of the standard installed OMakeroot file.

ABORT ON COMMAND ERROR If set to true, the construction of a target should be aborted whenever one of the commands to build it fail. This defaults to true, and should normally be left that way.

SCANNER MODE This variable should be defined as one of four values (defaults to enabled).

enabled Allow the use of default .SCANNER rules. Whenever a rule does not specify a :scanner: dependency explicitly, try to find a .SCANNER with the same target name.

disabled Never use default .SCANNER rules.

warning Allow the use of default .SCANNER rules, but print a warning whenever one is selected.

error Do not allow the use of default .SCANNER rules. If a rule does not specify a :scanner: dependency, and there is a default .SCANNER rule, the build will terminate abnormally.

12.4.2 System variables

INSTALL The command to install a program (install on Unix, cp on

Win32).

PATHSEP The normal path separator (: on Unix, ; on Win32).

DIRSEP The normal directory separator (/ on Unix, \ on Win32).

EXT LIB File suffix for a static library (default is .a on Unix, and .lib on

Win32).

EXT OBJ File suffix for an object file (default is .o on Unix, and .obj on

Win32).

EXT ASM File suffix for an assembly file (default is .s on Unix, and .asm

on Win32).

EXE File suffix for executables (default is empty for Unix, and .exe on Win32

and Cygwin).

12.5 Building C and C++ code

OMake provides extensive support for building C and C++ programs. In order to use the functions defined in this section, you need to make sure the line

open build/C

is present in your OMakeroot file.

12.5.1 Autoconfiguration variables

These variables will get defined based on the “autoconf-style” static. tests executed when you run OMake for the first time. You can use them to configure your project accordingly, and you should not redefine them.

You can use the --configure command line option (Section A.3.9) to force re-execution of all the tests.

A different set of autoconfiguration tests is performed depending on the build environment involved — one set of tests would be performed in a Win32 environment, and another — in a Unix-like environment (including Linux, OS X and Cygwin).

12.5.1.1 Unix-like systems

GCC FOUND A boolean flag specifying whether the gcc binary was found in your path.

GXX FOUND A boolean flag specifying whether the g++ binary was found in your path.

12.5.1.2 Win32

CL FOUND A boolean flag specifying whether the cl binary was found in your path.

LIB FOUND A boolean flag specifying whether the lib binary was found in your path.

12.5.2 C and C++ configuration variables

The following variables can be redefined in your project.

188 CHAPTER 12. BUILD FUNCTIONS AND UTILITIES CC The name of the C compiler (on Unix it defaults to gcc when gcc is present and to cc otherwise; on Win32 defaults to cl /nologo). CXX The name of the C++ compiler (on Unix it defaults to gcc when gcc is present and to c++ otherwise; on Win32 defaults to cl /nologo). CPP The name of the C preprocessor (defaults to cpp on Unix, and cl /E

on Win32).

CFLAGS Compilation flags to pass to the C compiler (default empty on

Unix, and /DWIN32 on Win32).

CXXFLAGS Compilation flags to pass to the C++ compiler (default empty

on Unix, and /DWIN32 on Win32).

INCLUDES Additional directories that specify the search path to the C and C++ compilers (default is .). The directories are passed to the C and C++ compilers with the -I option. The include path with -I prefixes is defined in

the PREFIXED_INCLUDES variable.

LIBS Additional libraries needed when building a program (default is empty). CCOUT The option to use for specifying the output file in C and C++ compilers (defaults to -o on Unix and /Fo on Win32). AS The name of the assembler (defaults to as on Unix, and ml on Win32). ASFLAGS Flags to pass to the assembler (default is empty on Unix, and

/c /coff on Win32).

ASOUT The option string that specifies the output file for AS (defaults to -o

on Unix and /Fo on Win32).

AR The name of the program to create static libraries (defaults to ar cq on

Unix, and lib on Win32).

LD The name of the linker (defaults to ld on Unix, and cl on Win32). LDFLAGS Options to pass to the linker (default is empty). LDOUT The option to use for specifying the output file in C and C++ linkers (defaults to -o on Unix and /Fe on Win32).

YACC The name of the yacc parser generator (default is yacc on Unix, empty

on Win32).

LEX The name of the lex lexer generator (default is lex on Unix, empty on

Win32).

12.5.3 Generated C files

Because the C scanners do not normally know anything about generated source files (such as generated header files), these files may need to be created before running the scanner.

12.5.3.1 CGeneratedFiles, LocalCGeneratedFiles

CGeneratedFiles(files)
LocalCGeneratedFiles(files)
   The CGeneratedFiles and LocalCGeneratedFiles functions specify files

that need to be generated before any C files are scanned for dependencies. For example, if config.h and inputs.h are both generated files, specify:

CGeneratedFiles(config.h inputs.h)

The CGeneratedFiles function is global — its arguments will be generated before any C files anywhere in the project are scanned for dependencies. The LocalCGeneratedFiles function follows the normal scoping rules of OMake.

12.5.4 Building C programs and Libraries

12.5.4.1 StaticCLibrary

The StaticCLibrary builds a static library.

   StaticCLibrary(<target>, <files>)

The <target> does not include the library suffix, and The <files> list does not include the object suffix. These are obtained from the EXT_LIB and EXT_OBJ variables.

This function returns the library filename.

The following command builds the library libfoo.a from the files a.o b.o c.o

on Unix, or the library libfoo.lib from the files a.obj b.obj c.obj on Win32.
StaticCLibrary(libfoo, a b c)
.DEFAULT: $(StaticCLibrary libbar, a b c d)

12.5.4.2 StaticCLibraryCopy

The StaticCLibraryCopy function copies the static library to an install location.

   StaticCLibraryCopy(<tag>, <dir>, <lib>)

The <tag> is the name of a target (typically a .PHONY target); the <dir> is the installation directory, and <lib> is the library to be copied (without the library suffix).

This function returns the filename of the library in the target directory.

For example, the following code copies the library libfoo.a to the /usr/lib directory.

.PHONY: install
StaticCLibraryCopy(install, /usr/lib, libfoo)

12.5.4.3 StaticCLibraryInstall

The StaticCLibraryInstall function builds a library, and sets the install location in one step. It returns the filename of the library in the target directory.

   StaticCLibraryInstall(<tag>, <dir>, <libname>, <files>)
StaticCLibraryInstall(install, /usr/lib, libfoo, a b c)

12.5.4.4 StaticCObject, StaticCObjectCopy, StaticCObjectInstall

These functions mirror the StaticCLibrary, StaticCLibraryCopy, and StaticCLibraryInstall

functions, but they build an object file (a .o file on Unix, and a .obj file on

Win32).

12.5.4.5 CProgram

The CProgram function builds a C program from a set of object files and libraries.

   CProgram(<name>, <files>)

The <name> argument specifies the name of the program to be built; the <files> argument specifies the files to be linked. The function returns the filename of the executable.

Additional options can be passed through the following variables.

CFLAGS Flags used by the C compiler during the link step.

LDFLAGS Flags to pass to the loader.

LIBS Additional libraries to be linked.

For example, the following code specifies that the program foo is to be produced by linking the files bar.o and baz.o and libraries libfoo.a.

section
   LIBS = libfoo
   LDFLAGS += -lbar
   CProgram(foo, bar baz)

12.5.4.6 CProgramCopy

The CProgramCopy function copies a file to an install location.

   CProgramCopy(<tag>, <dir>, <program>)
CProgramCopy(install, /usr/bin, foo)

12.5.4.7 CProgramInstall

The CProgramInstall function specifies a program to build, and a location to install, simultaneously.

   CProgramInstall(<tag>, <dir>, <name>, <files>)
section
   LIBS = libfoo
   LDFLAGS += -lbar
   CProgramInstall(install, /usr/bin, foo, bar baz)

12.5.4.8 CXXProgram, CXXProgramInstall

The CXXProgram and CXXProgramInstall functions are equivalent to their C counterparts, except that would use $(CXX) and $(CXXFLAGS) for linking instead

of $(CC) and $(CFLAGS).

12.5.4.9 StaticCXXLibrary, StaticCXXLibraryCopy, StaticCXXLibraryInstall

Similarly, StaticCXXLibrary, StaticCXXLibraryCopy and StaticCXXLibraryInstall
are the C++ equivalents of StaticCLibrary, StaticCLibraryCopy and StaticCLibraryInstall

functions.

12.6 Building OCaml code

OMake provides extensive support for building OCaml code, including support for tools like ocamlfind, ocamlyacc and menhir. In order to use the functions defined in this section, you need to make sure the line

open build/OCaml

is present in your OMakeroot file.

192 CHAPTER 12. BUILD FUNCTIONS AND UTILITIES

12.6.1 Autoconfiguration variables for OCaml compilation

These variables will get defined based on the “autoconf-style” tests executed when you run OMake for the first time. You can use them to configure your project accordingly, and you should not redefine them.

You can use the --configure command line option (Section A.3.9) to force re-execution of all the tests. OCAMLOPT EXISTS True when ocamlopt (or ocamlopt.opt) is available on your machine. OCAMLFIND EXISTS True when the ocamlfind is available on your machines. OCAMLDEP MODULES AVAILABLE True when a version of ocamldep that understands the -modules option is available on your machine. MENHIR AVAILABLE True when the Menhir parser-generator is available on your machine.

12.6.2 Configuration variables for OCaml compilation

The following variables can be redefined in your project. USE OCAMLFIND Whether to use the ocamlfind utility (default false) OCAMLC The OCaml bytecode compiler (default ocamlc.opt if it exists and USE_OCAMLFIND is not set, otherwise ocamlc). OCAMLOPT The OCaml native-code compiler (default ocamlopt.opt if it exists and USE_OCAMLFIND is not set, otherwise ocamlopt). CAMLP4 The camlp4 preprocessor (default camlp4). OCAMLLEX The OCaml lexer generator (default ocamllex). OCAMLLEXFLAGS The flags to pass to ocamllex (default -q). OCAMLYACC The OCaml parser generator (default ocamlyacc). OCAMLYACCFLAGS Additional options to pass to $(OCAMLYACC). OCAMLDEP The OCaml dependency analyzer (default ocamldep).

OCAMLDEP MODULES The OCaml dependency analyzer that understands the -module option (default ocamldep, if ocamldep -modules works,

or ocamlrun ocamldep-omake, if ocamlrun ocamldep-omake -modules works,

and empty when neither works).

OCAMLDEP MODULES ENABLED Instead of using OCAMLDEP in a traditional make-style fashion, run $(OCAMLDEP_MODULES) -modules and then postprocess the output internally to discover all the relevant generated .ml and .mli files. See Section 12.6.5 for more information on interactions between OMake, OCAMLDEP and generated files. This feature is currently considered highly experimental and is disabled by default.

OCAMLMKTOP The OCaml toploop compiler (default ocamlmktop).

OCAMLLINK The OCaml bytecode linker (default $(OCAMLC)).

OCAMLOPTLINK The OCaml native-code linker (default $(OCAMLOPT)).

OCAMLINCLUDES Search path to pass to the OCaml compilers (default .). The search path with the -I prefix is defined by the PREFIXED_OCAMLINCLUDES variable.

OCAMLFIND The ocamlfind utility (default ocamlfind if USE_OCAMLFIND

is set, otherwise empty).

OCAMLFINDFLAGS The flags to pass to ocamlfind (default empty, USE_OCAMLFIND must be set).

OCAMLPACKS Package names to pass to ocamlfind (USE_OCAMLFIND must be set).

BYTE ENABLED Flag indicating whether to use the bytecode compiler (default true, when no ocamlopt found, false otherwise).

NATIVE ENABLED Flag indicating whether to use the native-code compiler (default true, when ocamlopt is found, false otherwise). Both BYTE_ENABLED and NATIVE_ENABLED can be set to true; at least one should be set to true.

MENHIR ENABLED Define this as true if you wish to use menhir instead

of ocamlyacc (default false).

12.6.3 OCaml command flags

The following variables specify additional options to be passed to the OCaml tools.

194 CHAPTER 12. BUILD FUNCTIONS AND UTILITIES OCAMLDEPFLAGS Flags to pass to OCAMLDEP (but not to OCAMLDEP_MODULES). OCAMLPPFLAGS Flags to pass to CAMLP4. OCAMLCFLAGS Flags to pass to the byte-code compiler (default -g). OCAMLOPTFLAGS Flags to pass to the native-code compiler (default empty). OCAMLFLAGS Flags to pass to either compiler (default -warn-error A). OCAML BYTE LINK FLAGS Flags to pass to the byte-code linker (default empty). OCAML NATIVE LINK FLAGS Flags to pass to the native-code linker (default empty). OCAML LINK FLAGS Flags to pass to either linker. MENHIR FLAGS Additional flags to pass to menhir.

12.6.4 Library variables

The following variables are used during linking. OCAML LIBS Libraries to pass to the linker. These libraries become dependencies of the link step. OCAML OTHER LIBS Additional libraries to pass to the linker. These libraries are not included as dependencies to the link step. Typical use is for the OCaml standard libraries like unix or str. OCAML CLIBS C libraries to pass to the linker. OCAML LIB FLAGS Extra flags for the library linker. ABORT ON DEPENDENCY ERRORS OCaml linker requires the OCaml files to be listed in dependency order. Normally, all the functions presented in this section will automatically sort the list of OCaml modules passed in as the <files> argument. However, this variable is set to true, the order of the files passed into these function will be left as is, but OMake will abort with an error message if the order is illegal.

12.6.5 Generated OCaml Files

As of OCaml version 3.09.2, the standard ocamldep scanner is “broken”. The main issue is that it finds only those dependencies that already exist. If foo.ml contains a dependency on Bar,

foo.ml:
   open Bar

then the default ocamldep will only find the dependency if a file bar.ml or bar.ml exists in the include path. It will not find (or print) the dependency if, for example, only bar.mly exists at the time ocamldep is run, even though bar.ml and bar.mli can be generated from bar.mly.

OMake currently provides two methods for addressing this problem — one that requires manually specifying the generated files, and an experimental method for discovering such “hidden” dependencies automatically. The OCAMLDEP MODULES ENABLED variable controls which method is going to be used. When this variable is false, the manual specifications are expected and when it is true, the automated discovery will be attempted.

12.6.5.1 OCamlGeneratedFiles, LocalOCamlGeneratedFiles

OCamlGeneratedFiles(files)
LocalOCamlGeneratedFiles(files)

When the OCAMLDEP MODULES ENABLED variable variable is set to false, the

OCamlGeneratedFiles and LocalOCamlGeneratedFiles functions specify files

that need to be generated before any OCaml files are scanned for dependencies. For example, if parser.ml and lexer.ml are both generated files, specify:

OCamlGeneratedFiles(parser.ml lexer.ml)

The OCamlGeneratedFiles function is global — its arguments will be generated before any OCaml files anywhere in the project are scanned for dependencies. The LocalOCamlGeneratedFiles function follows the normal scoping rules of OMake.

These functions have no effect when the OCAMLDEP MODULES ENABLED variable is true.

12.6.5.2 Automatic discovery of generated files during dependency analysis

Having to specify the generated files manualy when OMake could discover them automatically is obviously suboptimal. To address this, we try to use a custom ocamldep that only finds the free module names in a file.

This functionality is experimental and is disabled by default for now. Set the

OCAMLDEP MODULES ENABLED variable to true (or to $(OCAMLDEP_MODULES_AVAILABLE))

in your project to enable it.

Note that the experimental ocamldep functionality this relies upon is not yet included in the standard OCaml (it is expected to be a part of the upcoming

OCaml 3.10 — see http://caml.inria.fr/mantis/view.php?id=4047). Tem-

porarily, we distribute a bytecode version ocamldep-omake of the appropriately modified ocamldep. The appropriate ocamldep will be discovered automatically

— see and the OCAMLDEP MODULES AVAILABLE and OCAMLDEP MODULES variables

will be set accordingly.

12.6.6 Using the Menhir parser generator

Menhir is a parser generator that is mostly compatible with ocamlyacc, but with many improvements. A few of these are listed here (excerpted from the

Menhir home page http://cristal.inria.fr/~fpottier/menhir/).

• Menhir’s explanations are believed to be understandable by mere humans.

• Menhir allows grammar specifications to be split over multiple files. It also allows several grammars to share a single set of tokens.

• Menhir is able to produce parsers that are parameterized by Objective Caml modules.

Added by jyh With the --infer option, Menhir can typecheck the semantic actions in your grammar at generation time.

What do you need to do to use Menhir instead of ocamlyacc?

1. Place the following definition before the relevant section of your project (or at the top of your project OMakefile if you want to use Menhir everywhere).

        MENHIR_ENABLED = true

2. Optionally, add any desired Menhir options to the MENHIR_FLAGS variable.

        MENHIR_FLAGS += --infer

With this setup, any file with a .mly suffix will be compiled with Menhir. If your grammar is split across several files, you need to specify it explicitly,

using the MenhirMulti function.

    MenhirMulti(target, sources)
        target : filename, without suffix
        sources : the files that define the grammar, without suffixes

For example, if you want to generate the parser files parse.ml and parse.mli, from the grammar specified in files a.mly and b.mly, you would use the following.

    MenhirMulti(parse, a b)

12.6.6.1 OCamlLibrary

The OCamlLibrary function builds an OCaml library.

   OCamlLibrary(<libname>, <files>)

The <libname> and <files> are listed without suffixes.

This function returns the list of all the targets that it defines the rules for

(including the $(name)$(EXT_LIB) file when NATIVE_ENABLED is set).

The following code builds the libfoo.cmxa library from the files foo.cmx

and bar.cmx (if NATIVE_ENABLED is set), and libfoo.cma from foo.cmo and
bar.cmo (if BYTE_ENABLED is set).
OCamlLibrary(libfoo, foo bar)

12.6.6.2 OCamlPackage

The OCamlPackage function builds an OCaml package.

   OCamlPackage(<name>, <files>)

The <name> and <files> are listed without suffixes. The <files> must have been compiled with the -for-pack <ident> flag to the OCaml compiler.

This function returns the list of all the targets that it defines the rules for

(including the $(name)$(EXT_LIB) file when NATIVE_ENABLED is set).

The following code builds the libfoo.cmx package from the files package.cmx

and bar.cmx (if NATIVE_ENABLED is set), and package.cmo from foo.cmo and
bar.cmo (if BYTE_ENABLED is set).
OCamlPackage(package, foo bar)

12.6.6.3 OCamlLibraryCopy

The OCamlLibraryCopy function copies a library to an install location.

   OCamlLibraryCopy(<tag>, <libdir>, <libname>, <interface-files>)

The <interface-files> specify additional interface files to be copied if the

INSTALL_INTERFACES variable is true.

12.6.6.4 OCamlLibraryInstall

The OCamlLibraryInstall function builds a library and copies it to an install location in one step.

   OCamlLibraryInstall(<tag>, <libdir>, <libname>, <files>)

12.6.6.5 OCamlProgram

The OCamlProgram function builds an OCaml program. It returns the array with all the targets for which it has defined the rules ($(name)$(EXE) and

$(name).run and/or $(name).opt, depending on the NATIVE_ENABLED and
BYTE_ENABLED variables).
   OCamlProgram(<name>, <files>)

Additional variables used:

OCAML LIBS Additional libraries passed to the linker, without suffix. These files become dependencies of the target program.

OCAML OTHER LIBS Additional libraries passed to the linker, without suffix. These files do not become dependencies of the target program.

OCAML CLIBS C libraries to pass to the linker.

OCAML BYTE LINK FLAGS Flags to pass to the bytecode linker.

OCAML NATIVE LINK FLAGS Flags to pass to the native code linker.

OCAML LINK FLAGS Flags to pass to both linkers.

12.6.6.6 OCamlProgramCopy

The OCamlProgramCopy function copies an OCaml program to an install location.

   OCamlProgramCopy(<tag>, <bindir>, <name>)

Additional variables used:

NATIVE ENABLED If the NATIVE ENABLED variable is set, the native-code executable is copied; otherwise the byte-code executable is copied.

12.6.6.7 OCamlProgramInstall

The OCamlProgramInstall function builds a programs and copies it to an install location in one step.

   OCamlProgramInstall(<tag>, <bindir>, <name>, <files>)

12.7 Building L^{A}TEX files

OMake provides support for building L^{A}TEX documents, including support for automatically running BiBTex and for producing PostScript and PDF files. In order to use the functions defined in this section, you need to make sure the line

open build/LaTeX

is present in your OMakeroot file.

12.7.1 Configuration variables

The following variables can be modified in your project.

LATEX The L^{A}TEX command (default latex).

TETEX2 ENABLED Flag indicating whether to use advanced L^{A}TEX options present in TeTeX v.2 (default value is determined the first time omake reads LaTeX.src and depends on the version of L^{A}TEX you have installed).

12.7. BUILDING L^{A}TEX FILES 199 LATEXFLAGS The L^{A}TEX flags (defaults depend on the TETEX2_ENABLED variable) BIBTEX The BibTeX command (default bibtex). MAKEINDEX The command to build an index (default makeindex). DVIPS The .dvi to PostScript converter (default dvips). DVIPSFLAGS Flags to pass to dvips (default -t letter). DVIPDFM The .dvi to .pdf converter (default dvipdfm). DVIPDFMFLAGS Flags to pass to dvipdfm (default -p letter). PDFLATEX The .latex to .pdf converter (default pdflatex). PDFLATEXFLAGS Flags to pass to pdflatex (default is $`(LATEXFLAGS)). USEPDFLATEX Flag indicating whether to use pdflatex instead of dvipdfm to generate the .pdf document (default false).

12.7.2 Building L^{A}TEX documents

12.7.2.1 LaTeXDocument The LaTeXDocument produces a L^{A}TEX document.

   LaTeXDocument(<name>, <texfiles>)

The document <name> and <texfiles> are listed without suffixes. This function returns the filenames for the generated .ps and .pdf files.

Additional variables used: TEXINPUTS The L^{A}TEX search path (an array of directories, default is taken from the TEXINPUTS environment variable). TEXDEPS Additional files this document depends on. TEXVARS An array of names of the environment variables that are to be updated based on the value of OMake’s TEXINPUTS variable. Defaults to

TEXINPUTS BIBINPUTS BSTINPUTS.

12.7.2.2 TeXGeneratedFiles, LocalTeXGeneratedFiles

TeXGeneratedFiles(files)
LocalTeXGeneratedFiles(files)
   The TeXGeneratedFiles and LocalTeXGeneratedFiles functions specify

files that need to be generated before any L^{A}TEXfiles are scanned for dependencies. For example, if config.tex and inputs.tex are both generated files, specify:

    TeXGeneratedFiles(config.tex inputs.tex)

The TeXGeneratedFiles function is global — its arguments will be generated before any TeX files anywhere in the project are scanned for dependencies. The LocalTeXGeneratedFiles function follows the normal scoping rules of OMake.

12.7.2.3 LaTeXDocumentCopy

The LaTeXDocumentCopy copies the document to an install location.

   LaTeXDocumentCopy(<tag>, <libdir>, <installname>, <docname>)

This function copies just the .pdf and .ps files.

12.7.2.4 LaTeXDocumentInstall

The LaTeXDocumentInstall builds a document and copies it to an install location in one step.

   LaTeXDocumentInstall(<tag>, <libdir>, <installname>, <docname>, <files>)

Autoconfiguration functions and variables

OMake standard library provides a number of functions and variables intended to help one write build specifications that need to be capable of autoconfiguring itself to adjust to different build environments.

13.1 General-purpose autoconfiguration functions

The following general-purpose functions can be used to discover the properties of your build environment in a fashion similar to the one used by GNU autoconf tool you may be familiar with. It is recommended that these function be used from an appropriate static. block (see Section 4.18 for more information).

In order to use the following general-purpose functions, you need to have the line

open configure/Configure
included in your OMakefile or OMakeroot.

13.1.1 ConfMsgChecking, ConfMsgResult

ConfMsgChecking(<msg>)
...
ConfMsgResult(<msg>)

The ConfMsgChecking function output message of the form --- Checking <msg>... without any trailing newline. After the test advertized by ConfMsgChecking is performed, the ConfMsgResult function should be used to output the result.

In certain cases users may want to redefine these function — for example, to use a different output formatting and/or to copy the messages to a log file.

Example:

202CHAPTER 13. AUTOCONFIGURATION FUNCTIONS AND VARIABLES

static. =
   ConfMsgChecking(which foo to use)
   foo = ...
   ConfMsgResult($(foo))

13.1.2 ConfMsgWarn, ConfMsgError

ConfMsgWarn(<msg>)
ConfMsgError(<msg>)

Print a warning or an error message respectively. ConfMsgError would then abort OMake.

13.1.3 ConfMsgYesNo, ConfMsgFound

flag = $(ConfMsgYesNo <bool expr>
flag = $(ConfMsgFound <bool expr>

The ConfMsgFound function expects to receive a boolean flag describing whether a test previously announced using the ConfMsgChecking function found what it was looking for. ConfMsgFound will output the appropriate result (“found” or “NOT found”) using the ConfMsgResult function and return its argument back.

The ConfMsgYesNo function is similar, outputting a simple (“yes” or “NO”).

13.1.4 TryCompileC, TryLinkC, TryRunC

success = $(TryCompileC <prog_text>)
success = $(TryLinkC <prog_text>)
success = $(TryRunC <prog_text>)

Given the text of a C program, the TryCompileC, TryLinkC, and TryRunC functions would try to compile / compile and link / compile, link, and run, the given program and return a boolean flag indicating whether the attempt was successful.

TryCompileC will use the CC, CFLAGS and INCLUDES variables to run the C compiler. TryLinkC and TryRunC will also use the LDFLAGS variable to run the C compiler and linker. However, the flags like /WX, -Werror and -warn-error will be not be passed to the compiler, even if they occur in CFLAGS.

These functions are silent and should normally be used with an appropriate

ConfMsgChecking . . . ConfMsgResult.

13.1.5 RunCProg

output = $(RunCProg <prog>)

RunCProg is similar to the RunCProg function, except that it returns the output of the function (will return false if the program fails to compile or run).

13.1.6 CheckCHeader, VerboseCheckCHeader

success = $(CheckCHeader <files>)
success = $(VerboseCheckCHeader <files>)

Use the TryCompileC function to check whether your C compiler can locate and process the specified headers files. Will incude <stdio.h> before including the header files.

Both functions return a boolean value. The CheckCHeader function is silent;

the VerboseCheckCHeader function will use the ConfMsgChecking and ConfMsgResult

functions to describe the test and the outcome.

Example:

static. =
   NCURSES_H_AVAILABLE = $(VerboseCheckCHeader ncurses.h)

13.1.7 CheckCLib, VerboseCheckCLib

success = $(CheckCLib <libs>, <functions>)
success = $(VerboseCheckCLib <libs>, <functions>)

Use the TryLinkC function to check whether your C compiler and linker can find the named functions when linking with the named libraries. Will pass the <libs> to the compiler using the -l flag.

Both functions return a boolean value. The CheckCLib function is silent; the

VerboseCheckCHeader function will use the ConfMsgChecking and ConfMsgResult

functions to describe the test and the outcome.

Example:

static. =
    NCURSES_LIB_AVAILABLE = $(VerboseCheckCLib ncurses, initscr setupterm tigetstr)

13.1.8 CheckProg

success = $(CheckProg <prog>)

Checks whether the program <prog> exists in your path. Will use the ConfMsgChecking and ConfMsgResult functions to describe the test and the outcome.

13.2 Translating autoconf scripts

Some of the functions described above are very similar to the ones present in autoconf. Below is a brief translation table for such functions.

 AC MSG CHECKING is very similar to ConfMsgChecking function.
 AC MSG RESULT is very similar to ConfMsgResult function.

AC MSG WARN is very similar to ConfMsgWarn function.

AC MSG ERROR is very similar to ConfMsgError function.

AC TRY COMPILE is somewhat similar to TryCompileC function, except the TryCompileC function returns a boolean value and only works for C. Similarly,

AC TRY LINK is approximated by TryLinkC function, and

AC TRY RUN is approximated by TryRunC function.

13.3 Predefined configuration tests

A number of configuration tests are already included in the standard library. In order to use them in your project, simply open (see Section 4.7) the corresponding build file in your OMakefile and the tests will run the first time OMake is executed. Note that it is not a problem to open these files from more than one place in your project — if you do that, the test will still run only once.

13.3.1 NCurses library configuration

Add open configure/ncurses line to your OMakefile to get access to the

following autoconfiguration variables.

NCURSES AVAILABLE A boolean flag that would be set when both the curses.h header, the term.h header, and the ncurses library very found.

NCURSES TERMH IN NCURSES A boolean flag that would be set when term.h has to be included as <ncurses/term.h> instead of <term.h>.

NCURSES CFLAGS The CFLAGS to use when compiling ncurses code. Will

include -DNCURSES and -DTERMH_IN_NCURSES, respectively when NCURSES_AVAILABLE
and NCURSES_TERMH_IN_NCURSES are true.

NCURSES CLIBS The LDFLAGS to use when linking ncurses code. Will normally contain -lncurses when ncurses is found and remain empty otherwise.

13.3.2 ReadLine library configuration

Add open configure/readline line to your OMakefile to get access to the

following autoconfiguration variables.

READLINE AVAILABLE A boolean flag that would be set when both

the readline/readline.h header, the readline/history.h header, and the

readline library very found.

READLINE GNU A boolean flag that would be set when the GNU version of the readline library is found (as opposed to the BSD one).

READLINE CFLAGS The CFLAGS to use when compiling readline code.

Will include -DREADLINE_ENABLED and -DREADLINE_GNU, respectively when READLINE_AVAILABLE
and READLINE_GNU are true.

READLINE CLIBS The LDFLAGS to use when linking readline code. Will normally contain -lncurses -lreadline when readline is found and remain empty otherwise.

13.3.3 Snprintf configuration

Add open configure/snprintf line to your OMakefile to get access to the

following autoconfiguration variables.

SNPRINTF AVAILABLE A boolean flag telling whether the snprintf function is available in the standard C library.

206CHAPTER 13. AUTOCONFIGURATION FUNCTIONS AND VARIABLES

The OSH shell

OMake also includes a standalone command-line interpreter osh that can be used as an interactive shell. The shell uses the same syntax, and provides the same features on all platforms omake supports, including Win32.

14.1 Startup

On startup, osh reads the file ~/.oshrc if it exists. The syntax of this file is the same as an OMakefile. The following additional variables are significant.

prompt The prompt variable specifies the command-line prompt. It can be a simple string.

    prompt = osh>

Or you may choose to define it as a function of no arguments.

    prompt() =
        return $"<$(USER):$(HOST) $(homename $(CWD))>"

An example of the latter prompt is as follows.

    <jyh:kenai.yapper.org ~>cd links/omake
    <jyh:kenai.yapper.org ~/links/omake>

If you include any ”invisible” text in the prompt (such as various terminal escape sequences), they must be wrapped using the prompt-invisible function. For example, to create a bold prompt on terminals that support it, you can use the following.

    prompt =
       bold-begin = $(prompt-invisible $(tgetstr bold))
       bold-end = $(prompt-invisible $(tgetstr sgr0))
       value $(bold-begin)$"osh>"$(bold-end)

ignoreeof If the ignoreeof is true, then osh will not exit on a terminal end-of-file (usually ^D on Unix systems).

14.2 Aliases

Command aliases are defined by adding functions to the Shell. object. The following alias adds the -AF option to the ls command.

    Shell. +=
       ls(argv) =
          "ls" -AF $(argv)

Quoted commands do not undergo alias expansion. The quotation "ls" prevents the alias from being recursive.

14.3 Interactive syntax

The interactive syntax in osh is the same as the syntax of an OMakefile, with one exception in regard to indentation. The line before an indented block must have a colon at the end of the line. A block is terminated with a . on a line by itself, or ^D. In the following example, the first line if true has no body, because there is no colon.

   # The following if has no body
   osh>if true
   # The following if has a body
   osh>if true:
   if>       if true:
   if>          println(Hello world)
   if>          .
   Hello world

Note that osh makes some effort to modify the prompt while in an indented body, and it auto-indents the text.

The colon signifier is also allowed in files, although it is not required.

Appendix A

Synopsis

omake [-j <count>] [-k] [-p] [-P] [-n] [-s] [-S] [-w] [-t] [-u] [-U] [-R] [--verbose]
[--project] [--depend] [--progress] [--print-status] [--print-exit] [--print-dependencies]
[--show-dependencies <target>] [--all-dependencies] [--verbose-dependencies]
[--force-dotomake] [--dotomake <dir>] [--flush-includes] [--configure]
[--save-interval <seconds>] [--install] [--install-all] [--install-force]
[--version] [--absname] [--output-normal] [--output-postpone] [--output-only-errors]
[--output-at-end] filename... [var-definition...]

A.1 General usage

For Boolean options (for example, -s, --progress, etc.) the option can include a prefix --no, which inverts the usual sense of the option. For example, the option --progress means “print a progress bar,” while the option --no--progress means “do not print a progress bar.”

If multiple instances of an option are specified, the final option determines the behavior of OMake. In the following command line, the final --no-S cancels the earlier -S.

    % omake -S --progress --no-S

A.2 Output control

A.2.1 -s

-s

Never not print commands as they are executed (be “silent”).

A.2.2 -S

-S

Do not print commands as they are executed unless they produce output and/or fail. This is the default.

A.2.3 -w

-w

Print directory information in make format as commands are executed. This is mainly useful for editors that expect make-style directory information for determining the location of errors.

A.2.4 --progress
--progress

Print a progress indicator. This option is enabled by default when the OMake’s output (stdout) is on a terminal and disabled by default (except on Windows) when the OMake’s output is redirected.

A.2.5 --print-status
--print-status

Print status lines (the + and - lines).

A.2.6 --print-exit
--print-exit

Print termination codes when commands complete.

A.2.7 --verbose
--verbose

Make OMake very verbose. This option is equivalent to --no-S --print-status --print-exit VERBOSE=true

A.2.8 --output-normal
--output-normal

As rule commands are executed, relay their output to the OMake output right away. This is enabled by default, unless --output-postpone or

--output-only-errors is enabled.
A.2.9 --output-postpone
--output-postpone

When a rule finishes, print the output as a single block. This is useful in combination -j option (see Section A.3.12), where the output of multiple subprocesses can be garbled. The diversion is printed as a single coherent unit.

Note that enabling --output-postpone will by default disable the --output-normal option. This might be problematic if you have a command that decides to ask for interactive input. If the --output-postpone is enabled, but the --output-normal is not, the prompt of such a command will not be visible and it may be hard to figure out why the build appears “stuck”. You might also consider using the --progress flag (see Section A.2.4) so that you can see when the build is active.

A.2.10 --output-only-errors
--output-only-errors

Similar to --output-postpone, except that the postponed output from commands that were successful will be discarded. This can be useful in reducing unwanted output so that you can concentrate on any errors.

A.2.11 --output-at-end
--output-at-end

If any rules/commands fail, re-print the output of the failed commands when OMake finishes the build. This is especially useful when any of the -k, -p, or -P options are enabled.

This option is off by default. However, when -k is enabled — either explicitly or via one of the -p/-P options — --output-at-end will be enabled by default.

A.2.12 -o

-o [01jwWpPxXsS]

For brevity, the -o option is also provided to duplicate the above output options. The -o option takes a argument consisting of a sequence of characters. The characters are read from left-to-right; each specifies a set of output options. In general, an uppercase character turns the option on; a lowercase character turns the option off.

0 Equivalent to -s --output-only-errors

This option specifies that omake should be as quiet as possible. If any errors occur during the build, the output is delayed until the build terminates. Output from successful commands is discarded.

1 Equivalent to -S --progress --output-only-errors

This is a slightly more relaxed version of “quiet” output. The output from successful commands is discarded. The output from failed commands is printed immediately after the command complete. The output from failed commands is displayed twice: once immediately after the command completes, and again when the build completes. A progress bar is displayed so that you know when the build is active. Include the ‘p’ option if you want to turn off the progress bar (for example omake -o 1p).

2 Equivalent to --progress --output-postpone

The is even more relaxed, output from successful commands is printed. This is often useful for deinterleaving the output when using -j.

212 APPENDIX A. SYNOPSIS W Equivalent to -w w Equivalent to --no-w P Equivalent to --progress

p Equivalent to --no--progress

X Equivalent to --print-exit

x Equivalent to --no-print-exit

S Equivalent to -S s Equivalent to --no-S

A.3 Build options

A.3.1 -k

-k

Do not abort when a build command fails; continue to build as much of the project as possible. This option is implied by both -p and -P options. In turn, this option would imply the --output-at-end option.

A.3.2 -n

-n

This can be used to see what would happen if the project were to be built.

A.3.3 -p

-p

Watch the filesystem for changes, and continue the build until it succeeds. If this option is specified, omake will restart the build whenever source files are modified. Implies -k.

A.3.4 -P

-P

Watch the filesystem for changes forever. If this option is specified, omake will restart the build whenever source files are modified. Implies -k.

A.3.5 -R

-R

Ignore the current directory and build the project from its root directory. When omake is run in a subdirectory of a project and no explicit targets are given on the command line, it would normally only build files within the current directory and its subdirectories (more precisely, it builds all the .DEFAULT targets in the current directory and its subdirectories). If the -R option is specified, the build is performed as if omake were run in the project root.

In other words, with the -R option, all the relative targets specified on the command line will be taken relative to the project root (instead of relative to the current directory). When no targets are given on the command line, all the .DEFAULT targets in the project will be built (regardless of the current directory).

A.3.6 -t

-t

Update the omake database to force the project to be considered up-to-date.

A.3.7 -U

-U

Do not trust cached build information. This will force the entire project to be rebuilt.

A.3.8 --depend
--depend

Do not trust cached dependency information. This will force files to be rescanned for dependency information.

A.3.9 --configure
--configure

Re-run static. sections of the included omake files, instead of trusting the cached results.

A.3.10 --force-dotomake
--force-dotomake

Always use the $HOME/.omake for the .omc cache files.

A.3.11 --dotomake
--dotomake <dir>

Use the specified directory instead of the $HOME/.omake for the placement of the .omc cache files.

A.3.12 -j

-j <count>

Run multiple build commands in parallel. The count specifies a bound on the number of commands to run simultaneously. In addition, the count may specify servers for remote execution of commands in the form server=count.

For example, the option -j 2:small.host.org=1:large.host.org=4 would

specify that up to 2 jobs can be executed locally, 1 on the server small.host.org and 4 on large.host.org. Each remote server must use the same filesystem location for the project.

Remote execution is currently an experimental feature. Remote filesystems like NFS do not provide adequate file consistency for this to work.

A.3.13 --print-dependencies
--print-dependencies

Print dependency information for the targets on the command line.

A.3.14 --show-dependencies
--show-dependencies <target>

Print dependency information if the target is built.

A.3.15 --all-dependencies
--all-dependencies
   If either of the options --print-dependencies or --show-dependencies

is in effect, print transitive dependencies. That is, print all dependencies re-

cursively. If neither option --print-dependencies, --show-dependencies is

specified, this option has no effect.

A.3.16 --verbose-dependencies
--verbose-dependencies
   If either of the options --print-dependencies or --show-dependencies

is in effect, also print listings for each dependency. The output is very verbose, consider redirecting to a file. If neither option --print-dependencies, --show-dependencies is specified, this option has no effect.

A.3.17 --install
--install

Install default files OMakefile and OMakeroot into the current directory. You would typically do this to start a project in the current directory.

A.3.18 --install-all
--install-all

In addition to installing files OMakefile and OMakeroot, install default OMakefiles into each subdirectory of the current directory. cvs(1) rules are used for filtering the subdirectory list. For example, OMakefiles are not copied into directories called CVS, RCCS, etc.

A.3.19 --install-force
--install-force

Normally, omake will prompt before it overwrites any existing OMakefile. If this option is given, all files are forcibly overwritten without prompting.

A.3.20 --absname
--absname

Filenames should expand to absolute pathnames.

N.B. This is an experimental option. It may become deprecated.

A.3.21 variable definition
name=[value]

omake variables can also be defined on the command line in the form name=value. For example, the CFLAGS variable might be defined on the command line with

the argument CFLAGS="-Wall -g".

A.4 Additional options

In addition, omake supports a number of debugging flags on the command line. Run omake --help to get a summary of these flags.

A.5 Environment variables

A.5.1 OMAKEFLAGS

If defines, the OMAKEFLAGS should specify a set of options exactly as they are specified on the command line.

A.5.2 OMAKELIB

If defined, the OMAKELIB environment variable should refer to the installed location of the OMake standard library. This is the directory that contains

Pervasives.om etc. On a Unix system, this is often /usr/lib/omake or /usr/local/lib/omake,

and on Win32 systems it is often c:\Program Files\OMake\lib.

If not defined, omake uses the default configured location. You should normally leave this unset.

A.6 Functions

A.6.1 OMakeFlags

The OMakeFlags function can be used within an OMakefile to modify the set of options. The options should be specified exactly as they are on the command line. For example, if you want some specific project to be silent and display a progress bar, you can add the following line to your OMakefile.

    OMakeFlags(-S --progress)

For options where it makes sense, the options are scoped like variables. For example, if you want OMake to be silent for a single rule (instead of for the entire project), you can use scoping the restrict the range of the option.

    section
        # Do not display command output when foo is constructed
        OMakeFlags(-S)
        foo: fee
           echo "This is a generated file" > foo
           cat fee >> foo
           chmod 555 foo

A.7 Option processing

When omake is invoked, the options are processed in the following order.

1. All options specified by the OMAKEFLAGS environment variable are defined globally.

2. All options from the command line are defined globally.

3. Any individual calls the the OMakeFlags function modify the options locally.

A.8 .omakerc

If the $(HOME)/.omakerc exists, it is read before any of the OMakefiles in your project. The .omakerc file is frequently used for user-specific customization. For example, instead of defining the OMAKEFLAGS environment variable, you could add a line to your .omakerc.

    $(HOME)/.omakerc:
        # My private options
        OMakeFlags(-S --progress)

218 APPENDIX A. SYNOPSIS

Appendix B

OMake grammar

B.1 OMake lexical conventions

The OMake language is based on the language for GNU/BSD make, where there are few lexical conventions. Strictly speaking, there are no keywords, and few special symbols.

B.1.1 Comments

Comments begin with the # character and continue to the end-of-line. Text within a comment is unrestricted.

Examples.

# This is a comment
# This $comment contains a quote " character

B.1.2 Special characters

The following characters are special in some contexts.

$    (    )    ,    .   =    :    "    '    `    \    #

• $ is used to denote a variable reference, or function application.

• Parentheses ), ( are argument deliminters.

• The command , is an argument separator.

• The period symbol . is a name separator.

• The equality symbol = denotes a definition.

• The colon symbol : is used to denote rules, and (optionally) to indicate that an expression is followed by an indented body.

• The quotation symbols " and ' delimit character strings.

• The symbol # is the first character of a constant.

• The escape symbol \ is special only when followed by another special character. In this case, the special status of the second character is removed, and the sequence denotes the second character. Otherwise, the \ is not special.

Examples:

– \$: the $ character (as a normal character).

– \#: the # character (as a normal character).

– \\: the \ character (as a normal character).

   – c\:\Windows\moo\#boo: the string c:\Windows\moo#boo.

B.1.3 Identifiers

Identifiers (variable names) are drawn from the ASCII alphanumeric characters as well as _, -, ~, @. Case is significant; the following identifiers are distinct: FOO, Foo, foo. The identifier may begin with any of the valid characters, including digits.

Using egrep notation, the regular expression for identifiers is defined as follows.

identifier ::= [-@~_A-Za-z0-9]+

The following are legal identifiers.

Xyz    hello_world    seventy@nine
79-32  Gnus~Gnats     CFLAGS

The following are not legal identifiers.

x+y    hello&world

B.1.4 Command identifiers

The following words have special significance when they occur as the first word of a program line. They are not otherwise special.

case     catch  class    declare    default
do       else   elseif   export     extends
finally  if     import   include    match
open     raise  return   section    switch
try      value  when     while

B.1.5 Variable references

A variable reference is denoted with the $ special character followed by an identifier. If the identifier name has more than one character, it must be enclosed in parentheses. The parenthesized version is most common. The following are legal variable references.

    $(Xyz)    $(hello_world)   $(seventy@nine)
    $(79-32)  $(Gnus~Gnats)    $(CFLAGS)

Single-character references also include several additional identifiers, including &*<^?][. The following are legal single-character references.

   $@   $&   $*   $<   $^$+   $?   $[   $]
   $A   $_$a   $b   $x   $1   $2   $3

Note that a non-parenthesized variable reference is limited to a single character, even if it is followed by additional legal identifier charqcters. Suppose the value of the $x variable is 17. The following examples illustrate evaluation.

    $x           evaluates to    17
    foo$xbar     evaluates to    foo17bar
    foo$(x)bar   evaluates to    foo17bar

The special sequence $$ represents the character literal $. That is, the twocharacter sequences \$ and $$ are normally equalivalent.

B.1.6 String constants

Literal strings are defined with matching string delimiters. A left string delimiter begins with the dollar-sign $, and a non-zero number of single-quote or double-quote characters. The string is terminated with a matching sequence of quotation symbols. The delimiter quotation may not be mixed; it must contain only single-quote characters, or double-quote characters. The following are legal strings.

    $'Hello world'
    $"""printf("Hello world\n")"""
    $''''
Large "block" of
text # spanning ''multiple'' lines''''

The string delimiters are not included in the string constant. In the singlequote form, the contents of the string are interpreted verbatim–there are no special characters.

The double-quote form permits expression evaluation within the string, denoted with the $ symbol. The following are some examples.

X = Hello
Y = $""$X world""             # Hello world
Z = $'''$X world'''           # $X world
I = 3
W = $"6 > $(add $I, 2)"       # 6 > 5

Note that quotation symbols without a leading $ are not treated specially by OMake. The quotation symbols is included in the sequence.

osh>println('Hello world')
'Hello world'
osh>println($'Hello world')
Hello world
osh>X = Hello
- : "Hello" : Sequence
osh>println('$X world')
Hello world

B.2 The OMake grammar

OMake programs are constructed from expressions and statements. Generally, an input program consists of a sequence of statements, each of which consists of one or more lines. Indentation is significant–if a statement consists of more than one line, the second and remaining lines (called the body) are usually indented relative to the first line.

B.2.1 Expressions

The following table lists the syntax for expressions.

expr ::=

(empty)

– Text (see note)

— text
— string-literal

– Applications

— dollar <char>
— dollar ( pathid args )

– Concatenation

— expr expr
dollar ::= $ — $` — $,
pathid ::=

id

— pathid . id
arg ::= expr – excluding special characters )(,)
args ::= (empty) — arg, ..., arg

An expression is a sequence composed of text, string-literals, variables references and function applications. Text is any sequence of non-special characters.

B.2.1.1 Inline applications

An application is the application of a function to zero-or-more arguments. Inline applications begin with one of the “dollar” sequences $, $`, or $,. The application itself is specified as a single character (in which case it is a variable reference), or it is a parenthesized list including a function identifier pathid, and zero-or-more comma-separated arguments args. The arguments are themselves a variant of the expressions where the special character )(, are not allowed (though any of these may be made non-special with the \ escape character). The following are some examples of valid expressions.

   • xyz abc

The text sequence “xyz abc”

   • xyz$wabc

A text sequence containing a reference to the variable w.

   • $(addsuffix .c, $(FILES))

An application of the function addsuffix, with first argument .c, and second argument $(FILES).

   • $(a.b.c 12)

This is a method call. The variable a must evaluate to an object with a field b, which must be an object with a method c. This method is called with argument 12.

The additional dollar sequences specify evaluation order, $` (lazy) and $, (eager), as discussed in the section on dollar modifiers (Section B.3).

B.2.2 Statements and programs

The following table lists the syntax of statements and programs.

params ::= (empty) — id, ..., id
target ::= expr – excluding special character :
program ::= stmt <eol> ... <eol> stmt
stmt ::=

– Special forms

— command expr optcolon-body
— command ( args ) optcolon-body
— catch id ( id ) optcolon-body
— class id ... id

– Variable definitions

— pathid {+}= expr
— pathid {+}= <eol> indented-body
— pathid[] {+}= expr
— pathid[] {+}= <eol> indented-exprs

– Functions

— pathid(args) optcolon-body
— pathid(params) = <eol> indented-body

– Objects

— pathid . {+}= <eol> indented-body

– Rules

— target : target rule-options <eol> indented-body
— target :: target rule-options <eol> indented-body
— target : target : target rule-options <eol> indented-body
— target :: target : target rule-options <eol> indented-body

– Shell commands

— expr
indented-body ::= (empty)
— indented-stmt <eol> ... <eol> indented-stmt
indented-exprs ::= (empty)
— indented-expr <eol> ... <eol> indented-expr
optcolon-body ::= (empty)
— <eol> indented-body
— : <eol> indented-body
rule-option ::= :id: target
rule-options ::= (empty)
— rule-options rule-option

B.2.2.1 Special forms

The special forms include the following.

Conditionals (see the section on conditionals — Section 4.9). The if command should be followed by an expression that represents the condition, and an indented body. The conditional may be followed by elseif and else blocks.

if expr
    indented-body
elseif expr
    indented-body
...
else
    indented-body

matching (see the section on matching — Section 4.10). The switch and match commands perform pattern-matching. All cases are optional. Each case may include when clauses that specify additional matching conditions.

match(expr)
case expr
   indented-body
when expr
   indented-body
...
case expr
   indented-body
default
   indented-body

Exceptions (see also the try function documentation). The try command introduces an exception handler. Each name is the name of a class. All cases, including catch, default, and finally are optional. The catch and default clauses contain optional when clauses.

try
    indented-body
catch name1(id1)
    indented-body
when expr
    indented-body
...
catch nameN(idN)
    indented-body
default
    indented-body
finally
    indented-body

The raise command is used to raise an exception.

    raise expr

section (see the section description in Section 4.8). The section command introduces a new scope.

    section
        indented-body

include, open (see also Section 4.7). The include command performs file inclusion. The expression should evaluate to a file name.

The open form is like include, but it performs the inclusion only if the inclusion has not already been performed. The open form is usually used to include library files. [jyh– this behavior will change in subsequent revisions.]

    include expr
    open expr

return (see the description of functions in Section 4.5). The return command terminates execution and returns a value from a function.

    return expr

value (see the description of functions in Section 4.5). The value command is an identity. Syntactically, it is used to coerce a n expression to a statement.

    value expr

export (see the section on scoping — Section 5.3). The export command exports a environment from a nested block. If no arguments are given, the entire environment is exported. Otherwise, the export is limited to the specified identifiers.

    export expr

while (see also the while function description). The while command introduces a while loop.

    while expr
        indented-body

class, extends (see the section on objects — Section 4.11). The class command specifies an identifier for an object. The extends command specifies a parent object.

    class id
    extends expr

B.2.2.2 Variable definitions

See the section on variables (Section 4.1). The simplest variable definition has the following syntax. The = form is a new definition. The += form appends the value to an existing definition.

id = expr
id += expr
osh> X = 1
- : "1" : Sequence
osh> X += 7
- : "1" " " "7" : Sequence

A multi-line form is allowed, where the value is computed by an indented body.

id {+}=
    indented-body
osh> X =
         Y = HOME
         println(Y is $Y)
         getenv($Y)
Y is HOME
- : "/home/jyh" : Sequence

The name may be qualified qith one of the public, prtected, or private modifiers. Public variables are dynamically scoped. Protected variables are fields in the current object. Private variables are statically scoped.

[jyh: revision 0.9.9 introduces modular namespaces; the meaning of these qualifiers is slightly changed.]

public.X = $(addsuffix .c, 1 2 3)
protected.Y = $(getenv HOME)
private.Z = $"Hello world"

B.2.2.3 Applications and function definitions

See the section on functions (Section 4.5). A function-application statement is specified as a function name, followed a parenthesized list of comma-separated arguments.

osh> println($"Hello world")
osh> FILES = 1 2 3
- : 1 2 3
osh> addsuffix(.c, $(FILES))

- : 1.c 2.c 3.c
# The following forms are equivalent
osh> value $(println $"Hello world")
osh> value $(addsuffix .c, $(FILES))
- : 1.c 2.c 3.c

If the function application has a body, the body is passed (lazily) to the function as its first argument. [jyh: in revision 0.9.8 support is incomplete.] When using osh, the application must be followed by a colon : to indicate that the application has a body.

# In its 3-argument form, the foreach function takes
# a body, a variable, and an array.  The body is evaluated
# for each element of the array, with the variable bound to
# the element value.
#
# The colon is required only for interactive sessions.
osh> foreach(x, 1 2 3):
        add($x, 1)
- : 2 3 4

Functions are defined in a similar form, where the parameter list is specified as a comma-separated list of identifiers, and the body of the function is indented.

osh> f(i, j) =
        add($i, $j)
- : <fun 2>
osh> f(3, 7)
- : 10 : Int

B.2.2.4 Objects

See the section on objects (Section 4.11). Objects are defined as an identifier with a terminal period. The body of the object is indented.

Obj. =
    class Obj
    X = 1
    Y = $(sub $X, 12)
    new(i, j) =
       X = $i
       Y = $j
       value $(this)
    F() =
       add($X, $Y)
    println($Y)

The body of the object has the usual form of an indented body, but new variable definitions are added to the object, not the global environment. The object definition above defines an object with (at least) the fields X and Y, and methods new and F. The name of the object is defined with the class command

as Obj.

The Obj itself has fields X = 1 and Y = -11. The new method has the typical form of a constructor-style method, where the fields of the object are initialized to new values, and the new object returned ($(this) refers to the current object).

The F method returns the sum of the two fields X and Y.

When used in an object definition, the += form adds the new definitions to an existing object.

   pair. =
      x = 1
      y = 2
   pair. +=
      y = $(add $y, 3)
   # pair now has fields (x = 1, and y = 5)

The extends form specifies inheritance. Multiple inheritance is allowed. At evaluation time, the extends directive performs inclusion of the entire parent object.

   pair. =
      x = 1
      y = 2
   depth. =
      z = 3
      zoom(dz) =
         z = $(add $z, $(dz))
         return $(this)
   triple. =
      extends $(pair)
      extends $(depth)
      crazy() =
         zoom($(mul $x, $y))

In this example, the triple object has three fields x, y, and z; and two methods zoom and crazy.

B.2.2.5 Rules

See the chapter on rules (Chapter 7). A rule has the following parts.

1. A sequence of targets;

2. one or two colons;

3. a sequence of dependencies and rule options;

4. and an indented body.

The targets are the files to be built, and the dependencies are the files it depends on. If two colons are specified, it indicates that there may be multiple rules to build the given targets; otherwise only one rule is allowed.

If the target contains a % character, the rule is called implicit, and is considered whenever a file matching that pattern is to be built. For example, the following rule specifies a default rule for compiling OCaml files.

    %.cmo: %.ml %.mli
       $(OCAMLC) -c $<

This rule would be consulted as a default way of building any file with a .cmo suffix. The dependencies list is also constructed based on the pattern match. For example, if this rule were used to build a file foo.cmo, then the dependency

list would be foo.ml foo.mli.

There is also a three-part version of a rule, where the rule specification has three parts.

    targets : pattern : dependencies rule-options
       indented-body

In this case, the pattern must contain a single % character. However this is considered to be a sequence of explicit rules, where each target is matched against the pattern, and a new rule is computed based on the pattern match. For example, the following rule specifies how to build the explicit targets a.cmo

and b.cmo.
    a.cmo b.cmo: %.cmo: %.ml %.mli
       $(OCAMLC) -c $<

This example is equivalent to the following two-rule sequence.

    a.cmo: a.ml a.mli
       $(OCAMLC) -c $<
    b.cmo: b.ml b.mli
       $(OCAMLC) -c $<

There are several special targets, including the following.

• .PHONY : declare a “phony” target. That is, the target does not correspond to a file.

• .ORDER : declare a rule for dependency ordering.

B.2. THE OMAKE GRAMMAR 231 • .INCLUDE : define a rule to generate a file for textual inclusion.

• .SUBDIRS : specify subdirectories that are part of the project.

• .SCANNER : define a rule for dependency scanning. There are several rule options.

• :optional: dependencies the subsequent dependencies are optional, it is acceptable if they do not exist.

• :exists: dependencies the subsequent dependencies must exist, but changes to not affect whether this rule is considered out-of-date.

• :effects: targets the subsequent files are side-effects of the rule. That is, they may be created and/or modified while the rule is executing. Rules with overlapping side-effects are never executed in parallel.

• :scanner: name the subsequent name is the name of the .SCANNER rule for the target to be built.

• :value: expr the expr is a “value” dependency. The rule is considered out-of-date whenever the value of the expr changes.

Several variables are defined during rule evaluation.

• $* : the name of the target with the outermost suffix removed.

• $> : the name of the target with all suffixes removed.

• $@ : the name of the target.

• $^: the explicit file dependencies, sorted alphabetically, with duplicates removed.

• $+ : all explicit file dependencies, with order preserved.

• $< : the first explicit file dependency.

• $& : the free values of the rule (often used in :value: dependencies).

B.2.2.6 Shell commands See the chapter on shell commands (Chapter 10).

While it is possible to give a precise specification of shell commands, the informal description is simpler. Any non-empty statement where each prefix is not one of the other statements, is considered to be a shell command. Here are some examples.

    ls                                 -- shell command
    echo Hello world > /dev/null       -- shell command
    echo(Hello world)                  -- function application
    echo(Hello world) > /dev/null      -- syntax error
    echo Hello: world                  -- rule
    X=1 getenv X                       -- variable definition
    env X=1 getenv X                   -- shell command
    if true                            -- special form
    \if true                           -- shell command
    "if" true                          -- shell command

B.3 Dollar modifiers

Inline applications have a function and zero-or-more arguments. Evaluation is normally strict: when an application is evaluated, the function identifier is evaluated to a function, the arguments are then evaluated and the function is called with the evaluated arguments.

The additional “dollar” sequences specify additional control over evaluation. The token $` defines a “lazy” application, where evaluation is delayed until a value is required. The $, sequence performs an “eager” application within a lazy context.

To illustrate, consider the expression $(addsuffix .c, $(FILES)). The addsuffix function appends its first argument to each value in its second argument. The following osh interaction demonstrates the normal bahavior.

osh> FILES[] = a b c
- : <array a b c>
osh> X = $(addsuffix .c, $(FILES))
- : <array ...>
osh> FILES[] = 1 2 3 # redefine FILES
- : <array 1 2 3>
osh> println($"$X")  # force the evaluation and print
a.c b.c c.c

When the lazy operator $` is used instead, evaluation is delayed until it is printed. In the following sample, the value for X has changed to the $(apply ..) form, but otherwise the result is unchanged because it it printed immediately.

osh> FILES[] = a b c
- : <array a b c>
osh> SUF = .c
- : ".c"
osh> X = $`(addsuffix $(SUF), $(FILES))
- : $(apply global.addsuffix ...)
osh> println($"$X")  # force the evaluation and print
a.c b.c c.c

However, consider what happens if we redefine the FILES variable after the definition for X. In the following sample, the result changes because evaluation occurs after the values for FILES has been redefined.

osh> FILES[] = a b c
- : <array a b c>
osh> SUF = .c
- : ".c"
osh> X = $`(addsuffix $(SUF), $(FILES))
- : $(apply global.addsuffix ...)
osh> SUF = .x
osh> FILES[] = 1 2 3
osh> println($"$X")  # force the evaluation and print
1.x 2.x 3.x

In some cases, more explicit control is desired over evaluation. For example, we may wish to evaluate SUF early, but allow for changes to the FILES variable. The $,(SUF) expression forces early evaluation.

osh> FILES[] = a b c
- : <array a b c>
osh> SUF = .c
- : ".c"
osh> X = $`(addsuffix $,(SUF), $(FILES))
- : $(apply global.addsuffix ...)
osh> SUF = .x
osh> FILES[] = 1 2 3
osh> println($"$X")  # force the evaluation and print
1.c 2.c 3.c

Index

–absname, 215 –all-dependencies, 214 –configure, 213 –depend, 213 –dotomake, 214 –force-dotomake, 213 –install, 215 –install-all, 215 –install-force, 215 –output-at-end, 211 –output-normal, 210 –output-only-errors, 211 –output-postpone, 210 –print-dependencies, 214 –print-exit, 210 –print-status, 210 –progress, 210 –show-dependencies, 214 –verbose, 210 –verbose-dependencies, 214 -P, 212 -R, 213 -S, 209 -U, 213 -j, 214 -k, 212 -n, 212 -o, 211 -p, 212 -s, 209 -t, 213 -w, 210 .BUILDORDER, 123 .BUILD BEGIN, 181 .BUILD FAILURE, 181 .BUILD SUCCESS, 181 .DEFAULT, 85, 90 .INCLUDE, 86 .ORDER, 123 .PHONY, 64, 86, 89–91 .RULE, 64 .SCANNER, 37, 83, 88 .SUBDIRS, 30, 85 .SUBDIRS bodies, 41 .omakerc, 217 :effects:, 82 :exists:, 81 :scanner:, 84 :value:, 82 $&, 231 $*, 79, 231 $+, 79, 231 $¡, 79, 231 $¿, 231 $@, 79, 231 $ˆ , 79, 231

ABORT ON COMMAND ERROR,

186 ABORT ON DEPENDENCY ERRORS,

194 absname, 119 AC MSG CHECKING, 203 AC MSG ERROR, 204 AC MSG RESULT, 203 AC MSG WARN, 204 AC TRY COMPILE, 204 AC TRY LINK, 204 AC TRY RUN, 204 accept, 143 add, 112 add-project-directories, 132 add-wrapper, 107 addprefix, 107 addsuffix, 105 addsuffixes, 106 aliases, 208 and, 95 apply, 114 applya, 114 AR, 188 Array, 172 array, 101 arrays, 48 AS, 188 ASFLAGS, 188 ASOUT, 188 asr, 112 awk, 38, 150

basename, 118 bg, 167 BIBTEX, 199 bind, 143 break, 112 build model, 29 BUILD SUMMARY, 94 BYTE ENABLED, 193

c-escaped, 104 CAMLP4, 192 capitalize, 109 case, 96 cat, 147 cats and dogs, 35 CC, 188 CCOUT, 188 cd, 166 CFLAGS, 188 CGeneratedFiles, 189 Channel, 174 CheckCHeader, 203 CheckCLib, 203 CheckProg, 203 chmod, 131 chown, 131 CL FOUND, 187 class, 226 classes, 54 close, 137 cmp-versions, 183 concat, 102 conditionals, 52 ConfMsgChecking, 201 ConfMsgError, 202 ConfMsgFound, 202 ConfMsgResult, 201 ConfMsgWarn, 202 ConfMsgYesNo, 202 connect, 143 CPP, 188 CProgram, 190 CProgramCopy, 191 CProgramInstall, 191 create-lazy-map, 115 create-map, 115 CWD, 186 CXX, 188 CXXFLAGS, 188 CXXProgram, 191 CXXProgramInstall, 191

decode-uri, 104 default, 96 DefineCommandVars, 183 defined, 98 defined-env, 99 dependencies, 183 dependencies-all, 183 dependencies-proper, 183 digest, 121 digest-path, 121 Dir, 174 dir, 117 dirname, 119 dirof, 119 DIRSEP, 186 div, 112 dup, 139 dup2, 139 DVIPDFM, 199 DVIPDFMFLAGS, 199 DVIPS, 199 DVIPSFLAGS, 199

echo, 166 else, 96, 225 elseif, 96, 225 EMPTY, 186 encode-uri, 104 eprint, 144 eprintln, 144 eprintv, 145 eprintvln, 145 eq, 113 equal, 95 Exception, 175 EXE, 187 exists-in-path, 121 exit, 98 export, 64, 110, 226 EXT ASM, 187 EXT LIB, 186 EXT OBJ, 186 extends, 226

fg, 167 fgets, 144 File, 174 file, 117 file-check-sort, 124 file-exists, 122 file-sort, 123 filter, 108 filter-exists, 122 filter-out, 109 filter-proper-targets, 122 filter-targets, 122 find, 135 find-build-targets, 184 find-in-path, 121 find-targets-in-path, 123 find-targets-in-path-optional, 123 Float, 171 float, 112 flush, 138 fopen, 136 foreach, 115 fprint, 144 fprintln, 144 fprintv, 145 fprintvln, 145 fsubst, 151 fullname, 119 Fun, 173 fun, 114 functions, 49

GCC FOUND, 187 ge, 113 get-registry, 100 getchar, 144 getenv, 99 getgrgid, 160 getgrnam, 160 gethostbyname, 141 getprotobyname, 141 getpwents, 160 getpwnam, 160 getpwuid, 160 gets, 144 getservbyname, 142 gettimeofday, 162 getvar, 100 glob, 126 gr gid, 160 gr group, 160 gr mem, 160 gr name, 160 grep, 148 Group, 160 gt, 113 GXX FOUND, 187

history, 167 HOME, 94 homename, 120 HOST, 94 Host, 141 html-escaped, 104 html-pre-escaped, 104 html-string, 105

id-escaped, 104 if, 52, 96, 225 ignoreeof, 208 in, 118 InChannel, 175 include, 50, 226 INCLUDES, 188 InetAddr, 141 inheritance, 55 INSTALL, 186 Int, 171 int, 112 intersection, 108 intersects, 108

jobs, 167 join, 103

kill, 167

land, 112 LATEX, 198 LaTeXDocument, 199 LaTeXDocumentCopy, 200 LaTeXDocumentInstall, 200 LATEXFLAGS, 199 LD, 188 LDFLAGS, 188 LDOUT, 188 le, 113 length, 102 LEX, 189 lex, 152 lex-search, 153 Lexer, 154 LIB FOUND, 187 LIBS, 188 link, 130 link-order sorting, 123 listen, 143 lnot, 112 LocalCGeneratedFiles, 189 LocalOCamlGeneratedFiles, 195 LocalTeXGeneratedFiles, 200 Location, 175 lockf, 140 lor, 112 lowercase, 110 ls, 127 lseek, 138 lsl, 112 lsr, 112 lstat, 129 lt, 113 lxor, 112

MACHINE, 94 MAKEINDEX, 199 Map, 170 mapprefix, 107 mapsuffix, 105 match, 52, 96, 225 max, 112 mem, 107 MENHIR AVAILABLE, 192 MENHIR ENABLED, 193 MENHIR FLAGS, 194 min, 112 mkdir, 128 mkfifo, 140 mod, 112 mul, 112

NATIVE ENABLED, 193 NCURSES AVAILABLE, 204 NCURSES CFLAGS, 204 NCURSES CLIBS, 204 NCURSES TERMH IN NCURSES,

204 neg, 112 Node, 174 NODENAME, 93 not, 95 nth, 102 nth-hd, 102 nth-tl, 103 Number, 171

Object, 169 objects, 54 ocaml-escaped, 104 OCAML BYTE LINK FLAGS, 194 OCAML CLIBS, 194 OCAML LIB FLAGS, 194 OCAML LIBS, 194 OCAML LINK FLAGS, 194 OCAML NATIVE LINK FLAGS, 194 OCAML OTHER LIBS, 194 OCAMLC, 192 OCAMLCFLAGS, 194 OCAMLDEP, 192 ocamldep-omake, 196 OCAMLDEP MODULES, 193 OCAMLDEP MODULES AVAILABLE,

192 OCAMLDEP MODULES ENABLED,

193 OCAMLDEPFLAGS, 194 OCAMLFIND, 193 OCAMLFIND EXISTS, 192 OCAMLFINDFLAGS, 193 OCAMLFLAGS, 194 OCamlGeneratedFiles, 195 OCAMLINCLUDES, 193 OCAMLLEX, 192 OCAMLLEXFLAGS, 192 OCamlLibrary, 197 OCamlLibraryCopy, 197 OCamlLibraryInstall, 197 OCAMLLINK, 193 OCAMLMKTOP, 193 OCAMLOPT, 192 OCAMLOPT EXISTS, 192 OCAMLOPTFLAGS, 194 OCAMLOPTLINK, 193 OCamlPackage, 197 OCAMLPACKS, 193 OCAMLPPFLAGS, 194 OCamlProgram, 197 OCamlProgramCopy, 198 OCamlProgramInstall, 198 OCAMLYACC, 192 OCAMLYACCFLAGS, 192 OMAKE VERSION, 93 OMakefile, 26, 30 OMAKEFLAGS, 215 OMakeFlags, 182 OMAKELIB, 216 OMAKEPATH, 93 OMakeroot, 26, 30, 185 OMakeVersion, 182 open, 50, 226 open-in-string, 136 open-out-string, 136 or, 95 OS VERSION, 94 OSTYPE, 93 out-string, 136 OutChannel, 175

Passwd, 159 PATHSEP, 186 PDFLATEX, 199 PDFLATEXFLAGS, 199 PID, 94 pipe, 139 Position, 175 print, 144 println, 144 printv, 145 printvln, 145 private., 56 project-directories, 184 prompt, 207 prompt-invisible, 161 prompt-invisible-begin, 161 prompt-invisible-end, 161 protected., 57 Protocol, 141 public., 58 pw dir, 159 pw gecos, 159 pw gid, 159 pw name, 159 pw passwd, 159 pw shell, 159 pw uid, 159

quotations, 48 quote, 105 quote-argv, 105

raise, 98 random, 112 random-init, 112 read, 137 READLINE AVAILABLE, 204 READLINE CFLAGS, 205 READLINE CLIBS, 205 READLINE GNU, 205 readlink, 130 regular expressions, 145 rehash, 120 remove-project-directories, 132 removeprefix, 106 removesuffix, 106 rename, 129 replacesuffixes, 106 return, 49, 226 rev, 103 rewind, 138 ROOT, 186 rootname, 119 Rule, 173 rule, 185 rule, options, 81 rule, scoping, 87 rules, bounded implicit, 80 rules, implicit, 80 RunCProg, 202 RuntimeException, 176

scan, 148 SCANNER MODE, 186 section, 50, 81, 226 section rule, 81 select, 140 Sequence, 172 Service, 142 set, 107 set-close-on-exec-mode, 139 set-diff, 108 set-nonblock, 139 setenv, 99 setvar, 101 Shell, 176 shell, 110 SNPRINTF AVAILABLE, 205 socket, 142 sorting (link-order), 123 split, 101 Stat, 128 stat, 129 stat-reset, 122 static., 58 StaticCLibrary, 189 StaticCLibraryCopy, 190 StaticCLibraryInstall, 190 StaticCObject, 190 StaticCObjectCopy, 190 StaticCObjectInstall, 190 StaticCXXLibrary, 191 StaticCXXLibraryCopy, 191 StaticCXXLibraryInstall, 191 stderr, 136 stdin, 136 STDLIB, 93 stdout, 136 STDROOT, 186 stop, 167 String, 173 string, 103 string-escaped, 104 sub, 112 subdirs, 127 subrange, 103 suffix, 120 switch, 52, 96 symlink, 130 SYSNAME, 93 system, 110

Target, 173 target, 184 target-exists, 122 target-is-proper, 122 TARGETS, 94 tell, 138 test, 133 TETEX2 ENABLED, 198 TEXDEPS, 199 TeXGeneratedFiles, 200 TEXINPUTS, 199 TEXVARS, 199 tgetstr, 161 tmpfile, 118 truncate, 131 try, 97, 225 TryCompileC, 202 TryLinkC, 202 TryRunC, 202 uge, 113 ugt, 113 ule, 113 ult, 113 umask, 131 UnbuildableException, 176 uncapitalize, 109 unlink, 129 unsetenv, 100 uppercase, 109 USE OCAMLFIND, 192 USEPDFLATEX, 199 USER, 94

value, 49, 226 variable definition, 215 VERBOSE, 94 VerboseCheckCHeader, 203 VerboseCheckCLib, 203 vmount, 27, 132

wait, 167 where, 120 which, 120 while, 111, 226 write, 137

xterm-escape, 161 xterm-escape-begin, 161 xterm-escape-end, 161

YACC, 189

Appendix C

References

C.1 See Also

omake(1) (Chapter 1), osh(1) (Chapter 14), make(1)

C.2 Version

Version: 0.9.8.1 of 15th March 2007.

C.3 License and Copyright

\begin{matrix}c \\ ◯ \end{matrix} 2003-2006, Mojave Group, Caltech

This program is free software; you can redistribute it and/or modify it under the terms of the GNU General Public License as published by the Free Software Foundation; version 2 of the License.

This program is distributed in the hope that it will be useful, but WITHOUT ANY WARRANTY; without even the implied warranty of MERCHANTABIL-ITY or FITNESS FOR A PARTICULAR PURPOSE. See the GNU General Public License for more details.

You should have received a copy of the GNU General Public License along with this program; if not, write to the Free Software Foundation, Inc., 675 Mass Ave, Cambridge, MA 02139, USA.

C.4 Author

Jason Hickey, Aleksey Nogin, et. al. Caltech 256-80 Pasadena, CA 91125, USA

Email: omake-devel@metaprl.org
WWW: http://www.cs.caltech.edu/~jyh and http://nogin.org/